### v3

# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

Abstract
Research in robotic planning with temporal logic specifications, such as syntactically co-safe Linear Temporal Logic (sc-LTL), has relied on single formulas. However, as task complexity increases, sc-LTL formulas become lengthy, making them difficult to interpret and generate, and straining the computational capacities of planners. To address this, we introduce a hierarchical structure to sc-LTL specifications with both syntax and semantics, proving it to be more expressive than flat counterparts. We conducted a user study that compared the flat sc-LTL with our hierarchical version and found that users could more easily comprehend complex tasks using the hierarchical structure. We develop a search-based approach to synthesize plans for multi-robot systems, achieving simultaneous task allocation and planning. This method approximates the search space by loosely interconnected sub-spaces, each corresponding to an sc-LTL specification. The search primarily focuses on a single sub-space, transitioning to another under conditions determined by the decomposition of automatons. We develop multiple heuristics to significantly expedite the search. Our theoretical analysis, conducted under mild assumptions, addresses completeness and optimality. Compared to existing methods used in various simulators for service tasks, our approach improves planning times while maintaining comparable solution quality.
Index Terms:
Formal Methods in Robotics and Automation; Planning, Scheduling and Coordination; Path Planning for Multiple Mobile Robots or Agents; Multi-Robot Systems
I Introduction
In the field of multi-robot systems, two challenges are consistently at the forefront of research interest: (1) task allocation [1], determining which robot should perform which task, and (2) planning [2], the strategy for executing these tasks. Traditionally, these problems have been tackled separately by researchers from various fields. To simplify these issues, certain assumptions are often employed, such as the existence of low-level controllers for task allocation or the use of pre-defined tasks in motion planning. These approaches, however, address only portions of the whole problem. Over the past two decades, methods have emerged that address these issues by combining task allocation and planning. Our work aligns with this integrative approach, but distinguishes itself by focusing on tasks defined by temporal logic specifications.
Formal methods, characterized by their mathematical rigor, are essential for specifying, developing, analyzing, and verifying software and hardware systems [3]. The recent trend of utilizing formal specifications, such as Linear Temporal Logic (LTL) [4], as high-level task specifications for multi-robots has garnered significant attention. This is because temporal logic specifications can encapsulate not just conventional point-to-point Boolean goals but also complex temporal requirements. While offering an expressive framework for task descriptions, temporal logic specifications also introduce computational challenges for solvers due to their complexity. Take LTL for an example. A typical method involves translating the formula into an automaton, a graphic representation. As shown in [5], which involves a task of collecting 5 keys and then opening 5 doors, it took approximately half an hour on a laptop with an Intel i7 CPU and 32 GB of RAM to convert the corresponding LTL into an automaton consisting of 65 nodes and 792 edges. Similarly, in our simulations, an automaton could not be generated within an hour on a laptop with Apple M2 Pro and 16G RAM. This issue stems from the usage of the “flat” form of LTL, that is, putting all requirements on robot behaviors into a single LTL formula, which tends to become difficult to solve for complex tasks. However, one key observation is that robot tasks often have a loose connection and can be broken down into smaller components. Studies suggest that humans prefer hierarchical task specification, which improves interpretability of planning and execution, making it easier to identify ongoing work and conveniently adjust unfeasible parts without affecting other components [6, 7]. Inspired by this, we propose a hierarchical form of LTL specifications.
Example 1
Consider a mobile manipulation task where robots are tasked with first picking and placing item $a$ and then fetching items $b$ and $c$ in any order. The “flat” specification is
ϕ=italic-$\phi\,=\,$ $\Diamond(s_{a}\wedge\Diamond(t_{a}\wedge\Diamond(s_{b}\wedge\Diamond t_{b})\wedge\Diamond(s_{c}\wedge\Diamond t_{c}))),$
[…]
The exact formulas can be found in Eq. (2), with syntax and semantics elaborated in Sec. IV. This representation effectively reduces the length of each formula, leading to smaller automatons and improving interpretability. $\square$
In this work, we first establish syntax rules for hierarchical syntactically co-safe LTL (sc-LTL) and then provide a framework for semantic interpretation. To do this, considering the context of multiple specifications, we introduce the concept of state-specification plan, which pairs the state of each robot with a specific specification it seeks to fulfill at a given time. Utilizing this concept, we devise an algorithm to check the satisfaction of a set of hierarchical sc-LTL specifications given a state-specification plan. Theoretically, we prove that the hierarchical form is more expressive compared to the flat form. Practically, we conducted a user study to show that using a hierarchical structure makes it easier to express complex tasks with temporal constraints. Next, we propose a planning algorithm with formal guarantees of completeness and optimality under mild assumptions. The bottom-up approach involves searching within a product graph that combines the environment and task space. This method’s advantage is it can simultaneously achieve task allocation and planning (STAP).
Inspired by [8], we approximate the entire search space as a set of loosely interconnected sub-spaces, each corresponding to an individual sc-LTL specification in the hierarchical structure. The search process primarily occurs within a single sub-space, with transitions to adjacent sub-spaces happening under conditions determined by the decomposition of the automatons. Under certain assumptions regarding the state-specification plan, we demonstrate that our planning algorithm is sound, complete, and optimal. Additionally, we introduce several heuristics to expedite the search process, including minimizing the loose connections between sub-spaces and leveraging task-level progress to guide the search. We evaluate on extensive simulations on robot service tasks. Comparisons between our approaches with and without heuristics highlight a trade-off between computational time and solution quality. Our method shows significant improvements in computational time than work [8]. Notably, for complex tasks, [8] failed to produce solutions within a one-hour timeout. In scalability tests, our method demonstrated the ability to generate sub-optimal solutions in approximately three minutes for complex tasks involving up to 30 robots. Note that the flat version of these tasks could not be solved using graph-theoretic methods due to the inability to generate an automaton within a one-hour timeout. A glossary of key notations is provided in Tab. I.
Contributions The contributions are listed as follows:
  1. 1.
    We introduce a hierarchical form of LTL and prove it is more expressive than the flat form;
  2. 2.
    We conduct a user study to show that the hierarchical sc-LTL is easier to interpret compared to the flat sc-LTL.
  3. 3.
    We develop a search-based planning algorithm, achieving simultaneous task allocation and planning. This planner is, to the best of our knowledge, the first to offer both completeness and optimality for hierarchical sc-LTL;
  4. 4.
    We devise multiple heuristics to expedite the search;
  5. 5.
    Under mild assumptions, we theoretically analyze the completeness and optimality of our approach;
  6. 6.
    We conduct extensive comparative simulations focusing on service tasks to showcase the efficiency and scalability of our proposed method.
II Related Work
II-A Multi-robots Planning under Temporal Logic Specifications
In the existing body of research on optimal control synthesis from LTL specifications, two primary approaches are present in handling LTL tasks for multi-robot systems. One approach, as seen in works [9, 10, 11], involves assigning LTL tasks locally to individual robots within the team. The other approach assigns a global LTL specification to the entire team. In scenarios where global LTL specifications are used, these specifications can either explicitly assign tasks to individual robots [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], or the tasks may not be explicitly designated to specific robots [22, 23, 24, 25], similar to our problem in this work.
As temporal logic formulas are used to tackle tasks involving multiple robots and complex environments, they inevitably become lengthy. Some attempts have been made to simplify this, such as merging multiple atomic propositions into one using logical operators [18, 21], or combining multiple sub-formulas into one formula using logical operators [14, 13, 26]. Despite these efforts, we still categorize these as the flat form. Notably, temporal operators, the feature distinguishing LTL from propositional logic, only appear on one side, either for the integration of merged propositions or inside the sub-formulas. Works [23, 26, 27, 28, 29, 30] have defined propositions involving more than one robot to encapsulate collaborative tasks. In our study, we introduce composite propositions that encompass more than one sub-formula and can be combined using temporal operators. Besides, various extensions of temporal logic have been introduced: [26] developed Counting Temporal Logic (CTL) to express the collective behavior of multiple robots. [31, 32] utilized Graph Temporal Logic (GTL) and Swarm Signal Temporal Logic (SwarmSTL) respectively to specify swarm characteristics like centroid positioning and density distribution. [27] introduced Capability Temporal Logic (CaTL) to encapsulate the diverse capabilities of robots. However, all works mentioned focus on using flat LTL specifications to describe the behaviors in multi-robot systems.
Global specifications that do not explicitly assign tasks to robots generally require decomposition to derive the task allocation, which can be accomplished in three ways: The most common method, utilized in works such as [8, 20, 33, 34, 35, 28, 36], involves decomposing a global specification into multiple tasks, which leverages the transition relations within the automaton, which is the graphical representation of an LTL formula. As demonstrated in [23, 26], the second approach builds on Bounded Model Checking (BMC) methods [37] to create a Boolean Satisfaction or Integer Linear Programming (ILP) model, which simultaneously addresses task allocation and implicit task decomposition in a unified formulation. Another method, proposed by [38], directly interacts with the syntax tree of LTL formulas, which segments the global specification into smaller, more manageable sub-specifications.
The works most closely related to ours are [8, 39, 40, 41, 42]. Works [8, 39] approached STAP by breaking down temporal logic tasks into independent tasks, each of which can be completed by a robot. [40] built upon this idea, incorporating environmental uncertainties and failures into the planning. [41] expanded further into a multi-objective setting, integrating conflicting objectives such as cost and success probability into a Markov Decision Process (MDP). Our work presents key differences: (i) While the aforementioned studies focus on flat LTL specifications, we consider hierarchical specifications and can solve tasks which are unsolvable by these earlier methods. (ii) In the mentioned studies, tasks between robots are independent thus robot can execute their tasks in parallel. In contrast, our approach also assigns tasks with temporal dependencies to different robots, thus robots may depend on others. Note that [42], developed around the same time, is the only existing work that investigated decision-making under hierarchical sc-LTL. This work differs from ours in that, algorithmically, [42] proposed a hierarchical approach that first focused on task allocation and then planning, unlike STAP here. In terms of task decompoistion, [42] treated each transition in the automaton as an individual task to be allocated. In contrast, in this work individual tasks can consist of several transitions or even span the entire automaton, making each task potentially equivalent to multiple tasks as defined in [42]. Theoretically, the planning algorithm introduced in [42] lacks formal guarantees of either completeness or optimality.
II-B Task Allocation and Planning
The domain of Multi-Robot Task Allocation (MRTA) is thoroughly studied, and various classifications of MRTA have been developed [1, 43], organizing the extensive research in this field. In terms of MRTA, the problem is this work can be is characterised by four aspects: (i) single-task robots (each robot is limited to executing one task at a time), (ii) single-robot tasks (only one robot is required for each task), (iii) instantaneous assignment (tasks are allocated under all available information at this moment), and (iv) cross-schedule interdependence (the assignment of a robot to a task is influenced not solely by its own schedule but also by the schedules of other robots within the system).
Methods for MRTA primarily fall into two categories: market-based [44] and optimization-based approaches [45]. Market-based methods use economic principles, employing auction and bidding techniques to distribute tasks according to cost and resource. Of these, auction-based methods see robots acting selfishly, each calculating bids for tasks based on their self-interest [46], while consensus-based methods [47] combine an auction stage with a consensus stage, using this as a conflict resolution mechanism to establish agreement on the final bids. Moreover, MRTA problems can also be addressed using various optimization techniques, such as Integer Linear Programming (ILP) or Mixed Integer Linear Programming (MILP) [48]. Additionally, meta-heuristic algorithms like genetic algorithms [49] or particle swarm optimization [50] are employed for more intricate MRTA challenges where exact methods are unfeasible. Generally, research on MRTA assumes that the cost of a robot executing a task is known or known with some uncertainty, focusing less on the planning.
The domain of integrated task assignment and path planning in multi-robot systems involves both task-level reasoning and motion-level planning. The primary objective is to create collision-free paths for robots, enabling them to accomplish a variety of reach-avoid tasks pending assignment. This field is often associated with the unlabeled version of multi-agent path planning (MAPF), as discussed by [51]. Various formulations have been proposed, including those by [52, 53, 54, 55], to name a few. [53] focused on simultaneous task assignment and path planning for Multi-agent Pickup and Delivery (MAPD) in warehouse settings, where the goal is to efficiently manage agents transporting packages. [54] tackled the combined challenge of task allocation and path planning, where an operator must assign multiple tasks to each vehicle in a fleet, ensuring collision-free travel and minimizing total travel costs. [55] explored this issue for teams of agents, each assigned the same number of targets as there are agents in the team. Our work differs from these studies in that: (i) We do not assume a predefined set of point-to-point navigation tasks. Instead, tasks, which may include navigation and manipulation, are implicitly defined within the temporal logic specifications. (ii) The presence of logical and temporal constraints between tasks adds complexity to our problem. (iii) In our work, a single robot might be assigned multiple tasks, while others might remain stationary. This contrasts with scenarios where each robot is assigned exactly one task in works mentioned above.
II-C Hierarchical Task Models
Hierarchical reasoning enhances human understanding of the world [6, 7]. In classical AI planning, researchers crafted task models reflecting hierarchical structures by employing domain control knowledge [56]. These models have proven to be superior to flat models in terms of interpretability and efficiency, due to the significant reduction in the search space. Hierarchical Task Network (HTN) [57], a commonly used task model, exemplifies this. It presents a hierarchy of tasks, each of which can be executed if it’s primitive, or broken down into finer sub-tasks if it’s complex. Its plan comprises a set of primitive tasks applicable to the initial world state. Owing to its expressiveness, HTN has been implemented in the planning [58]. There are other hierarchical models such as AND/OR graphs [59] and sequential/parallel graphs [60].
Research combining hierarchical task models with LTL includes studies that use LTL to express temporally extended preferences over tasks and sub-tasks in HTN [61], as well as research into the expressive power of HTN in combination with LTL [62]. However, despite the widespread use of hierarchical task models in classical AI planning, it’s intriguing to note the lack of specification hierarchy in temporal logic robotic planning. Our work differs from these studies as we follow an inverse direction; instead of integrating LTL into HTN to express the goal of the planning problem, we incorporate HTN into LTL, allowing for hierarchical structures within multiple LTL formulas, making them more capable of expressing complex tasks than a single flat LTL formula.
Notation Description Notation Description
$\pi$ atomic propositions $\sigma$ infinite word over the alphabet
[…]
$\Diamond$ eventual operator ${\mathcal{Q}}_{\mathcal{B}}$ set of automaton states
$\square$ global operator ${\mathcal{Q}}_{\mathcal{B}}^{0}$ set of initial automaton states
$\mathcal{U}$ until operator $\Sigma=2^{{\mathcal{A}}{\mathcal{P}}}$ alphabet
$\phi$ specification $\to_{\mathcal{B}}$ transition relation in ${\mathcal{B}}$
${\mathcal{A}}{\mathcal{P}}$ set of atomic propositions ${\mathcal{Q}}_{\mathcal{B}}^{F}$ set of final automaton states
${\mathcal{T}}(r)$ transition system of robot $r$ $\rho_{\mathcal{B}}$ infinite run
${\mathcal{S}}_{r}$ states of robot $r$ ${\mathcal{L}}_{\mathcal{B}}$ accepted language of ${\mathcal{B}}$
$s_{r}^{0}$ initial state of robot $r$ ${\mathcal{P}}(r,\phi)$ PBA of robot $r$ and specification $\phi$
$\to_{r}$ transition relation of robot $r$ ${\mathcal{Q}}_{\mathcal{P}}$ set of product states
${\mathcal{A}}{\mathcal{P}}_{r}$ atomic proposition of robot $r$ ${\mathcal{Q}}_{\mathcal{P}}^{0}$ set of initial product states
${\mathcal{L}}_{r}$ observation function of robot $r$ $\to_{\mathcal{P}}$ transition relation in ${\mathcal{P}}$
$\phi_{k}^{i}$ $i$-th specification at the $k$-th level ${\mathcal{Q}}_{\mathcal{P}}^{F}$ set of final states
${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$ specification hierarchy graph $c_{\mathcal{P}}$ cost function in ${\mathcal{P}}$
$\Phi^{k}$ set of specifications at the $k$-th level $\tau=\tau^{0}\ldots\tau^{h}$ state-specification plan
$\Phi^{K}$ set of specifications at the lowest level ${\mathcal{D}}_{\mathcal{B}}(\phi)$ decomposition set of specification $\phi$
$\Phi^{acc}$ set of satisfied specifications up to certain time ${\mathcal{P}}(\phi)$ product team model associated with specification $\phi$
$\zeta_{\text{in}}$ in-spec switch transition ${\mathcal{P}}$ hierarchical team models
$\zeta_{\text{inter}}^{1}$ type I inter-spec switch transition $\zeta_{\text{inter}}^{2}$ type II inter-spec switch transition
$v=(r,\phi,\mathbf{s},\mathbf{q})$ search state $\Lambda=v_{0}\ldots v_{n}$ search path
$\Lambda^{\prime}=\Lambda_{1}\ldots\Lambda_{l}$ search path after removing all switch states ${\mathcal{R}}_{k}$ active robots of the $k$-th path segment
TABLE I: Summary of notations.
III Preliminaries
Notation: Let $\mathbb{N}$ denote the set of all integers, $[K]=\{1,\ldots,K\}$ denote the set of integers from 1 to $K$ and ||$\cdot$|| ⋅ | denote the cardinality of a set.
In this section, we formally describe Linear Temporal Logic (LTL) by presenting its syntax and semantics. Also, we briefly review preliminaries of automata-based LTL model checking.
[…]
$\phi::=\text{true}~{}|~{}\pi~{}|~{}\phi_{1}\wedge\phi_{2}~{}|~{}\neg\phi~{}|~{}\bigcirc\phi~{}|~{}\phi_{1}~{}\mathcal{U}~{}\phi_{2}.$ (1)
For brevity, we abstain from deriving other Boolean and temporal operators, e.g., disjunction $\vee$, implication $\Rightarrow$, always $\square$, eventually $\lozenge$, which can be found in [63].
[…]
Definition III.1 (NBA)
[…]
  • ${\mathcal{Q}}_{{\mathcal{B}}}$ is the set of states;
  • ${\mathcal{Q}}_{{\mathcal{B}}}^{0}\subseteq{\mathcal{Q}}_{{\mathcal{B}}}$ is a set of initial states;
  • $\Sigma=2^{\mathcal{AP}}$ is an alphabet;
  • $\rightarrow_{{\mathcal{B}}}\,\subseteq\,{\mathcal{Q}}_{{\mathcal{B}}}\times\Sigma\times{\mathcal{Q}}_{{\mathcal{B}}}$ is the transition relation;
  • ${\mathcal{Q}}_{{\mathcal{B}}}^{F}\subseteq{\mathcal{Q}}_{{\mathcal{B}}}$ is a set of accepting/final states.
An infinite run $\rho_{{\mathcal{B}}}$ of ${\mathcal{B}}$ over an infinite word $\sigma=\pi_{0}\pi_{1}\pi_{2}\dots$, $\pi_{k}\in\Sigma=2^{\mathcal{AP}}$, $\forall k\in\mathbb{N}$, is a sequence $\rho_{{\mathcal{B}}}=q_{{\mathcal{B}}}^{0}q_{{\mathcal{B}}}^{1}q_{{\mathcal{B}}}^{2}\dots$ such that $q_{{\mathcal{B}}}^{0}\in{\mathcal{Q}}_{{\mathcal{B}}}^{0}$ and $(q_{{\mathcal{B}}}^{k},\pi_{k},q_{{\mathcal{B}}}^{k+1})\in\rightarrow_{{\mathcal{B}}}$, $\forall k\in\mathbb{N}$. An infinite run $\rho_{{\mathcal{B}}}$ is called accepting if $\texttt{Inf}(\rho_{{\mathcal{B}}})\cap{\mathcal{Q}}_{{\mathcal{B}}}^{F}\neq\varnothing$, where $\texttt{Inf}(\rho_{{\mathcal{B}}})$ represents the set of states that appear in $\rho_{{\mathcal{B}}}$ infinitely often. The words $\sigma$ that produce an accepting run of ${\mathcal{B}}$ constitute the accepted language of ${\mathcal{B}}$, denoted by ${\mathcal{L}}_{{\mathcal{B}}}$. Then [63] proves that the accepted language of ${\mathcal{B}}$ is equivalent to the words of $\phi$, i.e., ${\mathcal{L}}_{{\mathcal{B}}}=\texttt{Words}(\phi)$.
The dynamics of robot $r$ is captured by a Transition System (TS) defined as follows:
Definition III.2 (TS)
A Transition System for robot $r$ is a tuple ${\mathcal{T}}(r)=\{{\mathcal{S}}_{r},s_{r}^{0},\to_{r},{\mathcal{A}}{\mathcal{P}}_{r},{\mathcal{L}}_{r}\}$ where:
  • ${\mathcal{S}}_{r}=\{s_{r}\}$ is the set of discrete states of robot $r$;
  • $s_{r}^{0}$ is the initial state of robot $r$;
  • $\to_{r}\,\subseteq\,{\mathcal{S}}_{r}\times{\mathcal{S}}_{r}$ is the transition relation;
  • $\mathcal{AP}_{r}$ is the set of atomic propositions related to robot $r$;
  • ${\mathcal{L}}_{r}:{\mathcal{S}}_{r}\rightarrow 2^{\mathcal{AP}_{r}}$ is the observation (labeling) function that returns a subset of atomic propositions that are satisfied, i.e., ${\mathcal{L}}_{r}(s_{r})\subseteq\mathcal{AP}_{r}$.
[…]
Definition III.3 (PBA)
[…]
  • $\mathcal{Q}_{\mathcal{P}}={\mathcal{S}}_{r}\times\mathcal{Q}_{{\mathcal{B}}}$ is the set of product states;
  • $\mathcal{Q}_{\mathcal{P}}^{0}=\{s_{r}^{0}\}\times\mathcal{Q}_{\mathcal{B}}^{0}$ is a set of initial states;
  • $\longrightarrow_{{\mathcal{P}}}\subseteq\mathcal{Q}_{\mathcal{P}}\times\mathcal{Q}_{\mathcal{P}}$ is the transition relation defined by the rule: $\frac{(s_{r}\rightarrow_{r}s^{\prime}_{r})\wedge(q\xrightarrow{{\mathcal{L}}_{r}\left(s_{r}\right)}_{{\mathcal{B}}}q^{\prime})}{q_{{\mathcal{P}}}=\left(s_{r},q\right)\longrightarrow_{\mathcal{P}}q_{{\mathcal{P}}}^{\prime}=\left(s^{\prime}_{r},q^{\prime}\right)}$. The transition from the state $q_{\mathcal{P}}\in\mathcal{Q}_{\mathcal{P}}$ to $q_{\mathcal{P}}^{\prime}\in\mathcal{Q}_{\mathcal{P}}$, is denoted by $(q_{\mathcal{P}},q_{\mathcal{P}}^{\prime})\in\longrightarrow_{\mathcal{P}}$, or $q_{\mathcal{P}}\longrightarrow_{\mathcal{P}}q_{\mathcal{P}}^{\prime}$;
  • $\mathcal{Q}_{\mathcal{P}}^{F}={\mathcal{S}}_{r}\times\mathcal{Q}_{\mathcal{B}}^{F}$ is a set of accepting/final states;
IV Hierarchical Syntactically Co-safe LTL
In this work, we focus on a particular subset of LTL formulas known as syntactically co-safe LTL, or sc-LTL for short [65]. As established by [65], any LTL formula encompassing only the temporal operators $\Diamond$ and $\mathcal{U}$ and written in positive normal form (where negation is exclusively before atomic propositions) is classified under syntactically co-safe formulas. This category does not include the $\square$ operator. Sc-LTL formulas can be satisfied by finite sequences followed by any infinite repetitions. This characteristic makes sc-LTL apt for modeling and reasoning about systems with finite durations, such as those found in the robotics field. We first present the syntax of hierarchical sc-LTL, then we propose the interpretation of semantics of hierarchical sc-LTL specifications, along with the analysis of the expressiveness.
IV-A Syntax of Hierarchical sc-LTL
Definition IV.1 (Hierarchical sc-LTL)
Hierarchical sc-LTL is structured into $K$ levels, labeled as $L_{1},\ldots,L_{K}$, arranged from the highest to the lowest. Each level $L_{k}$, where $k\in[K]$, contains $n_{k}$ sc-LTL formulas. The hierarchical sc-LTL can be represented as $\Phi=\left\{\phi_{k}^{i}\,|\,k\in[K],i\in[n_{k}]\right\}$, where $\phi_{k}^{i}$ denotes the $i$-th sc-LTL formula at level $L_{k}$. The hierarchical sc-LTL adheres to the following rules:
  1. 1.
    Each formula at a given level $L_{k}$, for $k\in[K-1]$, is derived from the formulas at the next lower level $L_{k+1}$.
  2. 2.
    Every formula at any level other than the highest (i.e., $k=2,\ldots,K$) is included in exactly one formula at the next higher level $L_{k-1}$.
  3. 3.
    Atomic propositions are used exclusively within the formulas at the lowest level $L_{K}$.
Let $\Phi^{k}$ denote the set of formulas at level $L_{k}$ with $k\in[K]$. For any formula $\phi_{k}^{i}$ at the non-bottom level $k<K$, we slightly bend the notation to use $\phi_{k}^{i}$ to represent the same symbol at the higher level $L_{k-1}$, which we refer to as composite proposition. In this manner, $\phi_{k}^{i}$ represents not only the $i$-th formula at level $L_{k}$, but also the corresponding composite proposition at level $L_{k-1}$. In this work, when $\phi_{k}^{i}$ appears at the right side in a certain formula, we consider it as a composite proposition. On the other hand, when it appears at the left side as a standalone formula, we refer to it as a specification. The direct result of condition 3) is that non-bottom specifications consist solely of composite propositions.
Example 1
continued (Hierarchical sc-LTL) The hierarchical specifications for the task of first picking and placing item $a$ and then fetching items $b$ and $c$ in any order are:
[…]
where there are two levels, $L_{1}$ and $L_{2}$, with $L_{1}$ having one specification and $L_{2}$ having three specifications. Their automatons have a total of 8 states and 18 transitions. The symbol $\phi_{2}^{1}$ is a specification at level $L_{2}$ but is a composite proposition at level $L_{1}$. $\Phi^{1}=\{\phi_{1}^{1}\}$ and $\Phi^{2}=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}$. $\square$
As discussed in Section II-A, local and global forms of task specifications are distinguished based on whether robots receive tasks from individual specifications or from a unified specification. We highlight that the direction distinguishing local from global forms is complementary to the direction distinguishing flat from hierarchical forms. Thus, the hierarchical form can be integrated with both local and global forms. Here, we employ the global form of hierarchical sc-LTL.
Definition IV.2 (Specification Hierarchy Graph)
The specification hierarchy graph, denoted as ${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$, is a tree where each node represents a specification within the hierarchical sc-LTL, and an edge $(u,v)$ indicates that specification $u$ contains specification $v$ as a composite proposition.
(a) Specification hierarchy graph
(b) NBA of sc-LTL $\Diamond a\wedge\Diamond b$
Figure 1: The specification hierarchy graph corresponds to hierarchical sc-LTL specifications in (2). The NBA corresponds to the sc-LTL specification in Example 2, where non-trivial decomposition states are colored in gray.
Definition IV.3 (Leaf and Non-leaf Specifications)
A specification is termed as a leaf specification if the associated node in the graph ${\mathcal{G}}_{h}$ does not have any children; otherwise, it is referred to as a non-leaf specification.
Lemma IV.4
Any hierarchical sc-LTL specifications can be turned into a specification hierarchy graph.
Example 1
continued (Specification hierarchy graph)  The specification hierarchy graph is illustrated in Fig. 1(a). $\square$
IV-B Semantics of Hierarchical sc-LTL
Hierarchical sc-LTL cannot be interpreted by simply substituting composite propositions with their corresponding specifications. For example, replacing composite propositions in Eq. (2) at level $L_{1}$ generates a flat specification:
$\phi_{1}^{1}=\Diamond(\underbrace{\Diamond(s_{a}\wedge\Diamond t_{a})}_{\phi_{2}^{1}}\wedge\Diamond(\underbrace{\Diamond(s_{b}\wedge\Diamond t_{b})}_{\phi_{2}^{2}})\wedge\Diamond(\underbrace{\Diamond(s_{c}\wedge\Diamond t_{c})}_{\phi_{2}^{3}})),$
which is challenging to interpret.
Example 1
continued (Interpretation)  The way of interpretation is to start from the bottom level and progress upwards, verifying if the satisfaction of lower specifications allows the upper ones to move forward. We provide two separate cases. Case (i): If $s_{a}$ now holds true, then $\phi_{2}^{1}$ in (2) transitions to $\Diamond t_{a}$ based on the rules of progression [66] that generate sub-formulas to be satisfied, while $\phi_{1}^{1}$ remains unaffected as none of its composite propositions are fulfilled. If $t_{a}$ subsequently holds true, $\phi_{2}^{1}$ is satisfied. Consequently, $\phi_{1}^{1}$ then simplifies to $\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$. Following this logic, $\phi_{1}^{1}$ will be fulfilled when both $\phi_{2}^{2}$ and $\phi_{2}^{3}$ are satisfied, regardless of the order. Case (ii): Assuming the robot is capable of carrying multiple items. If $s_{b}$ is satisfied, then $\phi_{2}^{2}$ progresses to $\Diamond t_{b}$. Subsequently, as $s_{a}$ and $t_{a}$ sequentially become true, $\phi_{1}^{1}$ transitions to $\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$. At this stage, $\phi_{2}^{2}$ only requires $t_{b}$ to be satisfied. We allow the partial satisfaction of $\phi_{2}^{2}$ provided that it achieves full satisfaction no earlier than $\phi_{2}^{1}$. $\square$
Input: Hierarchical sc-LTL specifications $\Phi$, state-specification plan $\tau$
Output: True or False
1 $\Phi^{acc}=\varnothing$ ;
2 for $i\in\{0,\ldots,h\}$  do
3       for $k\in\{K,\ldots,1\}$  do
4             $\Phi^{acc,k}=\varnothing$ ;
5             for $\phi\in\Phi^{k}\setminus\Phi^{acc}$  do
6                   if $k=K$  then
7                         $S(i)=\left\{s_{r}^{i}\,|\,(s_{r}^{i},\phi)\in\tau^{i}\right\}$ ;
8                         if $S(i)=\varnothing$ then
9                               continue;
10
11                         $\pi^{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\,|\,s_{r}^{i}\in S(i)\right\}$ ;
12
13                  else
14                         $\pi^{i}=\Phi^{acc,k+1}$ ;
15
16                   ${\mathcal{Q}}_{\mathcal{B}}^{i+1}(\phi)=\left\{q^{\prime}\,|\,(q,\pi^{i},q^{\prime})\in\to_{\mathcal{B}}(\phi),\,\forall q\in{\mathcal{Q}}_{\mathcal{B}}^{i}(\phi),\right.$ $\left.\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{B}}(\phi)\right\}$ ;
17                   if ${\mathcal{Q}}^{i+1}_{\mathcal{B}}(\phi)\cap{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi)\neq\varnothing$  then
18                         $\Phi^{acc,k}=\Phi^{acc,k}\cup\{\phi\}$ ;
19
20            $\Phi^{acc}=\Phi^{acc}\cup\Phi^{acc,k}\,\cup\,\texttt{Children}(\Phi^{acc,k})$ ;
21
22      if $\phi_{1}^{1}\in\Phi^{acc}$ then
23             return True ;
24
25return False ;
Algorithm 1 Satisfaction check
Given the presence of multiple specifications in hierarchical sc-LTL, a robot could be addressing any specification at a specific state. To clarify the robot’s intention and facilitate the checking of satisfaction, we associate the state $s_{r}$ of a robot $r$ with a leaf specification $\phi_{r}\in\Phi^{K}$. This association forms a state-specification pair $(s_{r},\phi_{r})$, signifying that the robot $r$ is engaged in fulfilling the specification $\phi_{r}$ at state $s_{r}$.
Definition IV.5 (State-Specification Plan)
A state-specification plan with a horizon $h$, represented as $\tau$, is a timed sequence $\tau=\tau^{0}\tau^{1}\tau^{2}\ldots\tau^{h}$. Here, $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ is the collective state-specification pairs of $N$ robots at the $i$-th timestep, where $s_{r}^{i}\in{\mathcal{S}}_{r}$, and $\phi_{r}^{i}\in\Phi^{K}\cup\{\epsilon\}$, with $\epsilon$ indicating the robot’s non-involvement in any leaf specification at that time.
When considering the composite propositions in the context of long-horizon robot tasks, they correspond to sub-tasks. For example, one such task could be first clearing the table and then disposing of the trash, with each sub-task required to be completed exactly once. Another example involves retrieving a beverage from the refrigerator, such as a bottle of apple juice or orange juice, where completing just one sub-task suffices. Given this observation, we propose the following assumption:
Assumption IV.6 (Composite proposition)
For the successful completion of the robotic task, it is adequate that each composite proposition is satisfied no more than once.
Given a robot state-specification plan, the method for determining the satisfaction of hierarchical sc-LTL specifications is shown in Alg. 1. Because the truth of a non-leaf specification depends solely on the specifications at the immediately lower level, the algorithm iteratively assesses specifications in a bottom-up manner. Alg. 1 will stop early and return true if the highest-level specification is met [lines 1-1]. This is because that all specifications belong to co-safe LTL, the subsequent plan becomes irrelevant once the topmost specification is satisfied, eliminating the need for further satisfaction checks. Let $\Phi^{acc}$ represent the set of specifications that have been satisfied up to the current point [line 1]. Additionally, let $\Phi^{acc,k}$ denote the set of specifications at level $L_{k}$ that have just become satisfied at the current time step $i$ [line 1]. According to Assumption IV.6, each composite proposition is adequate to be satisfied no more than once, therefore, we iterate over each specification $\phi\in\Phi^{k}\setminus\Phi^{acc}$ at level $L_{k}$ that remain unsatisfied. Alg. 1 computes the reachable set of automaton states for the next time step, designated as ${\mathcal{Q}}^{i+1}_{\mathcal{B}}(\phi)$, to evaluate the satisfaction [lines 1-1]. This assessment varies depending on whether the specification $\phi$ is a leaf one or not. If the specification is a leaf specification, the algorithm aggregates the set of true propositions $\pi^{i}$, which are atomic propositions derived from the states of robots engaged in the specification $\phi$ [lines 1-1]. Otherwise, if the specification is not a leaf, the set of true propositions is identified as the set $\Phi^{acc,k+1}$ – these are the specifications at the immediate lower level that have just been satisfied [line 1]. The set $\Phi^{acc,k}$ gets updated if the reachable automaton states include at least one accepting state of $\phi$ [lines 1]. Following the completion of iterations over level $L_{k}$, the set $\Phi^{acc}$ is then updated [line 1], where $\texttt{Children}(\Phi^{acc,k})$ denotes the set of specifications that are children nodes in the specification hierarchy graph of any specification in $\Phi^{acc,k}$, that is, the set of specifications for which a path exists from any specification in $\Phi^{acc,k}$. There is no need to verify the satisfaction of a specification if its parent specification has already been satisfied. For instance, there is no need to verify the satisfaction of $\phi_{2}^{1}$ for the specification $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2})$ if $\phi_{2}^{2}$ becomes true.
Remark IV.7
Every sc-LTL formula can be transformed into a Deterministic Finite Automaton (DFA). We opt for NBA, a form of Nondeterministic Finite Automaton (NFA), because sc-LTLs are transformable into either DFAs or NFAs, making both equally expressive for sc-LTL in this sense. The work from [8] and the temporal order heuristic by [21] that our approach in Sec. VI is based on, uses NFAs. Nonetheless, our approach can easily be adapted to work with DFAs as well, as graphically speaking, NFAs are more complex than DFAs.
Remark IV.8
The syntax and semantics for hierarchical sc-LTL here are designed for robotics and may not be applicable to other domains. Our experience indicates that a top-down reasoning approach is effective in constructing hierarchical sc-LTL specifications from task descriptions. This entails recognizing the inclusion relationships among tasks and organizing them into various levels of abstraction, ranging from general to specific. At each level, tasks that are closely related should be combined into single formulas, while those with weaker connections should be grouped into separate formulas.
IV-C Expressiveness
Temporal logics are interpreted over paths in finite Kripke structures. The Kripke structure is identical to the transition system as presented in Def. III.2 in this work. A Kripke structure, denoted by ${\mathcal{K}}$, is deemed finite when it has a finite number of states. ${\mathcal{K}}$ is considered a model for a temporal logic formula $\phi$ if every initial infinite path within ${\mathcal{K}}$ satisfies $\phi$.
Definition IV.9 (Expressive power [67])
Considering two logics ${\mathcal{L}}_{1}$ and ${\mathcal{L}}_{2}$, along with two formulas $\phi_{1}\in{\mathcal{L}}_{1}$ and $\phi_{2}\in{\mathcal{L}}_{2}$, the formula $\phi_{1}$ is equivalent to $\phi_{2}$ if, for any finite Kripke structure ${\mathcal{K}}$, ${\mathcal{K}}$ models $\phi_{1}$ if and only if it models $\phi_{2}$. The logic ${\mathcal{L}}_{2}$ is subsumed by ${\mathcal{L}}_{1}$, denoted as ${\mathcal{L}}_{1}\geq{\mathcal{L}}_{2}$, when for every formula $\phi_{2}\in{\mathcal{L}}_{2}$, there is a corresponding formula $\phi_{1}\in{\mathcal{L}}_{1}$ that is equivalent to $\phi_{2}$. Additionally, ${\mathcal{L}}_{1}$ is (strictly) more expressive than ${\mathcal{L}}_{2}$ if ${\mathcal{L}}_{1}\geq{\mathcal{L}}_{2}$ and ${\mathcal{L}}_{2}\not\geq{\mathcal{L}}_{1}$.
Theorem IV.10 (Expressiveness)
Hierarchical sc-LTL is more expressive than the flat sc-LTL.
Proof:
We first show that, for any flat sc-LTL, there exists an equivalent hierarchical counterpart. Therefore, the flat sc-LTL is subsumed by the hierarchical sc-LTL. Given a flat sc-LTL $\phi$, an equivalent two-level hierarchical sc-LTL can be constructed in a straightforward manner, where the top level is $\Diamond\phi$ and the bottom level is $\phi$ itself, that is,
$L_{1}:\phi_{1}^{1}=\Diamond\phi_{2}^{1},\quad L_{2}:\phi_{2}^{1}=\phi.$
Next, we introduce a counter-example to demonstrate the absence of the flat equivalent for certain hierarchical sc-LTL specifications. Considering a Kripke structure with two atomic propositions $\{a,b\}$, the following hierarchical sc-LTL specifications include two mutually exclusive leaf specifications:
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\,\Diamond\,\phi_{2}^{2}$
$L_{2}:\quad$ $\phi_{2}^{1}=\neg b\,\mathcal{U}\,(a\wedge\neg b)$
$\phi_{2}^{2}=\neg a\,\mathcal{U}\,(b\wedge\neg a).$
While leaf specifications are mutually exclusive, the introduction of the state-specification plan enables the association of each state with one specification, facilitating the separate treatment of leaf specifications. For example, when these two specifications are assigned to separate robots, they can be realistically fulfilled. However, when solely using atomic propositions $a$ and $b$, combining two leaf specifications into a single sc-LTL specification is unfeasible, as it inevitably results in a contradiction, reducing the specification to False. ∎
Remark IV.11
Hierarchical sc-LTL retains the expressive capability of sc-LTL for defining robot tasks, while also enhancing the ability to specify tasks for multiple robots. Moreover, as detailed in Section VIII-A, we conducted a user study that showed hierarchical sc-LTL is more user-friendly than flat sc-LTL.
V Problem Formulation
We first introduce a variant of the product team model defined in work [8], which aggregates the set of PBA for a given specification. Following this, we develop the hierarchical team model that integrates a set of product team models to address hierarchical sc-LTL specifications.
V-A Product Team Model
Inspired by work [8], we make the following assumption.
Assumption V.1 (Decomposition)
Every leaf specification $\phi\in\Phi^{K}$ can be decomposed into a set of tasks such that
  • Independence: every task can be executed independently;
  • Completeness: completion of all tasks, irrespective of the order, satisfies the leaf specification $\phi$.
Definition V.2 (Decomposition Set [8])
The decomposition set ${\mathcal{D}}_{\mathcal{B}}\subseteq{\mathcal{Q}}_{\mathcal{B}}$ of the NBA ${\mathcal{B}}$ contains all states $q$ such that each state leads to a decomposition of tasks.
Example 2 (Decomposition Set)
Consider an sc-LTL specification $\phi=\Diamond a\wedge\Diamond b$, with its corresponding NBA depicted in Fig. 1(b). For any two-part satisfying word $\sigma=\sigma_{1}\sigma_{2}$, where $\sigma_{1}$ leads to a run ending at state 2, the reverse sequence $\sigma_{2}\sigma_{1}$ also satisfies the specification by leading to a run through state 1. This indicates that the tasks before and after state 2 can be completed in any order. Therefore, state 2 belongs to the decomposition set. Similarly, state 1 also belongs to the decomposition set. By default, all initial and accepting states in an NBA are considered decomposition states. $\square$
Figure 2: The product team model featuring two robots. PBAs ${\mathcal{P}}(1,\phi)$ and ${\mathcal{P}}(2,\phi)$ may be different reflecting the robots’ heterogeneity. Initial states are indicated with incoming arrows, and accepting states are depicted with double circles. The decomposition sets are highlighted in gray. In-spec switch transitions are represented by dashed arrows. A plan exclusive to robot 1 is traced by the green path, whereas the red path delineates a plan that involves the collaboration of both robots.
Definition V.3 (Product Team Model (modified from [8]))
Given a leaf specification $\phi\in\Phi^{K}$, the product team model ${\mathcal{P}}(\phi)$ consists of $N$ product models ${\mathcal{P}}(r,\phi)$ and is given by the tuple ${\mathcal{P}}(\phi):=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,s,q):r\in[N],(s,q)\in{\mathcal{Q}}_{\mathcal{P}}(r,\phi)\right\}$ is the set of state;
  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:r=1,(s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(1,\phi)\right\}$ is the set of initial states ;
  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation. One type of transition, referred to as in-spec transitions, occurs inside a PBA. That is, $(r,s,q)\to_{\mathcal{P}}(r,s^{\prime},q^{\prime})$ if $(s,q)\to_{{\mathcal{P}}(r,\phi)}(s^{\prime},q^{\prime})$; the other type of transition relation, denoted by $\zeta_{\text{in}}$, occurs between two PBAs and is defined below in Def. V.4;
  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)\}$ is the set of accepting/final states ;
Definition V.4 (In-Spec Switch Transition (modified from [8]))
Given a specification $\phi$, the set of in-spec switch transitions in ${\mathcal{P}}(\phi)$ is given by $\zeta_{\text{in}}\subset{\mathcal{Q}}_{\mathcal{P}}\times{\mathcal{Q}}_{\mathcal{P}}$. A transition $\left((r,s,q),(r^{\prime},s^{\prime},q^{\prime})\right)$ belongs to $\zeta_{\text{in}}$ if and only if it:
  • points to the next robot, $r^{\prime}=r+1$;
  • preserves the NBA progress, $q=q^{\prime}$;
  • represents a decomposition choice, $q\in{\mathcal{D}}_{\mathcal{B}}(\phi)$.
The distinction between in-spec transitions and in-spec switch transitions lies in their impact on task progress: in-spec transitions advance the task through the same robot, while in-spec switch transitions involve changing to a different robot without altering the task’s progress, as indicated by $q=q^{\prime}$. A visual representation of ${\mathcal{P}}(r,\phi)$ can be found in Fig. 2. The connections between product models are unidirectional, starting from the first robot and ending at the last robot. The specific arrangement of robots is irrelevant. Transitioning to the next product model occurs solely at in-spec switch transitions. Any path connecting an initial to an accepting state constitutes a feasible solution, comprising multiple path segments. Both the start and end state of each path segment have automaton states that belong to the decomposition set. Each segment results in a sequence of actions for an individual robot. These action sequences can be executed in parallel as they collectively decompose the task, as illustrated in Fig. 2.
Remark V.5
Note that we construct team models exclusively for leaf specifications which directly involve atomic propositions. The switch transitions in [8] point to the initial state of the next robot, that is, $s^{\prime}=s_{r^{\prime}}^{0}$, which is not the case for the in-spec switch transitions here. This difference arises because [8] addresses a single specification, where a robot typically starts a task from its initial state. In contrast, in scenarios involving multiple specifications, a robot can begin a new task from its last state at the end of the preceding task.
Figure 3: Hierarchical team models for two leaf specifications and two robots. The product team models for each leaf specification are depicted within areas shaded in light orange and green. To enhance clarity, some transitions within a single PBA are omitted. Inter-spec type I and II switch transitions are illustrated using blue and red dashed arrows, respectively. It’s important to note that type I switch transitions are bidirectional, while type II switch transitions are unidirectional. For the sake of readability, inter-spec type I switch transitions are specifically shown for ${\mathcal{P}}(1,\phi_{1})$ and ${\mathcal{P}}(1,\phi_{2})$. Inter-spec type II switch transitions are only shown from ${\mathcal{P}}(2,\phi_{1})$ to ${\mathcal{P}}(1,\phi_{2})$. The path marked in yellow highlights a specific plan. This plan originates from the initial state of ${\mathcal{P}}(1,\phi_{1})$ and concludes at the accepting state of ${\mathcal{P}}(2,\phi_{2})$. This path involves assigning robot 2 to fulfill the specification $\phi_{1}$, which includes one in-spec switch transition ①. Subsequently, robots 1 and 2 are tasked with satisfying specification $\phi_{2}$, connected through a type II inter-spec switch transition ② and involving another in-spec switch transition ③.
V-B Hierarchical Team Models
To extend to hierarchical sc-LTL, we construct a product team model per leaf specification. The goal here is to connect them to form a connectable space. We refer to the scenario of decomposition of tasks for one specifications as in-spec independence. In contrast, leaf specifications can exhibit dependencies. For example, one leaf specification $\phi$ might need to be completed before another $\phi^{\prime}$. This scenario is referred to as inter-spec dependence. Another scenario is inter-spec independence, where two leaf specifications can be satisfied independently. We note that in-spec dependence does not exist, as precluded by Asm. V.1. To capture inter-spec dependency and independence, we introduce two additional types of switch transitions between different product team models.
Definition V.6 (Hierarchical Team Models)
Given the hierarchical sc-LTL specifications $\Phi$, the hierarchical team models ${\mathcal{P}}$ consists of a set of product team models ${\mathcal{P}}(\phi)$ with $\phi\in\Phi^{K}$ and is denoted by the tuple ${\mathcal{P}}:=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,\phi,s,q):\phi\in\Phi^{K},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}(\phi)\right\}$ is the set of state;
  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:\phi\in\Phi^{K},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(\phi)\right\}$ is the set of initial states ;
  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation. One type of transition occurs inside a product team model, that is, $(r,\phi,s,q)\to_{\mathcal{P}}(r^{\prime},\phi,s^{\prime},q^{\prime})$ if $(r,s,q)\to_{{\mathcal{P}}(\phi)}(r^{\prime},s^{\prime},q^{\prime})$; another two types of transition relation, denoted by $\zeta_{\text{inter}}^{1}$ and $\zeta_{\text{inter}}^{2}$, occur between two product teams models and are defined below in Def. V.7 and V.8;
  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)\}$ is the set of accepting/final states.
Definition V.7 (Inter-Spec Type I Switch Transition)
A transition, denoted by $\zeta_{\text{inter}}^{1}=\left((r,\phi,s,q),(r^{\prime},\phi^{\prime},s^{\prime},q^{\prime})\right)$, is type I switch transition if and only if it:
  • connects the same robots, $r=r^{\prime}$;
  • points to two different leaf specifications, $\phi\neq\phi^{\prime}$;
  • points to the same robot state, $s=s^{\prime}$;
  • connects decomposition states, $q\in{\mathcal{D}}_{\mathcal{B}}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})$.
Definition V.8 (Inter-Spec Type II Switch Transition)
A transition, denoted by $\zeta_{\text{inter}}^{2}=\left((r,\phi,s,q),(r^{\prime},\phi^{\prime},s^{\prime},q^{\prime})\right)$, is type II switch transition if and only if it:
  • points to the first robot, $r^{\prime}=1$;
  • points to two different leaf specifications, $\phi\neq\phi^{\prime}$;
  • connects an accepting state to a decomposition state, $q\in{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})$.
Inter-spec type I switch transitions enable switching to another PBA with a different leaf specification, while retaining the same robot and pausing the current task. Conversely, inter-spec type II switch transitions create a connection from the accepting state of one PBA to the decomposition state of another PBA associated with the first robot, indicating the completion of one specification and the resumption of another. These hierarchical team models can be graphically represented in a 3D space, with each product team model occupying a 2D plane; see Fig. 3. This structure has three axes: one representing the task progress for a specification (the $(s,q)$-axis), another indicating the task assignment among robots (the $r$-axis), and the last one tracking the overall progress of tasks across different specifications (the $\phi$-axis).
Remark V.9
The construction is not a product of all NBA and robot states. Instead, if each product team model is considered a distinct sub-space, the inter-spec switch transitions facilitate the connections between these sub-spaces. As a result, two consecutive product models are only loosely interconnected. The majority of states across different sub-spaces remain disconnected. Such a design significantly reduces the overall search space, enhancing the efficiency of the search.
Given a state-specification plan $\tau=\tau^{0}\tau^{1}\ldots\tau^{h}$ where $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$, the cost for robot $r$, such as energy consumption or completion time, is represented as $c_{r}=\sum_{i=0}^{h-1}c_{r}(s^{i}_{r},s^{i+1}_{r})$, where $c_{r}(s^{i}_{r},s^{i+1}_{r})$ denotes the cost incurred transitioning between states for robot $r$. The goal is to minimize the additive cost, expressed as
$J(\tau)=\sum_{r=1}^{N}c_{r}$ (3)
Finally, the problem can be formulated as follow:
Problem 1
Given the hierarchical team models ${\mathcal{P}}$, derived from transition systems of robots and the NBAs of hierarchical sc-LTL specifications $\Phi$, find an optimal state-specification plan $\tau^{*}$ that minimizes $J(\tau^{*})$ and satisfies $\Phi$.
VI Planning under Hierarchical sc-LTL
This section begins by introducing a search-based approach, which constructs hierarchical team models on the fly, ensuring both completeness and optimality; see Sec. VII. Subsequently, we discuss several heuristics designed to expedite the search.
VI-A On-the-fly Search
The on-the-fly search algorithm is built upon the Dijkstra’s algorithm, as detailed in Alg. 2. The search state is defined as $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, where ${\mathbf{s}}\in{\mathcal{S}}_{1}\otimes\ldots\otimes{\mathcal{S}}_{N}$ and ${\mathbf{q}}={\mathcal{Q}}_{\mathcal{B}}(\phi_{K}^{1})\otimes\ldots\otimes{\mathcal{Q}}_{\mathcal{B}}(\phi_{1}^{1})$. This state is comprised of three parts: (i) $(r,\phi)$ indicates the specific PBA currently being searched; (ii) ${\mathbf{s}}$ represents the states of all robots; and (iii) ${\mathbf{q}}$ denotes the automaton states of all specifications (including leaf and non-leaf ones). Two states are considered equivalent if they share the same $r$, $\phi$, ${\mathbf{s}}$, and ${\mathbf{q}}$ values. The notation ${\mathbf{s}}(r)$ and ${\mathbf{q}}(\phi_{k}^{i})$ refer to the state of robot $r$ and the automaton state of specification $\phi_{k}^{i}$ reached so far, respectively. We use ${\mathcal{V}}_{\text{explored}}$ and ${\mathcal{V}}_{\text{seen}}$ to denote the sets of states that have been explored and seen, and ${\mathcal{H}}_{\text{front}}$ for the states awaiting exploration. States within ${\mathcal{H}}_{\text{front}}$ are ordered by their costs. Lastly, $\Lambda_{\text{explored}}$ maps each explored state to the shortest path that led to it.
Definition VI.1 (Initial State)
A state $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$ is an initial state if
  • $(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N],\forall\phi\in\Phi^{K}$;
  • ${\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi^{K}$.
That is, all robots begin at their initial states, and every specification starts from its initial automaton state [line 2]. Whenever a state $v$ is popped out from the set ${\mathcal{H}}_{\text{front}}$ for exploration, it is added to the set of explored states if it hasn’t been already [lines 2-2]. The iteration process concludes when an accepting state of the topmost-level specification is reached. At this point, the function ExtractPlan is employed to parallelize the sequentially searched path, generating a state-specification plan [line 2]. If the accepting state is not yet reached, the algorithm proceeds to determine the successor states of $v$, along with their associated transition costs [line 2], through the function GetSucc, outlined in Alg. 3. For each successor state that is either unseen or has a total cost lower than previously found, it is queued into ${\mathcal{H}}_{\text{front}}$ for future exploration. Concurrently, the corresponding path leading to this state is recorded [lines 2-2].
Input: Hierarchical sc-LTL specifications $\Phi$, a set of robots $[N]$
Output: state-specification plan
1 ${\mathcal{V}}_{\text{seen}}=\varnothing,{\mathcal{V}}_{\text{explored}}=\varnothing,{\mathcal{H}}_{\text{front}}=\varnothing,\Lambda_{\text{explored}}=\varnothing$ ;
2 ${\mathcal{V}}_{\text{init}}=\{v\,|\,(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N],\forall\phi\in\Phi^{K}\wedge{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi^{K}\}$ ;
3 ${\mathcal{H}}_{\text{front}}.\texttt{push}(({\mathcal{V}}_{\text{init}},0))$;
4 $\tau_{\text{explored}}(v)=v,\forall v\in{\mathcal{V}}_{\text{init}}$;
5 Build hierarchy specification graph ${\mathcal{G}}_{h}$;
6 while ${\mathcal{H}}_{\text{front}}\neq\varnothing$ do
7       $(v,c_{v})={\mathcal{H}}_{\text{front}}.{\texttt{pop}}()$;
8       if $v\in{\mathcal{V}}_{\text{explored}}$ then
9             continue;
10
11      ${\mathcal{V}}_{\text{explored}}.\texttt{add}(v)$;
12       if $v.{\mathbf{q}}(1,1)\in{\mathcal{Q}}_{\mathcal{B}}^{F}(1,1)$ then
13             return $\texttt{ExtractPlan}(\Lambda_{\text{explored}}(v))$ ;
14
15      ${\mathcal{V}}_{\text{succ}}=\texttt{GetSucc}(v,{\mathcal{P}},{\mathcal{G}}_{h})$ ;
16       for $(u,c_{vu})\in{\mathcal{V}}_{\text{succ}}$  do
17             $c_{u}=c_{v}+c_{vu}$;
18             if $u\not\in{\mathcal{V}}_{\text{seen}}$ or $c_{u}<{\mathcal{V}}_{\text{seen}}(u)$  then
19                   ${\mathcal{V}}_{\text{seen}}(u)=c_{u}$;
20                   ${\mathcal{H}}_{\text{front}}.\texttt{push}((u,c_{u}))$;
21                   $\Lambda_{\text{explored}}(u)=\Lambda_{\text{explored}}(v)+u$ ;
22
23return $\varnothing$;
Algorithm 2 Search-based STAP
Input: state $v$, hierarchical team models ${\mathcal{P}}$, hierarchy specification graph ${\mathcal{G}}_{h}$
Output: Set of successor states ${\mathcal{V}}_{\text{succ}}$
1 ${\mathcal{V}}_{\text{succ}}=\varnothing$;
2 $(r,\phi,s,q)=(v.r,v.\phi,v.{\mathbf{s}}(v.r),v.{\mathbf{q}}(v.\phi))$ ;
3
[…]
5 if $q\not\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)$ and $v.{\mathbf{q}}(\phi_{\text{prt}})\not\in{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi_{\text{prt}}),\forall\phi_{\text{prt}}\in\Phi_{\text{prt}}$ then
        $\triangleright$ In-spec transitions in ${\mathcal{P}}(r,\phi)$
6       for $((s,q),(s^{\prime},q^{\prime}))\in\to_{{\mathcal{P}}(r,\phi)}$ do
7             $u=v,{\mathcal{V}}_{u}=\varnothing$;
8             $u.{\mathbf{s}}(r)=s^{\prime},\;u.{\mathbf{q}}(\phi)=q^{\prime}$ ;
9             $\pi=\{\phi\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)$ else $\varnothing$ ;
[…]
11             ${\mathcal{V}}_{\text{succ}}.\texttt{add}(({\mathcal{V}}_{u},c_{r}(s,s^{\prime})))$;
12
        $\triangleright$ In-spec switch transitions in ${\mathcal{P}}(\phi)$
13       if $r\neq N$ and $((r,\phi,s,q),(r+1,\phi,v.{\mathbf{s}}(r+1),q))\in\zeta_{\text{in}}$ then
14             $u=v$;
15             $u.r=r+1$ ;
16             ${\mathcal{V}}_{\text{succ}}.\texttt{add}((u,0))$;
17
  $\triangleright$ Type I switch transitions in ${\mathcal{P}}$
18 for $\phi^{\prime}\in\Phi^{K}\setminus\{\phi\}$  do
19       if $((r,\phi,s,q),(r,\phi^{\prime},s,v.{\mathbf{q}}(\phi^{\prime}))\in\zeta_{\text{inter}}^{1}$ then
20             $u=v$;
21             $u.\phi=\phi^{\prime}$ ;
22             ${\mathcal{V}}_{\text{succ}}.\texttt{add}((u,0))$;
23
  $\triangleright$ Type II switch transitions in ${\mathcal{P}}$
24 for $\phi^{\prime}\in\Phi^{K}\setminus\{\phi\}$  do
25       if $((r,\phi,s,q),(1,\phi^{\prime},v.{\mathbf{s}}(1),v.{\mathbf{q}}(\phi^{\prime}))\in\zeta_{\text{inter}}^{2}$ then
26             $u=v$;
27             $u.r=1,\;u.\phi=\phi^{\prime}$ ;
28             ${\mathcal{V}}_{\text{succ}}.\texttt{add}((u,0))$;
29
30return ${\mathcal{V}}_{\text{succ}}$;
Algorithm 3 $\texttt{GetSucc}(v,{\mathcal{P}},{\mathcal{G}}_{h})$
The function GetSucc determines the successor states of a given state $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$ by considering in-spec transitions in ${\mathcal{P}}(r,\phi)$, in-spec switch transitions in ${\mathcal{P}}(v.\phi)$, and inter-spec type I and II switch transitions in ${\mathcal{P}}$. To identify in-spec transitions in ${\mathcal{P}}(r,\phi)$, the function initially checks if the specification $\phi$ or any of its non-leaf parent specifications are satisfied. This is achieved using Parents(${\mathcal{G}}_{h}$, $\phi$), which returns the path from the leaf specification $\phi$ to the topmost specification in the hierarchical specification graph ${\mathcal{G}}_{h}$, excluding $\phi$ itself [line 3]. If $\phi$ or any of its parent specifications are satisfied, the search in ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(\phi)$ ceases, as further exploration is unnecessary. This is based on condition (2) in Definition IV.1 and Assumption IV.6, which state that each leaf specification appears only in one non-leaf specification and satisfying each specification at most once is adequate. If a parent specification is satisfied, the leaf specification no longer contributes to task progress.
Input: True composite propositions $\pi$, state $u$, parent specifications $\Phi_{\text{prt}}$
Output: Updated set of states ${\mathcal{V}}_{u}$
1 $\phi_{\text{prt}}=\Phi_{\text{prt}}.\texttt{pop}()$ ;
2 ${\mathcal{Q}}_{\mathcal{B}}^{\text{succ}}=\left\{q^{\prime}\,|\,(u.{\mathbf{q}}(\phi_{\text{prt}}),\pi,q^{\prime})\in\to_{\mathcal{B}}(\phi_{\text{prt}}),\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{B}}(\phi_{\text{prt}})\right\}$;
3 for $q^{\prime}\in{\mathcal{Q}}^{\text{succ}}_{\mathcal{B}}$ do
4       $u.{\mathbf{q}}(\phi_{\text{prt}})=q^{\prime}$ ;
5       if $\phi_{\text{prt}}=\phi_{1}^{1}$ then
6             ${\mathcal{V}}_{u}.\texttt{add}(u)$ ;
7
8      else
9             $\pi=\{\phi_{\text{prt}}\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi_{\text{prt}})$ else $\varnothing$;
[…]
11
[…]
If none of the parent specifications are satisfied, the state of robot $r$ in $v.{\mathbf{s}}$ and the automaton state of leaf specification $\phi$ in $v.{\mathbf{q}}$ are updated based on in-spec transitions in $\to_{{\mathcal{P}}(r,\phi)}$ [line 3]. Subsequently, the automaton states of all parent specifications of $\phi$ are also updated, as detailed in lines 3-3 and further elaborated in Alg. 4. For the three types of switch transitions, only the robot $r$ (in the case of in-spec switch transitions [line 3)], the leaf specification $\phi$ (for inter-spec type I switch transitions [line 3]), or both robot $r$ and specification $\phi$ (for inter-spec type II switch transitions [line 3]) are updated. This implies a shift in the search space to a different PBA, without modifying $v.{\mathbf{s}}$ and $v.{\mathbf{q}}$. Since these transitions do not change robot states and automaton states, the function UpdateNonLeafSpecs is not invoked.
Given a state $u$ and the set of parent specifications $\Phi_{\text{prt}}$ of the leaf specification $u.\phi$ currently being explored, the function UpdateNonLeafSpecs utilizes a depth-first strategy to refresh the automaton states of non-leaf specifications within $\Phi_{\text{prt}}$. When considering a specific parent specification $\phi_{\text{prt}}\in\Phi_{\text{prt}}$, the function initially identifies all states that can be reached based on the acceptance of the preceding specification at the immediate lower level [lines 4-4]. The automaton state of $\phi_{\text{prt}}$ in $u.{\mathbf{q}}$ is then updated using these reachable states [line 4]. This process continues until the automaton state of the topmost specification is updated [line 4].
Input: Searched path $\Lambda$
Output: state-specification plan $\tau$
1 Remove switch states from $\Lambda$ and divide $\Lambda$ into $l$ path segments with each sharing the same specification ;
2 for $k\in[l]$ do
3       for $v\in\Lambda_{k}$ do
4             $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),\ldots,(s_{r}^{i},\phi_{r}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ where $(s_{r}^{i},\phi_{r}^{i})=(v.{\mathbf{s}}(r),v.\phi)$ and $(s_{r^{\prime}}^{i},\phi_{r^{\prime}}^{i})=(v.{\mathbf{s}}(r^{\prime}),\epsilon),\forall r^{\prime}\not=v.r$ ;
5
6return $\tau$;
Algorithm 5 $\texttt{ExtractPlan}(\Lambda)$
Figure 4: The process of plan extraction: The top bar represents the path, segmented by inter-spec switch transitions, with each segment corresponding to distinct specifications. The middle bar illustrates the path after removing switch states. At the bottom, the derived plan for a team of five robots is displayed. Here, orange segments indicate active periods for the robots, while gray segments represent their inactive phases.
The ExtractPlan function in Alg. 5 processes the path $\Lambda=v_{0}v_{1}\ldots v_{n}$, which leads to the state satisfying the topmost specification, and outputs a state-specification plan $\tau$. This process is visually represented in Fig. 4. A state $v$ within $\Lambda$ is identified as a switch state if it is the end state of an inter-spec switch transition $\zeta_{\text{inter}}^{1}\cup\zeta_{\text{inter}}^{2}$, which has the same robot state and automaton state as the corresponding start state. The path $\Lambda$ is segmented by these switch states, with each segment corresponding to states associated with the same specification, indicative of a search within a specific product team model. To create a concise path $\Lambda^{\prime}$, all switch states are removed, as they merely facilitate transitions between product team models and don’t influence the plan’s progress. This refined path is denoted as $\Lambda^{\prime}=\Lambda_{1}\Lambda_{2}\ldots\Lambda_{l}$, where $\Lambda_{k}$ represents the $k$-th path segment, as shown in line 5. The next step involves converting these sequential path segments into parallel executions. For each segment $\Lambda_{k}$, we identify all active robots ${\mathcal{R}}_{k}$ participating in the specification $\phi_{k}$, represented as ${\mathcal{R}}_{k}=\left\{v.r\,|\,v\in\Lambda_{k}\right\}$. Due to task decomposition [8], each robot in ${\mathcal{R}}_{k}$ can execute its part of the plan independently. Robots not in ${\mathcal{R}}_{k}$ are considered inactive for that segment. The plan horizon post-parallelization is determined by the longest path among all active robots. Note that an active robot might not maintain the active mode during the entire horizon if it completes its actions in a shorter time frame. For each robot state, we pair it with the relevant specification $\phi_{k}$ if the robot is active, or with a null specification $\epsilon$ if it’s inactive [line 5].
Remark VI.2
Our method aligns with the decoupled approaches in the domain MAPF [68, 69], where initially, individual agents independently plan their paths without considering potential collisions. Subsequently, the system assesses for collisions and modifies the paths to prevent conflicts. Although this approach is more scalable and demands less computational effort than fully coupled methods, it may not ensure an optimal outcome, could lead to deadlocks, and might face challenges in dense areas. Resolving conflicts in temporal logic planning proves more difficult because, unlike MAPF, the actions of robots may be temporally correlated. When addressing collisions, it is crucial not only to consider robots that are physically close but also those that are temporally correlated. Therefore, we leave the systemic exploration of a complete collision resolution strategy to future studies.
VI-B Search Acceleration with Heuristics
We introduce three heuristics aimed at expediting the search.
VI-B1 Temporal Order between Specifications
[…]
VI-B2 Essential Switch Transitions
We introduce essential switch transitions to refine the connections between PBAs.
Definition VI.3 (Essential State)
An essential state for robot $r$, denoted as $(r,s)$, is defined either (a) $s=s_{r}^{0}$ or (b) there exists a leaf specification $\phi$ fulfilling the conditions:
  • $((\phi,r,s,q),(\phi,r,s,q^{\prime}))\in\to_{{\mathcal{P}}(r,\phi)}$;
  • $q\neq q^{\prime}$;
  • $q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi)$.
This definition implies that an essential state is either (a) the initial state of robot $r$ or (b) a state where robot $r$ causes an in-spec transition to a decomposition state of a leaf specification for the first time. In the latter case, the intuition is that essential states are pivotal for achieving substantial progress in the task.
Definition VI.4 (Essential Switch Transition)
A switch transition $\zeta=\left((\phi,r,s,q),(\phi^{\prime},r^{\prime},s^{\prime},q^{\prime})\right)\in\zeta_{\text{in}}\cup\zeta_{\text{inter}}^{1}\cup\zeta_{\text{inter}}^{2}$ is an essential switch transition if both $(r,s)$ and $(r^{\prime},s^{\prime})$ are essential states.
The leaf specifications $\phi$ and $\phi^{\prime}$ might not be the exact ones that render the states $(r,s)$ and $(r^{\prime},s^{\prime})$ essential. In other words, $\phi$ and $\phi^{\prime}$ could represent any leaf specifications, which is explained in Ex. 3 below. Consequently, any switch transitions that are not deemed essential will be excluded.
Example 3
[…]
VI-B3 Guidance by Automaton States
Unlike the first two heuristics, which focus on reducing the graph size, the last heuristic aims to speed up the search by prioritizing the expansion of more promising states — those that are closer to achieving satisfaction. For a state $u=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, we define its task-level cost as summation over all leaf specifications:
$c_{q}({\mathbf{q}})&=\sum_{\phi^{\prime}\in\Phi^{K}}c_{q}({\mathbf{q}}(\phi^{\prime}))\\ =\sum_{\phi^{\prime}\in\Phi^{K}}&\max\left\{\texttt{len}(q^{0}_{\mathcal{B}}\to{\mathbf{q}}(\phi^{\prime}))\,|\,\forall\,q^{0}_{\mathcal{B}}\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi^{\prime})\right\}.\end{aligned}$ (4)
The cost $c_{q}({\mathbf{q}}(\phi^{\prime}))$ is determined by the length of the simplest path from any initial state $q^{0}_{\mathcal{B}}$ to the automaton state ${\mathbf{q}}(\phi^{\prime})$ within the NBA ${\mathcal{B}}(\phi)$. A simple path is defined as one that does not include any repeated states, and the function len measures the path’s length in terms of the number of transitions. Essentially, $c_{q}({\mathbf{q}}(\phi))$ captures the maximal progress made in the task $\phi$ since its beginning. A similar heuristic has been employed in [18, 21]. The cost $c_{vu}$ is now calculated as
$c_{vu}=c_{r}\left(v.{\mathbf{s}}(r),u.{\mathbf{s}}(r)\right)+w\left[c_{q}(u.{\mathbf{q}})-c_{q}(v.{\mathbf{q}})\right],$ (5)
where $c_{r}$ is the cost between robot states, $c_{q}(u.{\mathbf{q}})-c_{q}(v.{\mathbf{q}})$ is the cost between automaton states and $w$ is the weight.
Remark VI.5
The first two heuristics essentially alter the structure of the hierarchical team models. As a result, the completeness and optimality of the search approach, when these heuristics are applied, are no longer assured. However, employing exclusively the third heuristic still maintains the completeness of the search, as it ensures the full exploration of the search space eventually. Additionally, this approach with heuristic of automaton state remains optimal with respect to the newly defined cost (5). Extensive simulations presented in Sec. VIII provide empirical evidence that our approach, incorporating all three heuristics, can generate solutions for complex tasks. The rationale behind this empirical completeness is twofold. Firstly, [28] prove that the inference on temporal order is correct for a broad range of flat sc-LTL specifications. By eliminating switch transitions that conflict with precedence orders, transitions that are in accordance with temporal orders remain unaffected. Secondly, the concept of essential switch transitions ensures that the search process only transitions to a different product model when there has been task progress in the current product model, which is reasonable as it is unnecessary to switch to a different task if the ongoing task has not been completed.
VII Theoretic Analysis
We conduct a theoretical analysis of the complexity, soundness, completeness, and optimality of our approach.
Proposition VII.1
For every search state $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, among all the automaton states in ${\mathbf{q}}$ related to leaf specifications, at most one automaton state does not belong to the decomposition set.
Proof:
Based on Def. VI.1, all automaton states within an initial search state are initial automaton states. By default, these are recognized as decomposition states. When a PBA ${\mathcal{P}}(r,\phi)$ is chosen for exploration, only the corresponding automaton state ${\mathbf{q}}(\phi)$ is updated, leaving the automaton states from other leaf specifications unaffected. The search can exit ${\mathcal{P}}(r,\phi)$ only through three types of switch transitions, each, by definition, associated with decomposition states. Upon switching to a different PBA ${\mathcal{P}}(r^{\prime},\phi^{\prime})$, the automaton state of $\phi$, which is a decomposition state, remains unchanged. Therefore, by inductive reasoning, it follows that at any given point, at most one automaton state from the leaf specifications does not belong to the decomposition set. ∎
Note that while the set of automaton states ${\mathbf{q}}$ encompasses states from non-leaf specifications, these are not considered into our analysis of planning complexity, as they entirely depend on the leaf specifications.
Theorem VII.2 (Complexity)
The total number of transitions in the hierarchical team models ${\mathcal{P}}$ is given as follows:
$|\to_{\mathcal{P}}|=\sum_{\phi\in\Phi^{K}}\sum_{r\in[N]}\left|\to_{{\mathcal{P}}(r,\phi)}\right|+\underbrace{|\zeta_{\text{in}}|+|\zeta_{\text{inter}}^{1}|+|\zeta_{\text{inter}}^{2}|}_{|\zeta|},$ (6)
which is the total number of in-spec transitions and switch transitions. According to [8], $|\to_{{\mathcal{P}}(r,\phi)}|$ is bounded by $|\to_{{\mathcal{P}}(r,\phi)}|\leq\left(|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right).$ Furthermore, $|\zeta_{\text{in}}|$, $|\zeta_{\text{inter}}^{1}|$ and $|\zeta_{\text{inter}}^{2}|$ are, respectively, upper bounded by
$|\zeta_{\text{in}}|$ $\leq\sum_{\phi\in\Phi^{K}}\sum_{r\in[N-1]}\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r+1}|\right),$
$|\zeta_{\text{inter}}^{1}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi^{K}}\sum_{r\in[N]}\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{r}|\right),$ (7)
$|\zeta_{\text{inter}}^{2}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi^{K}}\sum_{r\in[N]}\left(|{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{1}|\right).$
Proof:
For a given PBA ${\mathcal{P}}(r,\phi)$, the number of states is upper bounded by $|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|$. Its squared term gives the upper bound for the number of in-spec transitions $|\to_{{\mathcal{P}}(r,\phi)}|$, since any pair of states within the PBA can potentially be connected. Regarding in-spec switch transitions between two consecutive PBAs, ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(r+1,\phi)$, each decomposition state can be associated with any robot state, which sets the upper bound on the number of possible starting states to $|{\mathcal{D}}_{\mathcal{B}}(\phi)|\times|{\mathcal{S}}_{r}|$. Similarly, the maximum number of end states is determined by $|{\mathcal{D}}_{\mathcal{B}}(\phi)|\times|{\mathcal{S}}_{r+1}|$, subsequently defining the upper bound for $|\zeta_{\text{in}}|$. Following the same logic, the upper bound for inter-spec switch transitions can be deduced as outlined in Eq. (VII.2). ∎
Corollary VII.3 (Complexity of Search with Heuristics)
The upper bound on the number of in-spec transitions $|\zeta_{\text{in}}|$ is the same as that in Eq. (VII.2), but the upper bound on the number of switch transitions $|\zeta_{\text{inter}}|$ is obtained by replacing ${\mathcal{S}}_{r}$ in Eq. (VII.2) with $\tilde{{\mathcal{S}}}_{r}$, where $\tilde{{\mathcal{S}}}_{r}$ is the set of essential states of robot $r$.
Theorem VII.4 (Soundness)
The state-specification plan given by Alg. 2 satisfies hierarchical sc-LTL specifications $\Phi$.
Proof:
The soundness of Alg. 2 stems directly from its correct-by-construction design. To prove it, we use Alg. 1 to validate the correctness of the returned state-specification plan $\tau$. This plan $\tau$ is divided into segments $\tau=\tau_{1}\tau_{2}\ldots\tau_{l}$, with each $\tau_{k}$ aligning with a corresponding path segment $\Lambda_{k}$ from the searched path $\Lambda$. Beginning with the first segment $\tau_{1}$, and as the algorithm iterates over each $\phi\in\Phi^{K}\setminus\Phi^{acc}$ in line 1, the loop continues as long as $\phi\not=\phi_{1}$, where $\phi_{1}$ is the leaf specification associated with $\tau_{1}$. This is because $S(i)=\varnothing$ in line 1 for any $\phi\not=\phi_{1}$. When $\phi=\phi_{1}$, the set of reachable states ${\mathcal{Q}}_{\mathcal{B}}^{i+1}$ in NBA ${\mathcal{B}}(\phi_{1})$ are updated accordingly. Consequently, progress is only made for $\phi_{1}$, while all other leaf specifications remain idle. Upon completing the iteration over $\tau_{1}$ segment, $\phi_{1}$ reaches a decomposition state in NBA ${\mathcal{B}}(\phi_{1})$, as per the design of the hierarchical team models. In line 1 of Alg. 1, the non-leaf parent specifications of $\phi_{1}$ are updated based on the progress made solely in $\phi_{1}$. This update mirrors the operations performed by the UpdateNonLeafSpecs function in Alg. 4. The same procedure is applied to the subsequent segments $\tau_{2},\tau_{3},\ldots$, with each segment facilitating progress on a single leaf specification while pausing others. When an accepting state for a particular specification is reached, all PBAs associated with its child leaf specifications cease to be searched, as specified in Alg. 3, line 3. This is in alignment with the process in Alg. 1, line 1, where child leaf specifications of a satisfied parent specification are not further examined. Ultimately, upon reaching the end of $\tau_{l}$, the topmost specification is fulfilled, as noted in Alg. 2, line 2. This completion signifies that the hierarchical sc-LTL specifications are satisfied in accordance with Alg. 1, line 1, thereby confirming the correctness of the entire process. ∎
Assumption VII.5 (Segmented Plan)
Assuming that there exists a feasible state-specification plan $\tau=\tau^{0}\tau^{1}\ldots\tau^{h}$ that can be sequentially divided into $\ell$ segments with $\ell\leq|\Phi^{K}|$, represented as $\tau=\langle\tau\rangle_{1}\langle\tau\rangle_{2}\ldots\langle\tau\rangle_{\ell}$ where no two segments share the same specification. Each segment $\langle\tau\rangle_{l}$ fulfills a single leaf specification. That is, for every aggregated state-specification pair $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ in the segment $\langle\tau\rangle_{l}=\tau^{m}\tau^{m+1}\ldots\tau^{n}$, each $\phi^{i}_{j}$ is either the same leaf specification $\phi$ that $\langle\tau\rangle_{l}$ satisfies, or it is the null specification $\epsilon$.
Note that the number of segments $\ell$ might be less than the number of leaf specifications $|\Phi^{K}|$, since it is possible that not all leaf specifications are required to be satisfied, such as leaf specifications that are connected by disjunction. The following properties only apply to Alg. 2 without heuristics.
Theorem VII.6 (Completeness)
Assuming that a state-specification plan $\tau$ satisfying Asm. VII.5 exists, Alg. 2 is complete in that it is guaranteed to return a feasible plan that fulfills the hierarchical sc-LTL specifications $\Phi$.
Proof:
We proved the completeness of Alg. 2 by first disregarding inter-spec type I switch transitions. The inclusion of type I transitions merely enlarges the search space, without impacting the algorithm’s completeness. In this case, only inter-spec type II switch transitions exist between team models of any two leaf specifications $\phi$ and $\phi^{\prime}$. Type II transitions connect accepting states of every PBA within the product team model of $\phi$ (or $\phi^{\prime}$) to decomposition states of the first PBA of the team model of $\phi^{\prime}$ (or $\phi$). As a result, the search does not advance to product team models of other leaf specifications until it discovers a path segment that satisfies the current one, e.g., $\phi$. Furthermore, once an accepting state of a PBA for $\phi$ is reached, the search only moves to initial states, not other decomposition states, of the first PBA in another team model, e.g., $\phi^{\prime}$, as the search on this first PBA has not started yet.
Following this logic, Alg. 2 initiates from an initial state of a certain specification $\phi$. Upon reaching one of its accepting states, the search transitions to the initial state of another specification $\phi^{\prime}$. This process is repeated until the termination condition is satisfied. Given that we employ a Dijkstra-based approach and the search within a product team model is complete as per [8], Alg. 2 is complete even without considering inter-spec type I switch transitions, concluding the proof. ∎
Theorem VII.7 (Optimality)
Assuming the existence of a state-specification plan $\tau^{*}$ that meets Asm. VII.5 with least cost, Alg. 2 is optimal in that the cost of the plan produced by Alg. 2 is equal to or less than the cost of $\tau^{*}$.
Proof:
To prove the optimality, we first set aside inter-spec type I switch transitions. Following the same logic as in the proof of completeness, and employing a Dijkstra-based search algorithm, we can assert that searching within a product team model is optimal, as established in [8]. The addition of inter-spec type I switch transitions expands the search space, allowing for the potential discovery of a feasible plan that might not adhere to Asm. VII.5 yet could yield a lower cost. ∎
VIII Simulation Experiments
Figure 5: Pick-and-place platform for user study.
We use Python 3.10.12 on a computer with 3.5 GHz Apple M2 Pro and 16G RAM. Our code can be available via the https://github.com/XushengLuo92/Hierarchical-LTL-STAP.
VIII-A User Study on Understanding Hierarchical sc-LTL
[…]
VIII-A1 Survey
We carry out the survey on Prolific (www.prolific.com), recruiting 25 anonymous participants. The screening criteria specify that participants must have a STEM background and be at least undergraduate students, ensuring familiarity with mathematical logic. We begin the survey with an introduction to linear temporal logic, using the sorting task as a demonstration. Following this, participants answer two types of questions: subjective and objective. The subjective section consists of 7 questions where participants are presented with natural language instructions for 7 tasks. For each task, we provide a flat sc-LTL formula and hierarchical formulas. The participants are asked to choose which formula is easier to understand and clearer in representing the task described, of which task 5 is, “First place the circle, and then place diamond and star in any order, and lastly place the heart.” The flat sc-LTL is
$&\Diamond(s_{\text{cc}}\wedge\Diamond t_{\text{cc}})\wedge\Diamond(s_{\text{dm}}\wedge\Diamond t_{\text{dm}})\wedge\Diamond(s_{\text{st}}\wedge\Diamond t_{\text{st}})\wedge\Diamond(s_{\text{ht}}\wedge\Diamond t_{\text{ht}})\\ &\wedge\neg s_{\text{dm}}\mathcal{U}t_{\text{cc}}\wedge\neg s_{\text{dm}}\mathcal{U}t_{\text{cc}}\wedge\neg s_{\text{st}}\mathcal{U}t_{\text{cc}}\wedge\neg s_{\text{ht}}\mathcal{U}t_{\text{dm}}\wedge\neg s_{\text{ht}}\mathcal{U}t_{\text{st}}\end{aligned}$
and the hierarchical sc-LTL is
$L_{1}:\quad&\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond(\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}))\\ L_{2}:\quad&\phi_{2}^{1}=\Diamond(s_{\text{cc}}\wedge\Diamond t_{\text{cc}})\\ &\phi_{2}^{2}=\Diamond(s_{\text{dm}}\wedge\Diamond t_{\text{dm}})\wedge\Diamond(s_{\text{st}}\wedge\Diamond t_{\text{st}})\\ &\phi_{2}^{3}=\Diamond(s_{\text{ht}}\wedge\Diamond t_{\text{ht}})\end{aligned}$ (8)
For the objective section, we prepare 4 pairs of flat and hierarchical formulas, each pair corresponding to the same task. This results in a total of 8 questions, with each one featuring either a flat sc-LTL formula or hierarchical sc-LTL formulas. We randomly order the formulas so that adjacent questions do not necessarily involve the same task. Participants are required to select from four natural language instructions the one that correspond to the given formula.
Figure 6: The count of preferences per task.
VIII-A2 Results analysis
Responses with low accuracy on objective questions were excluded, indicating that respondents lack a solid understanding of temporal logics. This results in 20 valid surveys. For the subjective questions, we compute the percentage of participants who prefer flat versus hierarchical formulas across 140 instances (20 participants, 7 questions each). For the objective questions, we calculate the accuracy for each type of sc-LTL across 80 instances per sc-LTL type (20 participants, 4 questions per sc-LTL type). We track the time taken by each participant to respond to each question. A detailed analysis of the survey results is provided below.
For the subjective questions, 27.1% of all responses show a preference for flat sc-LTL, while 72.9% favor hierarchical sc-LTL. A detailed count of preferences for each task is illustrated in Fig. 6, where respondents showed a clear preference over hierarchical sc-LTL except for task 4. The specification is that “Place the circle, diamond, star and heart. Diamond should be after circle and heart should be after star.” The temporal constraints are simpler compared to those in task 5 in (8), making flat sc-LTL adequate for this scenario. Regarding the objective questions, the accuracy rates are 65.0% for flat sc-LTL and 86.3% for hierarchical sc-LTL. Additionally, the average response time for questions involving flat sc-LTL formulas is 78.3 seconds, compared to 47.1 seconds for hierarchical sc-LTL formulas. These findings further validate the user-friendliness of hierarchical sc-LTL, as evidenced by both higher accuracy and reduced response times. Details are presented in Fig. 7, where respondents achieved higher accuracy rates in all tasks specified under hierarchical sc-LTL.
Figure 7: The accuracy and response time per task.
VIII-B Scenarios
We use the office environment described in [8] for service tasks, as depicted in Fig. 8. This 30×$\times 7$ grid map features 14 desk areas, 6 meeting rooms, and several other functional rooms. Additionally, we distribute $m$ robots at various locations within this simulated environment.
The first three scenarios are derived from the work [8]. We outline the hierarchical specifications for these scenarios below, while the corresponding flat versions of these specifications are referred to [8]. Note that the hierarchical specifications were derived directly from task descriptions, rather than from flat specifications. In Def. IV.1, we limit the sc-LTL specifications to co-safe LTL. However, all hierarchical sc-LTL specifications below contain a sub-formula expressed as $\square(a\Rightarrow b)$. Even with this sub-formula, it is still sufficient to assert satisfaction upon reaching the accepting states.
VIII-B1 Scenario 1
Empty a paper bin located at desk $d_{5}$. The robot must avoid the public area while transporting the bin.
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}$
$L_{2}:\quad$ $\phi_{2}^{1}=\Diamond(d_{5}\wedge\mathsf{default}\wedge\bigcirc((\mathsf{carrybin}\;\mathcal{U}\;\mathsf{dispose})\wedge\Diamond\mathsf{default}))$
$\quad\quad\quad\quad\wedge\square(\mathsf{carrybin\Rightarrow\neg\mathsf{public}})$
$\phi_{2}^{2}=\Diamond(d_{5}\wedge\mathsf{emptybin}\wedge\bigcirc(d_{5}\wedge\mathsf{default}))$
Figure 8: The office building in a grid-based layout, where areas $d_{1}$ to $d_{14}$ represent desks, $m_{1}$ to $m_{6}$ are meeting rooms, $e$ stands for the elevator, $g$ for the garbage room, $p$ for the printer room, and $k$ for the coffee kitchen. Areas marked as “public” indicate public spaces. Obstacles are illustrated in gray. The locations of robots are shown as numbered red dots.
VIII-B2 Scenario 2
Distribute documents to desks $d_{10}$, $d_{7}$, and $d_{5}$, and avoid public areas while carrying the document.
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$
$L_{2}:\quad$ $\phi_{2}^{1}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{10}\wedge\bigcirc\neg\mathsf{carry}))\wedge\mathsf{notpublic}$
$\phi_{2}^{2}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{7}\wedge\bigcirc\neg\mathsf{carry}))\wedge\mathsf{notpublic}$
$\phi_{2}^{3}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{5}\wedge\bigcirc\neg\mathsf{carry}))\wedge\mathsf{notpublic}$
$\mathsf{notpublic}:=\square(\mathsf{carry}\Rightarrow\neg\mathsf{public})$
scenario $t$ $t_{\text{heur}}$ $t_{\text{heur}}^{1}$ $t_{\text{heur}}^{2}$ $t_{\text{heur}}^{3}$ $c$ $c_{\text{heur}}$ $c_{\text{heur}}^{1}$ $c_{\text{heur}}^{2}$ $c_{\text{heur}}^{3}$
1 242.0$\pm$3.1 2.5$\pm$0.8 242.0$\pm$3.1 11.1$\pm$0.2 41.7$\pm$33.5 63.7$\pm$6.9 77.4$\pm$1.2 63.7$\pm$6.9 66.8$\pm$8.7 80.3$\pm$8.2
1 232.9$\pm$18.0 2.3$\pm$0.4 93.7$\pm$14.6 20.8$\pm$6.7 27.9$\pm$20.4 81.4$\pm$8.8 82.8$\pm$6.7 82.1$\pm$6.9 82.7$\pm$9.2 82.4$\pm$9.7
2 1036.7$\pm$13.6 4.1$\pm$0.9 1036.7$\pm$13.6 46.0$\pm$1.8 21.1$\pm$13.8 91.5$\pm$6.7 98.5$\pm$5.6 91.5$\pm$6.7 94.7$\pm$ 8.6 95.4$\pm$7.2
2 859.6$\pm$15.4 2.1$\pm$0.2 283.2$\pm$2.4 39.8$\pm$0.8 289.7$\pm$211.1 93.3$\pm$3.2 94.4$\pm$2.2 96.3$\pm$5.9 95.0$\pm$5.9 93.6$\pm$5.4
3 timeout 10.2$\pm$3.4 timeout 594.6$\pm$23.0 69.6$\pm$40.1 120.2$\pm$3.4 106.5$\pm$4.6 116.1$\pm$13.6
3 timeout 19.1$\pm$10.4 1371.0$\pm$28.6 572.5$\pm$20.9 timeout 120.0$\pm$5.7 119.4$\pm$7.3 117.6$\pm$6.4
TABLE II: Comparison of results between methods using heuristics and those not employing heuristics. Scenarios 1, 2, and 3 are modifications of the first three scenarios by arranging non-leaf specifications sequentially. The symbol $t$ represents the runtime without using any heuristics. The terms $t_{\text{heur}}$, $t_{\text{heur}}^{1}$, $t_{\text{heur}}^{2}$, and $t_{\text{heur}}^{3}$ correspond to the runtimes when applying all heuristics, the heuristic of temporal order, the heuristic of essential switch transitions, and the heuristic of automaton state, respectively. This notation is similarly used for the cost, denoted as $c$.
VIII-B3 Scenario 3
Take a photo in meeting rooms $m_{1}$, $m_{4}$, and $m_{6}$. The camera should be turned off for privacy reasons when not in meeting rooms. Deliver a document from desk $d_{5}$ to $d_{3}$, ensuring it does not pass through any public areas, as the document is internal and confidential. Guide a person waiting at desk $d_{11}$ to meeting room $m_{6}$.
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$
$L_{2}:\quad$ $\phi_{2}^{1}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}\wedge\Diamond\phi_{3}^{3}$
$\phi_{2}^{2}=\Diamond(d_{5}\wedge\mathsf{carry}\;\mathcal{U}\;(d_{3}\wedge\bigcirc\neg\mathsf{carry}))\wedge\mathsf{notpublic}$
$\phi_{2}^{3}=\Diamond(d_{11}\wedge\mathsf{guide}\;\mathcal{U}\;(m_{6}\wedge\bigcirc\neg\mathsf{guide}))$
$L_{3}:\quad$ $\phi_{3}^{1}=\Diamond(m_{1}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$
$\phi_{3}^{2}=\Diamond(m_{4}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$
$\phi_{3}^{3}=\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$
$\mathsf{meeting}:=\;m_{1}\vee m_{2}\vee m_{3}\vee m_{4}\vee m_{5}\vee m_{6}$
VIII-B4 Combinations of scenarios 1, 2 and 3
We examine combinations of any two of these tasks as well as the combination of all three. Due to space constraints, we only detail the scenario involving the combination of all three tasks below.
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$
$L_{2}:\quad$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{1}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{2}=\Diamond\phi_{3}^{3}\wedge\Diamond\phi_{3}^{4}\wedge\Diamond\phi_{3}^{5}}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{3}=\Diamond\phi_{3}^{6}\wedge\Diamond\phi_{3}^{7}\wedge\Diamond\phi_{3}^{8}}$
$L_{3}:\quad$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{1}=\Diamond(d_{5}\wedge\mathsf{default}}\wedge\bigcirc((\mathsf{carrybin}\;\mathcal{U}\;\mathsf{dispose})\wedge\Diamond\mathsf{default}))$
$\quad\quad\quad\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\wedge\square(\mathsf{carrybin\Rightarrow\neg\mathsf{public}})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{2}=\Diamond(d_{5}\wedge\mathsf{emptybin}\wedge\bigcirc(d_{5}\wedge\mathsf{default})})$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{3}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{10}\wedge\bigcirc\neg\mathsf{carry}))}\wedge\mathsf{notpublic}$ (9)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{4}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{7}\wedge\bigcirc\neg\mathsf{carry}))}\wedge\mathsf{notpublic}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{5}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{5}\wedge\bigcirc\neg\mathsf{carry}))}\wedge\mathsf{notpublic}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{6}=\Diamond\phi_{4}^{1}\wedge\Diamond\phi_{4}^{2}\wedge\Diamond\phi_{4}^{3}}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{7}=\Diamond(d_{5}\wedge\mathsf{carry}\;\mathcal{U}\;(d_{3}\wedge\bigcirc\neg\mathsf{carry}))}\wedge\mathsf{notpublic}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{8}=\Diamond(d_{11}\wedge\mathsf{guide}\;\mathcal{U}\;(m_{6}\wedge\bigcirc\neg\mathsf{guide}))}$
$L_{4}:\quad$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{4}^{1}=\Diamond(m_{1}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{4}^{2}=\Diamond(m_{4}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})}$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{4}^{3}=\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})}$
scenario $l_{\text{flat}}$ $l_{\text{hier}}$ ${\mathcal{A}}_{\text{flat}}$ ${\mathcal{A}}_{\text{hier}}$ $t_{\text{flat}}$ $t_{\text{hier}}$ $c_{\text{flat}}$ $c_{\text{hier}}$
1 18 19 (17, 39) (12, 13) 14.1$\pm$3.7 4.9$\pm$2.3 71.0$\pm$6.3 69.0$\pm$5.7
2 24 35 (56, 326) (20, 31) 39.6$\pm$2.5 7.0$\pm$3.1 90.4$\pm$4.0 88.6$\pm$7.0
3 35 52 (180, 1749) (30, 49) timeout 14.5$\pm$4.1 97.1$\pm$5.9
1 $\wedge$ 2 43 57 (868, 12654) (36, 49) timeout 16.4$\pm$7.1 148.8$\pm$8.4
1 $\wedge$ 3 54 74 (2555, 69858) (46, 67) timeout 47.9$\pm$22.3 166.4$\pm$8.0
2 $\wedge$ 3 60 90 (6056, 325745) (54, 85) timeout 42.5$\pm$16.9 175.4$\pm$6.9
1 $\wedge$ 2 $\wedge$ 3 79 111 timeout (70, 112) timeout 89.6$\pm$28.1 246.6$\pm$8.3
TABLE III: The comparative analysis focuses on two different types of sc-LTL specifications. We denote the lengths of the flat and hierarchical sc-LTL specifications as $l_{\text{flat}}$ and $l_{\text{hier}}$, respectively. The sizes of the corresponding NBAs are represented by ${\mathcal{A}}_{\text{flat}}$ and ${\mathcal{A}}_{\text{hier}}$, which detail the number of nodes and edges, with the node count listed first. In terms of solutions, the runtimes for the flat and hierarchical sc-LTL specifications are indicated by $t_{\text{flat}}$ and $t_{\text{hier}}$, respectively. Additionally, the plan horizons, or the lengths of the solutions, for the flat and hierarchical specifications are denoted by $c_{\text{flat}}$ and $c_{\text{hier}}$, respectively.
VIII-C Ablation Study on Effects of Heuristics
We set $m=2$. The results of the effects of different heuristics in scenarios 1 to 3 are detailed in Tab. II, which presents the average runtimes and cost over 20 trials per scenario. During each trial, the locations of the robots are randomly sampled within areas without labels. Our method was evaluated under three conditions: without any heuristics, with a single heuristic (either the temporal order as in Sec. VI-B1, critical switch transitions as in Sec. VI-B2, or automaton state as in Sec. VI-B3), and with the combination of all heuristics. It was observed that in scenarios 1 to 3 where non-leaf specifications have no precedence temporal order, the impact of solely using the temporal order heuristic is equivalent to not using any heuristic at all. To further investigate this, three additional scenarios were introduced where non-leaf specifications were arranged sequentially. For example, in scenario 1, $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}$ was modified to $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2})$. Each navigational and manipulative action carried out incurs a cost of 1, but transitions between switches do not incur any costs. A maximum time limit of one hour was set for these tests, same as other simulations.
As expected, the method without heuristics yields solutions with minimum overall costs. However, employing all three heuristics simultaneously results in a substantial acceleration of the search process, roughly by two orders of magnitude. This significant improvement in runtimes outweighs the minor increase in costs associated with the use of heuristics. In six different scenarios, the heuristic based on essential switch transitions outperforms the other two heuristics in four scenarios. Its effectiveness is attributed to its ability to significantly reduce the number of switch transitions among team models, thus keeping the search confined within individual team models at the most time. Conversely, the temporal order heuristic is the least effective in five of the six scenarios. This inefficiency stems from the absence of precedent constraints in the non-leaf specifications of scenarios 1, 2, and 3. The cost differences between these heuristics are insignificant. In conclusion, each heuristic demonstrates improvement over the baseline method, and all of them should be applied in order to maximize the performance gain.
VIII-D Comparison with Existing Works
We use $m=6$ robots and compare our method with the approach in [8], both using all heuristics. Note that [8]’s method is tailored to flat sc-LTL specifications. Due to the absence of open-source code, we implemented their method. Robot locations in each scenario are randomly assigned within the free space. The performance, in terms of average runtimes and costs over 20 runs, is detailed in Tab. III and includes the length of formulas and sizes of automatons. The length of a formula is the total number of logical and temporal operators. We observed that, contrary to simulation findings in [42], where hierarchical sc-LTL shows syntactical conciseness compared to flat sc-LTL, our hierarchical sc-LTL specifications are longer in length. We hypothesize that the syntactical succinctness of hierarchical sc-LTL becomes more evident in scenarios with complex interdependencies between tasks at non-bottom levels, which are more challenging to express through flat sc-LTL specifications. This is not the case in our scenarios, where tasks at non-bottom levels are temporally independent. Upon reviewing the results, [8]’s method failed to generate solutions for the last four tasks within the one-hour limit. For tasks 1 and 2, our method produced solutions more quickly and with comparable costs. The failure of  [8]’s approach is attributed to the excessively large automatons it generates, sometimes with hundreds of thousands of edges, e.g., 325745 edges for scenario $2\wedge 3$, making the computation of the decomposition set time-consuming as it requires iterating over all possible runs. For the most complex scenario $1\wedge 2\wedge 3$, generating an automaton within one hour is impossible. In contrast, our method was able to find a solution in around 90 seconds.
flattened specifications ${\mathcal{A}}_{\text{hier}}$ $t_{\text{hier}}$ $c_{\text{hier}}$
None (70, 112) 89.6$\pm$28.1 246.6$\pm$8.3
$\phi_{3}^{6}$ (64, 109) 98.4$\pm$50.0 245.4$\pm$10.3
$\phi_{3}^{6},\phi_{2}^{1}$ (69, 135) 173.5$\pm$118.1 247.9$\pm$8.7
$\phi_{3}^{6},\phi_{2}^{1},\phi_{2}^{2}$ (105, 430) 309.4$\pm$487.4 (7) 237.6$\pm$12.7
$\phi_{3}^{6},\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}$ (261, 2133) timeout
TABLE IV: Results illustrating the impact of hierarchy granularity on performance. The first column lists the set of non-leaf specifications that have been flattened. The number within the parentheses in the third column indicates the number of runs that resulted in a timeout.
(a) Runtimes w.r.t number of robots
(b) Cost w.r.t number of robots
Figure 9: Scalability results of runtimes and cost w.r.t number of robots.
VIII-E Effects of Granularity of Hierarchy
For the same task, the performance of the planning algorithm is influenced by the granularity of hierarchical specifications, given the variability in humans’ levels of hierarchical reasoning granularity. To investigate this, we evaluate the most complex task, which combines scenarios 1, 2, and 3 as defined in (VIII-B4). The specifications in (VIII-B4) represent the finest granularity. We then progressively increase the coarseness by flattening more non-leaf specifications through the merging of their respective leaf specifications, thus gradually transforming the hierarchical structure towards a more flattened form. Eventually, this process results in a single flat specification. For example, flattening specification $\phi_{3}^{6}$ results in $\phi_{3}^{6}=\Diamond(m_{1}\wedge\mathsf{photo})\wedge\Diamond(m_{4}\wedge\mathsf{photo})\wedge\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$. The outcomes from 20 runs are presented in Tab. IV. Generally, as more non-leaf specifications are flattened, the automatons increase in size, resulting in longer runtimes, while the costs remain relatively stable. Consequently, finer granularity in the hierarchy is associated with enhanced planning performance.
VIII-F Scalability
For assessing scalability, we varied the number of robots from 10 to 30. The statistical findings are presented in Fig. 9. Across different scenarios, a consistent pattern emerged: as the number of robots increased, the runtimes tended to increase, while the overall costs decreased. This trend is intuitive since a greater number of robots introduces a wider range of potential solutions. In the most complex scenario, our method was capable of identifying solutions in approximately 200 seconds, even with 30 robots. For tasks 1, 2, and 3, increasing the number of robots to 20 or 30 does not result in a substantial cost reduction compared to using 10 robots, indicating that 10 robots are sufficient for tasks of this complexity. However, when tasks are combined, the cost reduction becomes more significant with 20 or 30 robots. Additionally, 20 robots are capable of handling tasks of this complexity effectively, with the marginal benefit of increasing to 30 robots being minor.
Figure 10: The presented keyframes illustrate the execution of the derivative task described in Eq. (VIII-G) within the AI2-THOR framework, indicating the crucial phases in the completion of task. Each of the four robots is inside yellow boxes with distinct numbers. The upper-left corner of the keyframes marks the time steps, and the objects being manipulated are encircled in blue. Note that the robots 2 and 4 are simultaneously manipulating the box and the laptop.
VIII-G Evaluation on simulator AI2-THOR
We conduct our experiments using the AI2-THOR simulator and the ALFRED dataset. AI2-THOR [70] provides a realistic 3D simulated environment for robots to perform interactions within various household settings such as kitchens and living rooms. This setting allows for the assessment of algorithms in tasks that include navigation and object manipulation. The ALFRED dataset [71], used within AI2-THOR, focuses on natural language instructions for simple, sequentially ordered tasks. Examples of such tasks in a living room setting include (i) “Turn off the floor lamp,” (ii) “Move the box to a sofa,” (iii) “Move the laptop from the table to an armchair,” and (iv) “Dispose of the credit card in the garbage can.” We refer to these tasks as base tasks. Then we construct derivative tasks by combining several base tasks within the same floor plan. One derivative task from base tasks (i)-(iv) is that “First turn off the floor lamp. After that, move the box to a sofa and the laptop to an armchair. Finally, dispose of the credit card in the garbage can.” The hierarchical sc-LTL formulas are:
$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond(\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}))$
$L_{2}:\quad$ $\phi_{2}^{1}=\Diamond(\mathsf{floorlamp}\wedge\bigcirc\mathsf{turnoff})$
$\phi_{2}^{2}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}$ (10)
$\phi_{2}^{3}=\Diamond(\mathsf{creditcard}\wedge\bigcirc(\mathsf{pickup}\wedge\Diamond(\mathsf{garbagecan}\wedge\bigcirc\mathsf{place})))$
$L_{3}:\quad$ $\phi_{3}^{1}=\Diamond(\mathsf{box}\wedge\bigcirc(\mathsf{pickup}\wedge\Diamond((\mathsf{sofa1}\vee\mathsf{sofa2})\wedge\bigcirc\mathsf{place})))$
$\phi_{3}^{2}=\Diamond(\mathsf{laptop}\wedge\bigcirc(\mathsf{pickup}\wedge\Diamond((\mathsf{armchair1}$
$\quad\quad\vee\mathsf{armchair2}\vee\mathsf{armchair3}\vee\mathsf{armchair4})\wedge\bigcirc\mathsf{place}))).$
We represent navigation destinations using nouns and the manipulative actions to be performed at these destinations with verbs. In $\phi_{3}^{2}$, given that there are four armchair available and the instructions do not specify a particular one, placing the laptop on any of these armchairs meets the task requirements.
We vary the number of base tasks included in each derivative task, selecting from the set {1, 2, 3, 4}. A total of 35 derivative tasks are constructed, each comprising an equal number of randomly chosen base tasks from the dataset. For each derivative, we assign 1, 2, or 4 robots, each placed at random starting locations within the floor plan. This results in $4\times 35\times 3=420$ testing scenarios. The floor map is transformed into a grid world where each grid unit measures 0.25 meters. The grid world’s dimensions vary from 10 to 40 units per side due to various floor plans. Recall that in objective (3), $c_{r}$ denotes the cost for robot $r$. In this part, we modified the goal (3) as
$J(\tau)=\omega\max_{r\in[N]}c_{r}+(1-\omega)\sum_{r=1}^{N}c_{r}$ (11)
where $\omega\in[0,1]$. Here, setting $\omega=0$ simplifies (11) to objective (3), while $\omega=1$ seeks to reduce the maximum cost among robots, with the secondary term regularizing the costs of the remaining robots. We set $\omega=0.9$ to prioritize minimizing the maximum path length among the robots. The results, detailed in Table V, indicate an increase in runtime for solutions as the number of base tasks or robots increases. Additionally, with more robots engaged, the maximum cost per robot is reduced due to the ability to distribute tasks more evenly among multiple robots. The execution of derivative task (VIII-G) is illustrated in Fig. 10.
# base tasks # robots $t$ $c$
1 1 2.6$\pm$0.9 11.7$\pm$8.7
2 3.0$\pm$1.2 11.7$\pm$9.8
4 4.1$\pm$1.3 9.0$\pm$7.6
2 1 4.3$\pm$2.9 23.4$\pm$21.0
2 4.8$\pm$3.2 17.1$\pm$15.4
4 6.5$\pm$4.7 14.8$\pm$12.7
3 1 10.8$\pm$7.7 38.2$\pm$25.4
2 12.1$\pm$9.2 27.9$\pm$21.6
4 14.6$\pm$9.1 22.0$\pm$14.2
4 1 13.0$\pm$13.9 50.1$\pm$29.6
2 13.3$\pm$14.9 30.7$\pm$16.9
4 14.9$\pm$11.4 23.9$\pm$14.0
TABLE V: Results comparing performance across different numbers of base tasks and robots. Here, $t$ represents the runtimes required to return solutions, and $c$ denotes the maximal cost incurred by any robot.
IX Conclusion and Discussion
In this work, we addressed the problem of STAP for multiple robots subject to hierarchical sc-LTL specifications. First, we presented the syntax and semantics for hierarchical sc-LTL. Next, we developed a search-based planning algorithm. To further expedite the search, we introduced several heuristics leveraging the task structure. We also provided theoretical analysis on the completeness and optimality.
Our simulations, focused on service tasks involving navigation and manipulation, included an ablation study on heuristics. This study demonstrated that each heuristic variably accelerates the search, with their combination yielding a significant speedup. Comparative studies with existing works illustrate that our approach can handle complex tasks beyond the reach of current methods. Scalability tests reveal that our method can manage up to 30 robots within 200 seconds. We attribute the computational efficiency to two factors. First, the adoption of hierarchical specifications results in smaller automatons, as indicated in Tab. III. We contend that hierarchical specifications do not necessarily increase cognitive load, given that humans naturally excel at hierarchical thinking [6, 7]. The results from the user study also highlighted the user-friendliness of hierarchical sc-LTL. The second factor is the design of the planning algorithm, particularly its heuristics. As shown in Tab. II, the planning algorithm performs faster with any heuristic than without. Furthermore, employing a combination of all three heuristics results in shorter runtimes compared to using any single heuristic.
While the hierarchical sc-LTL introduced has proven effective, there are still several areas that require additional investigation. Firstly, the planning algorithm assumes that robots operate independently, without considering scenarios where coordination, such as jointly carrying heavy items, is necessary. Therefore, it is crucial to include such situations in the algorithm. Another area of interest is addressing uncertainty arising from multiple sources, including partially observable environments, or the costs of actions, which might not be predetermined. Since hierarchical sc-LTL can encompass multiple conflicting specifications, a future research avenue could be to define and maximize degree of satisfaction when all specifications cannot be concurrently met. Lastly, integrating natural languages, as in work [72], to embed task hierarchies from human instructions into structured planning languages suitable for existing planners, presents an intriguing prospect.

### v4

@@
 Abstract
-Research in robotic planning with temporal logic specifications, such as syntactically co-safe Linear Temporal Logic (sc-LTL), has relied on single formulas.
-However, as task complexity increases, sc-LTL formulas become lengthy, making them difficult to interpret and generate, and straining the computational capacities of planners.
-To address this, we introduce a hierarchical structure to sc-LTL specifications with both syntax and semantics, proving it to be more expressive than flat counterparts.
-We conducted a user study that compared the flat sc-LTL with our hierarchical version and found that users could more easily comprehend complex tasks using the hierarchical structure.
+Research in robotic planning with temporal logic specifications, such as Linear Temporal Logic (LTL), has relied on single formulas.
+However, as task complexity increases, LTL formulas become lengthy, making them difficult to interpret and generate, and straining the computational capacities of planners.
+To address this, we introduce a hierarchical structure for a widely used specification type—LTL on finite traces (LTLf).
+The resulting language, termed H-LTLf, is defined with both its syntax and semantics.
+We further prove that H-LTLf is more expressive than its standard “flat” counterparts.
+Moreover, we conducted a user study that compared the standard LTLf with our hierarchical version and found that users could more easily comprehend complex tasks using the hierarchical structure.
 We develop a search-based approach to synthesize plans for multi-robot systems, achieving simultaneous task allocation and planning.
-This method approximates the search space by loosely interconnected sub-spaces, each corresponding to an sc-LTL specification.
-The search primarily focuses on a single sub-space, transitioning to another under conditions determined by the decomposition of automatons.
+This method approximates the search space by loosely interconnected sub-spaces, each corresponding to an LTLf specification.
+The search primarily focuses on a single sub-space, transitioning to another under conditions determined by the decomposition of automata.
 We develop multiple heuristics to significantly expedite the search.
@@
 As shown in [5], which involves a task of collecting 5 keys and then opening 5 doors, it took approximately half an hour on a laptop with an Intel i7 CPU and 32 GB of RAM to convert the corresponding LTL into an automaton consisting of 65 nodes and 792 edges.
-Similarly, in our simulations, an automaton could not be generated within an hour on a laptop with Apple M2 Pro and 16G RAM.
-This issue stems from the usage of the “flat” form of LTL, that is, putting all requirements on robot behaviors into a single LTL formula, which tends to become difficult to solve for complex tasks.
+Similarly, in our simulations, an automaton could not be generated for a complex task within an hour on a laptop with Apple M2 Pro and 16G RAM.
+This issue stems from the usage of the standard “flat” form of LTL, that is, putting all requirements on robot behaviors into a single LTL formula, which tends to become difficult to solve for complex tasks.
 However, one key observation is that robot tasks often have a loose connection and can be broken down into smaller components.
 Studies suggest that humans prefer hierarchical task specification, which improves interpretability of planning and execution, making it easier to identify ongoing work and conveniently adjust unfeasible parts without affecting other components [6, 7].
-Inspired by this, we propose a hierarchical form of LTL specifications.
-Example 1
-Consider a mobile manipulation task where robots are tasked with first picking and placing item $a$ and then fetching items $b$ and $c$ in any order. The “flat” specification is
-ϕ=italic-$\phi\,=\,$ $\Diamond(s_{a}\wedge\Diamond(t_{a}\wedge\Diamond(s_{b}\wedge\Diamond t_{b})\wedge\Diamond(s_{c}\wedge\Diamond t_{c}))),$
-[…]
-[…]
-[…]
-[…]
-[…]
-The exact formulas can be found in Eq.
-(2), with syntax and semantics elaborated in Sec.
-IV.
-This representation effectively reduces the length of each formula, leading to smaller automatons and improving interpretability.
-$\square$
-In this work, we first establish syntax rules for hierarchical syntactically co-safe LTL (sc-LTL) and then provide a framework for semantic interpretation.
-To do this, considering the context of multiple specifications, we introduce the concept of state-specification plan, which pairs the state of each robot with a specific specification it seeks to fulfill at a given time.
-Utilizing this concept, we devise an algorithm to check the satisfaction of a set of hierarchical sc-LTL specifications given a state-specification plan.
-Theoretically, we prove that the hierarchical form is more expressive compared to the flat form.
+Inspired by this, we propose a hierarchical form of a widely used specification type—LTL on finite traces (LTLf[8], which can be satisfied by finite sequence of system states.
+This characteristic makes LTLf apt for modeling and reasoning about systems with finite durations, such as those found in the robotics field.
+In this work, we first establish syntax and semantics for H-LTLf.
+To do this, considering the context of multiple specifications, we introduce the concept of state-specification sequence, which pairs the state of a system with a specification it seeks to fulfill at a given time.
+Theoretically, we prove that the hierarchical form is more expressive compared to the standard form.
 Practically, we conducted a user study to show that using a hierarchical structure makes it easier to express complex tasks with temporal constraints.
-Next, we propose a planning algorithm with formal guarantees of completeness and optimality under mild assumptions.
-The bottom-up approach involves searching within a product graph that combines the environment and task space.
+Next, we propose a bottom-up planning algorithm that involves searching within a product graph that combines the environment and task space.
 This method’s advantage is it can simultaneously achieve task allocation and planning (STAP).
-Inspired by [8], we approximate the entire search space as a set of loosely interconnected sub-spaces, each corresponding to an individual sc-LTL specification in the hierarchical structure.
-The search process primarily occurs within a single sub-space, with transitions to adjacent sub-spaces happening under conditions determined by the decomposition of the automatons.
-Under certain assumptions regarding the state-specification plan, we demonstrate that our planning algorithm is sound, complete, and optimal.
+Particularly, inspired by [9], we approximate the entire search space as a set of loosely interconnected sub-spaces, each corresponding to an individual LTLf specification in the hierarchical structure.
+The search process primarily occurs within a single sub-space, with transitions to adjacent sub-spaces happening under conditions determined by the decomposition of the automata.
+Under certain assumptions regarding the state-specification sequence, we demonstrate that our planning algorithm is sound, complete, and optimal.
 Additionally, we introduce several heuristics to expedite the search process, including minimizing the loose connections between sub-spaces and leveraging task-level progress to guide the search.
@@
 In scalability tests, our method demonstrated the ability to generate sub-optimal solutions in approximately three minutes for complex tasks involving up to 30 robots.
-Note that the flat version of these tasks could not be solved using graph-theoretic methods due to the inability to generate an automaton within a one-hour timeout.
+The standard version of these tasks could not be solved using graph-theoretic methods due to the inability to generate an automaton within a one-hour timeout.
 A glossary of key notations is provided in Tab.
@@
   1. 1.
-    We introduce a hierarchical form of LTL and prove it is more expressive than the flat form;
+    We introduce a hierarchical form of LTLf and prove it is more expressive than the standard form;
   2. 2.
-    We conduct a user study to show that the hierarchical sc-LTL is easier to interpret compared to the flat sc-LTL.
+    We conduct a user study to show that the H-LTLf  is easier to interpret compared to the standard LTLf;
   3. 3.
     We develop a search-based planning algorithm, achieving simultaneous task allocation and planning.
-This planner is, to the best of our knowledge, the first to offer both completeness and optimality for hierarchical sc-LTL;
+This planner is, to the best of our knowledge, the first to offer both completeness and optimality for H-LTLf;
   4. 4.
@@
 Some attempts have been made to simplify this, such as merging multiple atomic propositions into one using logical operators [19, 22], or combining multiple sub-formulas into one formula using logical operators [15, 14, 27].
-Despite these efforts, we still categorize these as the flat form.
+Despite these efforts, we still categorize these as the standard form.
 Notably, temporal operators, the feature distinguishing LTL from propositional logic, only appear on one side, either for the integration of merged propositions or inside the sub-formulas.
@@
 In our study, we introduce composite propositions that encompass more than one sub-formula and can be combined using temporal operators.
-Besides, various extensions of temporal logic have been introduced: [26] developed Counting Temporal Logic (CTL) to express the collective behavior of multiple robots.
+Besides, various extensions of temporal logic have been introduced: [27] developed Counting Linear Temporal Logic (cLTL) to express the collective behavior of multiple robots.
 [32, 33] utilized Graph Temporal Logic (GTL) and Swarm Signal Temporal Logic (SwarmSTL) respectively to specify swarm characteristics like centroid positioning and density distribution.
 [28] introduced Capability Temporal Logic (CaTL) to encapsulate the diverse capabilities of robots.
-However, all works mentioned focus on using flat LTL specifications to describe the behaviors in multi-robot systems.
+However, all works mentioned focus on using standard LTL specifications to describe the behaviors in multi-robot systems.
 Global specifications that do not explicitly assign tasks to robots generally require decomposition to derive the task allocation, which can be accomplished in three ways: The most common method, utilized in works such as [9, 21, 34, 35, 36, 29, 37], involves decomposing a global specification into multiple tasks, which leverages the transition relations within the automaton, which is the graphical representation of an LTL formula.
@@
 [42] expanded further into a multi-objective setting, integrating conflicting objectives such as cost and success probability into a Markov Decision Process (MDP).
-Our work presents key differences: (i) While the aforementioned studies focus on flat LTL specifications, we consider hierarchical specifications and can solve tasks which are unsolvable by these earlier methods.
+Our work presents key differences: (i) While the aforementioned studies focus on standard LTL specifications, we consider hierarchical specifications and can solve tasks which are unsolvable by these earlier methods.
 (ii) In the mentioned studies, tasks between robots are independent thus robot can execute their tasks in parallel.
 In contrast, our approach also assigns tasks with temporal dependencies to different robots, thus robots may depend on others.
-Note that [42], developed around the same time, is the only existing work that investigated decision-making under hierarchical sc-LTL.
+Note that [43], developed around the same time, is the only existing work that investigated decision-making under H-LTLf.
 This work differs from ours in that, algorithmically, [43] proposed a hierarchical approach that first focused on task allocation and then planning, unlike STAP here.
-In terms of task decompoistion, [42] treated each transition in the automaton as an individual task to be allocated.
-In contrast, in this work individual tasks can consist of several transitions or even span the entire automaton, making each task potentially equivalent to multiple tasks as defined in [42].
 Theoretically, the planning algorithm introduced in [43] lacks formal guarantees of either completeness or optimality.
@@
 The domain of Multi-Robot Task Allocation (MRTA) is thoroughly studied, and various classifications of MRTA have been developed [1, 44], organizing the extensive research in this field.
-In terms of MRTA, the problem is this work can be is characterised by four aspects: (i) single-task robots (each robot is limited to executing one task at a time), (ii) single-robot tasks (only one robot is required for each task), (iii) instantaneous assignment (tasks are allocated under all available information at this moment), and (iv) cross-schedule interdependence (the assignment of a robot to a task is influenced not solely by its own schedule but also by the schedules of other robots within the system).
+In terms of MRTA, the problem is this work can be is characterized by four aspects: (i) single-task robots (each robot is limited to executing one task at a time), (ii) single-robot tasks (only one robot is required for each task), (iii) instantaneous assignment (tasks are allocated under all available information at this moment), and (iv) cross-schedule interdependence (the assignment of a robot to a task is influenced not solely by its own schedule but also by the schedules of other robots within the system).
 Methods for MRTA primarily fall into two categories: market-based [45] and optimization-based approaches [46].
@@
 II-C Hierarchical Task Models
-Hierarchical reasoning enhances human understanding of the world [6, 7].
+Studies [6, 7] suggest that hierarchical reasoning enhances humans’ ability to understand the world more effectively.
 In classical AI planning, researchers crafted task models reflecting hierarchical structures by employing domain control knowledge [57].
@@
 However, despite the widespread use of hierarchical task models in classical AI planning, it’s intriguing to note the lack of specification hierarchy in temporal logic robotic planning.
-Our work differs from these studies as we follow an inverse direction; instead of integrating LTL into HTN to express the goal of the planning problem, we incorporate HTN into LTL, allowing for hierarchical structures within multiple LTL formulas, making them more capable of expressing complex tasks than a single flat LTL formula.
+Our work differs from these studies as we follow an inverse direction; instead of integrating LTLf into HTN to express the goal of the planning problem, we incorporate HTN into LTLf, allowing for hierarchical structures within multiple LTLf formulas, making them more capable of expressing complex tasks than a single LTLf formula.
 Notation Description Notation Description
-$\pi$ atomic propositions $\sigma$ infinite word over the alphabet
-[…]
-$\Diamond$ eventual operator ${\mathcal{Q}}_{\mathcal{B}}$ set of automaton states
-$\square$ global operator ${\mathcal{Q}}_{\mathcal{B}}^{0}$ set of initial automaton states
-$\mathcal{U}$ until operator $\Sigma=2^{{\mathcal{A}}{\mathcal{P}}}$ alphabet
-$\phi$ specification $\to_{\mathcal{B}}$ transition relation in ${\mathcal{B}}$
-${\mathcal{A}}{\mathcal{P}}$ set of atomic propositions ${\mathcal{Q}}_{\mathcal{B}}^{F}$ set of final automaton states
-${\mathcal{T}}(r)$ transition system of robot $r$ $\rho_{\mathcal{B}}$ infinite run
-${\mathcal{S}}_{r}$ states of robot $r$ ${\mathcal{L}}_{\mathcal{B}}$ accepted language of ${\mathcal{B}}$
+$\sigma$ finite word over the alphabet ${\mathcal{A}}(\phi)$ NBA of specification $\phi$
+${\mathcal{Q}}_{\mathcal{A}}^{0}$ set of initial automaton states ${\mathcal{Q}}_{\mathcal{A}}$ set of automaton states
+$\Sigma=2^{{\mathcal{A}}{\mathcal{P}}}$ alphabet $\to_{\mathcal{A}}$ transition relation in ${\mathcal{A}}$
+${\mathcal{A}}{\mathcal{P}}$ set of atomic propositions ${\mathcal{Q}}_{\mathcal{A}}^{F}$ set of final automaton states
+${\mathcal{T}}(r)$ transition system of robot $r$ $\rho_{\mathcal{A}}$ finite run
+${\mathcal{S}}_{r}$ states of robot $r$ ${\mathcal{L}}_{\mathcal{A}}$ accepted language of ${\mathcal{A}}$
 $s_{r}^{0}$ initial state of robot $r$ ${\mathcal{P}}(r,\phi)$ PBA of robot $r$ and specification $\phi$
@@
 $\phi_{k}^{i}$ $i$-th specification at the $k$-th level ${\mathcal{Q}}_{\mathcal{P}}^{F}$ set of final states
-${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$ specification hierarchy graph $c_{\mathcal{P}}$ cost function in ${\mathcal{P}}$
-$\Phi^{k}$ set of specifications at the $k$-th level $\tau=\tau^{0}\ldots\tau^{h}$ state-specification plan
-$\Phi^{K}$ set of specifications at the lowest level ${\mathcal{D}}_{\mathcal{B}}(\phi)$ decomposition set of specification $\phi$
-$\Phi^{acc}$ set of satisfied specifications up to certain time ${\mathcal{P}}(\phi)$ product team model associated with specification $\phi$
+${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$ specification hierarchy tree $c_{\mathcal{P}}$ cost function in ${\mathcal{P}}$
+$\Phi_{k}$ set of specifications at the $k$-th level $\tau=\tau_{0}\ldots\tau_{h}$ state-specification sequence
+$\Phi_{K}$ set of specifications at the lowest level ${\mathcal{D}}_{\mathcal{A}}(\phi)$ decomposition set of specification $\phi$
+$\Phi_{\text{leaf}}$ set of leaf specifications ${\mathcal{P}}(\phi)$ product team model associated with specification $\phi$
 $\zeta_{\text{in}}$ in-spec switch transition ${\mathcal{P}}$ hierarchical team models
@@
 III Preliminaries
-Notation: Let $\mathbb{N}$ denote the set of all integers, $[K]=\{1,\ldots,K\}$ denote the set of integers from 1 to $K$ and ||$\cdot$|| ⋅ | denote the cardinality of a set.
-In this section, we formally describe Linear Temporal Logic (LTL) by presenting its syntax and semantics. Also, we briefly review preliminaries of automata-based LTL model checking.
+Notation: Let $\mathbb{N}$ denote the set of all integers, $[K]=\{0,\ldots,K\}$ and $[K]_{+}=\{1,\ldots,K\}$ represent the sets of integers from 0 to $K$ and from 1 to $K$, respectively, and ||$\cdot$|| ⋅ | denote the cardinality of a set.
+In this section, we introduce LTLf by presenting its syntax and semantics, and automata-based LTLf model checking.
 […]
@@
 For brevity, we abstain from deriving other Boolean and temporal operators, e.g., disjunction $\vee$, implication $\Rightarrow$, always $\square$, eventually $\lozenge$, which can be found in [64].
-[…]
-Definition III.1 (NBA)
-[…]
-  • ${\mathcal{Q}}_{{\mathcal{B}}}$ is the set of states;
-  • ${\mathcal{Q}}_{{\mathcal{B}}}^{0}\subseteq{\mathcal{Q}}_{{\mathcal{B}}}$ is a set of initial states;
+LTLf [8] is a variant of LTL that is interpreted over finite sequences of states while maintaining the same syntax as standard LTL.
+A finite word $\sigma$ over the alphabet $2^{\mathcal{AP}}$ is defined as a finite sequence $\sigma=\sigma_{0}\sigma_{1}\ldots\sigma_{h}$ with $\sigma_{i}\in 2^{{\mathcal{A}}{\mathcal{P}}}$ for $i\in[h]$.
+The satisfaction of an LTLf formula $\phi$ over a sequence $\sigma$ at an instant $i$, for $i\in[h]$, is inductively defined as follows:
+  • $\sigma,i\models\pi\;\;\text{iff}\;\pi\in\sigma_{i}$.
+  • $\sigma,i\models\neg\phi\;\;\text{iff}\;\sigma,i\not\models\phi$.
+  • $\sigma,i\models\phi_{1}\wedge\phi_{2}\;\;\text{iff}\;\sigma,i\models\phi_{1}\;\text{and}\;\sigma,i\models\phi_{2}$.
+  • $\sigma,i\models\bigcirc\phi\;\;\text{iff}\;i<h\;\text{and}\;\sigma,i+1\models\phi$.
+  • $\sigma,i\models\phi_{1}\,{\mathcal{U}}\,\phi_{2}$ iff for some $j$ such that $i\leq j\leq h$, we have that $\sigma,j\models\phi_{2}$, and for all $k,i\leq k<j$, we have that $\sigma,k\models\phi_{1}$.
+A formula $\phi$ is satisfied by $\sigma$, denoted by $\sigma\models\phi$, if $\sigma,0\models\phi$.
+The language $\texttt{Words}(\phi)=\left\{\sigma|\sigma\models\phi\right\}$ is defined as the set of words that satisfy the formula $\phi$.
+An LTLf formula $\phi$ can be translated into a Nondeterministic Finite Automaton:
+Definition III.1 (NFA)
+A Nondeterministic Finite Automaton (NFA) ${\mathcal{A}}$ of an LTLf formula $\phi$ over $2^{\mathcal{AP}}$ is defined as a tuple ${\mathcal{A}}(\phi)=\left({\mathcal{Q}}_{{\mathcal{A}}},{\mathcal{Q}}_{{\mathcal{A}}}^{0},\Sigma,\rightarrow_{{\mathcal{A}}},\mathcal{Q}_{\mathcal{A}}^{F}\right)$, where
+  • ${\mathcal{Q}}_{{\mathcal{A}}}$ is the set of states;
+  • ${\mathcal{Q}}_{{\mathcal{A}}}^{0}\subseteq{\mathcal{Q}}_{{\mathcal{A}}}$ is a set of initial states;
   • $\Sigma=2^{\mathcal{AP}}$ is an alphabet;
-  • $\rightarrow_{{\mathcal{B}}}\,\subseteq\,{\mathcal{Q}}_{{\mathcal{B}}}\times\Sigma\times{\mathcal{Q}}_{{\mathcal{B}}}$ is the transition relation;
-  • ${\mathcal{Q}}_{{\mathcal{B}}}^{F}\subseteq{\mathcal{Q}}_{{\mathcal{B}}}$ is a set of accepting/final states.
-An infinite run $\rho_{{\mathcal{B}}}$ of ${\mathcal{B}}$ over an infinite word $\sigma=\pi_{0}\pi_{1}\pi_{2}\dots$, $\pi_{k}\in\Sigma=2^{\mathcal{AP}}$, $\forall k\in\mathbb{N}$, is a sequence $\rho_{{\mathcal{B}}}=q_{{\mathcal{B}}}^{0}q_{{\mathcal{B}}}^{1}q_{{\mathcal{B}}}^{2}\dots$ such that $q_{{\mathcal{B}}}^{0}\in{\mathcal{Q}}_{{\mathcal{B}}}^{0}$ and $(q_{{\mathcal{B}}}^{k},\pi_{k},q_{{\mathcal{B}}}^{k+1})\in\rightarrow_{{\mathcal{B}}}$, $\forall k\in\mathbb{N}$.
-An infinite run $\rho_{{\mathcal{B}}}$ is called accepting if $\texttt{Inf}(\rho_{{\mathcal{B}}})\cap{\mathcal{Q}}_{{\mathcal{B}}}^{F}\neq\varnothing$, where $\texttt{Inf}(\rho_{{\mathcal{B}}})$ represents the set of states that appear in $\rho_{{\mathcal{B}}}$ infinitely often.
-The words $\sigma$ that produce an accepting run of ${\mathcal{B}}$ constitute the accepted language of ${\mathcal{B}}$, denoted by ${\mathcal{L}}_{{\mathcal{B}}}$.
-Then [63] proves that the accepted language of ${\mathcal{B}}$ is equivalent to the words of $\phi$, i.e., ${\mathcal{L}}_{{\mathcal{B}}}=\texttt{Words}(\phi)$.
-The dynamics of robot $r$ is captured by a Transition System (TS) defined as follows:
+  • $\rightarrow_{{\mathcal{A}}}\,\subseteq\,{\mathcal{Q}}_{{\mathcal{A}}}\times\Sigma\times{\mathcal{Q}}_{{\mathcal{A}}}$ is the transition relation;
+  • ${\mathcal{Q}}_{{\mathcal{A}}}^{F}\subseteq{\mathcal{Q}}_{{\mathcal{A}}}$ is a set of accepting/final states.
+A finite run $\rho_{{\mathcal{A}}}$ of ${\mathcal{A}}$ over a finite word $\sigma=\sigma_{0}\sigma_{1}\dots\sigma_{h}$, $\sigma_{i}=2^{\mathcal{AP}}$, $\forall i\in[h]$, is a sequence $\rho_{{\mathcal{A}}}=q_{{\mathcal{A}}}^{0}q_{{\mathcal{A}}}^{1}\dots q_{{\mathcal{A}}}^{h+1}$ such that $q_{{\mathcal{A}}}^{0}\in{\mathcal{Q}}_{{\mathcal{A}}}^{0}$ and $(q_{{\mathcal{A}}}^{i},\sigma_{i},q_{{\mathcal{A}}}^{i+1})\in\rightarrow_{{\mathcal{A}}}$, $\forall i\in[h]$.
+A run $\rho_{{\mathcal{A}}}$ is called accepting if $q_{{\mathcal{A}}}^{h+1}\in{\mathcal{Q}}_{{\mathcal{A}}}^{F}$.
+The words $\sigma$ that produce an accepting run of ${\mathcal{A}}$ constitute the accepted language of ${\mathcal{A}}$, denoted by ${\mathcal{L}}_{{\mathcal{A}}}$.
+Then [64] proves that the accepted language of ${\mathcal{A}}$ is equivalent to the words of $\phi$, i.e., ${\mathcal{L}}_{{\mathcal{A}}}=\texttt{Words}(\phi)$.
+The dynamics of robot $r$ is captured by a Transition System:
 Definition III.2 (TS)
 A Transition System for robot $r$ is a tuple ${\mathcal{T}}(r)=\{{\mathcal{S}}_{r},s_{r}^{0},\to_{r},{\mathcal{A}}{\mathcal{P}}_{r},{\mathcal{L}}_{r}\}$ where:
-  • ${\mathcal{S}}_{r}=\{s_{r}\}$ is the set of discrete states of robot $r$;
-  • $s_{r}^{0}$ is the initial state of robot $r$;
+  • ${\mathcal{S}}_{r}$ is the set of discrete states of robot $r$, and $s_{r}\in{\mathcal{S}}_{r}$ denotes a specific state;
+  • $s_{r}^{0}\in{\mathcal{S}}_{r}$ is the initial state of robot $r$;
   • $\to_{r}\,\subseteq\,{\mathcal{S}}_{r}\times{\mathcal{S}}_{r}$ is the transition relation;
@@
   • ${\mathcal{L}}_{r}:{\mathcal{S}}_{r}\rightarrow 2^{\mathcal{AP}_{r}}$ is the observation (labeling) function that returns a subset of atomic propositions that are satisfied, i.e., ${\mathcal{L}}_{r}(s_{r})\subseteq\mathcal{AP}_{r}$.
+A finite sequence of states $s_{r}^{0},s_{r}^{1},\ldots,s_{r}^{h}$ is referred to as a trace.
+Given the transition system ${\mathcal{T}}(r)$ of robot $r$ and the NFA ${\mathcal{A}}$ of an LTLf formula $\phi$, we can define the Product Automaton (PA) ${\mathcal{P}}(r,\phi)={\mathcal{T}}(r)\times{\mathcal{A}}(\phi)$ as follows [64]:
+Definition III.3 (PA)
+For a robot $r$ and an LTLf formula $\phi$, the Product Automaton is defined by the tuple ${\mathcal{P}}(r,\phi)=(\mathcal{Q}_{\mathcal{P}},\mathcal{Q}_{\mathcal{P}}^{0},\longrightarrow_{{\mathcal{P}}},\mathcal{Q}_{\mathcal{P}}^{F})$, where
+  • $\mathcal{Q}_{\mathcal{P}}={\mathcal{S}}_{r}\times\mathcal{Q}_{{\mathcal{A}}}$ is the set of product states;
+  • $\mathcal{Q}_{\mathcal{P}}^{0}=\{s_{r}^{0}\}\times\mathcal{Q}_{\mathcal{A}}^{0}$ is a set of initial states;
+  • $\longrightarrow_{{\mathcal{P}}}\subseteq\mathcal{Q}_{\mathcal{P}}\times\mathcal{Q}_{\mathcal{P}}$ is the transition relation defined by the rule: $\frac{(s_{r}\rightarrow_{r}s^{\prime}_{r})\wedge(q\xrightarrow{{\mathcal{L}}_{r}\left(s_{r}\right)}_{{\mathcal{A}}}q^{\prime})}{q_{{\mathcal{P}}}=\left(s_{r},q\right)\longrightarrow_{\mathcal{P}}q_{{\mathcal{P}}}^{\prime}=\left(s^{\prime}_{r},q^{\prime}\right)}$.
+The transition from the state $q_{\mathcal{P}}\in\mathcal{Q}_{\mathcal{P}}$ to $q_{\mathcal{P}}^{\prime}\in\mathcal{Q}_{\mathcal{P}}$, is denoted by $(q_{\mathcal{P}},q_{\mathcal{P}}^{\prime})\in\longrightarrow_{\mathcal{P}}$, or $q_{\mathcal{P}}\longrightarrow_{\mathcal{P}}q_{\mathcal{P}}^{\prime}$;
+  • $\mathcal{Q}_{\mathcal{P}}^{F}={\mathcal{S}}_{r}\times\mathcal{Q}_{\mathcal{A}}^{F}$ is a set of accepting/final states;
+IV Hierarchical LTLf
+We first present the syntax and semantics of H-LTLf, then we analyze its expressiveness.
+IV-A Syntax of H-LTLf
+Definition IV.1 (Hierarchical LTLf)
+H-LTLf is structured into $K$ levels, labeled $L_{1},\ldots,L_{K}$, arranged from the highest to the lowest.
+Each level $L_{k}$ with $k\in[K]_{+}$ contains $n_{k}$ LTLf formulas.
+The H-LTLf specification is represented as $\Phi=\left\{\phi_{k}^{i}\,|\,k\in[K]_{+},i\in[n_{k}]_{+}\right\}$, where $\phi_{k}^{i}$ denotes the $i$-th LTLf formula at level $L_{k}$.
+Let $\Phi_{k}$ denote the set of formulas at level $L_{k}$, and let $\texttt{Prop}(\phi_{k}^{i})$ represent the set of propositions appearing in formula $\phi_{k}^{i}$.
+The H-LTLf follows these rules:
+  1. 1.
+    There is exactly one formula at the highest level: $n_{1}=1$.
+  2. 2.
+    Each formula at level $L_{k}$ consists either entirely of atomic propositions, i.e., $\texttt{Prop}(\phi_{k}^{i})\subseteq{\mathcal{A}}{\mathcal{P}}$, or entirely of formulas from the next lower level, i.e., $\texttt{Prop}(\phi_{k}^{i})\subseteq\Phi_{k+1}$.
+  3. 3.
+    Each formula at level $L_{k+1}$ appears in exactly one formula at the next higher level: $\phi_{k+1}^{i}\in\bigcup_{j\in[n_{k}]_{+}}\texttt{Prop}(\phi_{k}^{j})$ and $\texttt{Prop}(\phi_{k}^{j_{1}})\cap\texttt{Prop}(\phi_{k}^{j_{2}})=\varnothing$, for $j_{1},j_{2}\in[n_{k}]_{+}$ and $j_{1}\not=j_{2}$.
+For a formula $\phi_{k}^{i}$ at a non-highest level, we slightly bend the notation to use $\phi_{k}^{i}$ to represent the same symbol at the higher level $L_{k-1}$, which we refer to as composite proposition.
+$\phi_{k}^{i}$ represents not only the $i$-th formula at level $L_{k}$, but also the corresponding composite proposition at level $L_{k-1}$.
+When $\phi_{k}^{i}$ appears at the right side in a certain formula, we consider it as a composite proposition.
+When it appears at the left side as a standalone formula, we refer to it as a specification.
+Example 1 (H-LTLf)
+Consider a task where robots are tasked with first picking and placing item $a$ and then fetching items $b$ and $c$ in any order. The LTLf specification is
+ϕ=italic-$\phi\,=\,$ $\Diamond(s_{a}\wedge\Diamond(t_{a}\wedge\Diamond(s_{b}\wedge\Diamond t_{b})\wedge\Diamond(s_{c}\wedge\Diamond t_{c}))),$
+where sub-formula $\Diamond(s_{x}\wedge\Diamond t_{x})$ denotes the event of first picking item $x$ from its source location $s_{x}$ and then placing it at the target location $t_{x}$.
+This formula is less interpretable by arranging all sub-formulas side by side.
+Furthermore, its automaton has 17 states and 102 transitions, which is unexpectedly large for a task of this complexity.
+The H-LTLf specifications of the same task are
 […]
-Definition III.3 (PBA)
-[…]
-  • $\mathcal{Q}_{\mathcal{P}}={\mathcal{S}}_{r}\times\mathcal{Q}_{{\mathcal{B}}}$ is the set of product states;
-  • $\mathcal{Q}_{\mathcal{P}}^{0}=\{s_{r}^{0}\}\times\mathcal{Q}_{\mathcal{B}}^{0}$ is a set of initial states;
-  • $\longrightarrow_{{\mathcal{P}}}\subseteq\mathcal{Q}_{\mathcal{P}}\times\mathcal{Q}_{\mathcal{P}}$ is the transition relation defined by the rule: $\frac{(s_{r}\rightarrow_{r}s^{\prime}_{r})\wedge(q\xrightarrow{{\mathcal{L}}_{r}\left(s_{r}\right)}_{{\mathcal{B}}}q^{\prime})}{q_{{\mathcal{P}}}=\left(s_{r},q\right)\longrightarrow_{\mathcal{P}}q_{{\mathcal{P}}}^{\prime}=\left(s^{\prime}_{r},q^{\prime}\right)}$.
-The transition from the state $q_{\mathcal{P}}\in\mathcal{Q}_{\mathcal{P}}$ to $q_{\mathcal{P}}^{\prime}\in\mathcal{Q}_{\mathcal{P}}$, is denoted by $(q_{\mathcal{P}},q_{\mathcal{P}}^{\prime})\in\longrightarrow_{\mathcal{P}}$, or $q_{\mathcal{P}}\longrightarrow_{\mathcal{P}}q_{\mathcal{P}}^{\prime}$;
-  • $\mathcal{Q}_{\mathcal{P}}^{F}={\mathcal{S}}_{r}\times\mathcal{Q}_{\mathcal{B}}^{F}$ is a set of accepting/final states;
-IV Hierarchical Syntactically Co-safe LTL
-In this work, we focus on a particular subset of LTL formulas known as syntactically co-safe LTL, or sc-LTL for short [65].
-As established by [65], any LTL formula encompassing only the temporal operators $\Diamond$ and $\mathcal{U}$ and written in positive normal form (where negation is exclusively before atomic propositions) is classified under syntactically co-safe formulas.
-This category does not include the $\square$ operator.
-Sc-LTL formulas can be satisfied by finite sequences followed by any infinite repetitions.
-This characteristic makes sc-LTL apt for modeling and reasoning about systems with finite durations, such as those found in the robotics field.
-We first present the syntax of hierarchical sc-LTL, then we propose the interpretation of semantics of hierarchical sc-LTL specifications, along with the analysis of the expressiveness.
-IV-A Syntax of Hierarchical sc-LTL
-Definition IV.1 (Hierarchical sc-LTL)
-Hierarchical sc-LTL is structured into $K$ levels, labeled as $L_{1},\ldots,L_{K}$, arranged from the highest to the lowest.
-Each level $L_{k}$, where $k\in[K]$, contains $n_{k}$ sc-LTL formulas.
-The hierarchical sc-LTL can be represented as $\Phi=\left\{\phi_{k}^{i}\,|\,k\in[K],i\in[n_{k}]\right\}$, where $\phi_{k}^{i}$ denotes the $i$-th sc-LTL formula at level $L_{k}$.
-The hierarchical sc-LTL adheres to the following rules:
-  1. 1.
-    Each formula at a given level $L_{k}$, for $k\in[K-1]$, is derived from the formulas at the next lower level $L_{k+1}$.
-  2. 2.
-    Every formula at any level other than the highest (i.e., $k=2,\ldots,K$) is included in exactly one formula at the next higher level $L_{k-1}$.
-  3. 3.
-    Atomic propositions are used exclusively within the formulas at the lowest level $L_{K}$.
-Let $\Phi^{k}$ denote the set of formulas at level $L_{k}$ with $k\in[K]$.
-For any formula $\phi_{k}^{i}$ at the non-bottom level $k<K$, we slightly bend the notation to use $\phi_{k}^{i}$ to represent the same symbol at the higher level $L_{k-1}$, which we refer to as composite proposition.
-In this manner, $\phi_{k}^{i}$ represents not only the $i$-th formula at level $L_{k}$, but also the corresponding composite proposition at level $L_{k-1}$.
-In this work, when $\phi_{k}^{i}$ appears at the right side in a certain formula, we consider it as a composite proposition.
-On the other hand, when it appears at the left side as a standalone formula, we refer to it as a specification.
-The direct result of condition 3) is that non-bottom specifications consist solely of composite propositions.
-Example 1
-continued (Hierarchical sc-LTL) The hierarchical specifications for the task of first picking and placing item $a$ and then fetching items $b$ and $c$ in any order are:
-[…]
-where there are two levels, $L_{1}$ and $L_{2}$, with $L_{1}$ having one specification and $L_{2}$ having three specifications.
-Their automatons have a total of 8 states and 18 transitions.
-The symbol $\phi_{2}^{1}$ is a specification at level $L_{2}$ but is a composite proposition at level $L_{1}$.
-$\Phi^{1}=\{\phi_{1}^{1}\}$ and $\Phi^{2}=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}$.
-$\square$
-As discussed in Section II-A, local and global forms of task specifications are distinguished based on whether robots receive tasks from individual specifications or from a unified specification.
-We highlight that the direction distinguishing local from global forms is complementary to the direction distinguishing flat from hierarchical forms.
-Thus, the hierarchical form can be integrated with both local and global forms.
-Here, we employ the global form of hierarchical sc-LTL.
-Definition IV.2 (Specification Hierarchy Graph)
-The specification hierarchy graph, denoted as ${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$, is a tree where each node represents a specification within the hierarchical sc-LTL, and an edge $(u,v)$ indicates that specification $u$ contains specification $v$ as a composite proposition.
-(a) Specification hierarchy graph
-(b) NBA of sc-LTL $\Diamond a\wedge\Diamond b$
-Figure 1: The specification hierarchy graph corresponds to hierarchical sc-LTL specifications in (2).
-The NBA corresponds to the sc-LTL specification in Example 2, where non-trivial decomposition states are colored in gray.
+There are two levels, $L_{1}$ and $L_{2}$, with $L_{1}$ having one formula and $L_{2}$ having three formulas.
+The symbol $\phi_{2}^{1}$, appearing on the left side of the equal sign, is a specification at level $L_{2}$ but is a composite proposition at level $L_{1}$ since it is on the right side of the equal sign.
+$\Phi_{1}=\{\phi_{1}^{1}\}$ and $\Phi_{2}=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}$, $\texttt{Prop}(\phi_{1}^{1})=\{\phi_{2}^{1},\phi_{2}^{2},\phi_{2}^{3}\}\subseteq\Phi_{2}$ and $\texttt{Prop}(\phi_{2}^{1})=\{s_{a},t_{a}\}\subseteq{\mathcal{A}}{\mathcal{P}}$.
+This representation reduces the length of each formula, leading to smaller automata and improving interpretability.
+Their automata have a total of 8 states and 18 transitions.
+Note that the hierarchical representation is not unique.
+Depending on the level of granularity in the hierarchy, the same task can be represented as:
+$L_{1}:\quad$ $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2})$
+$L_{2}:\quad$ $\phi_{2}^{1}=\Diamond(s_{a}\wedge\Diamond t_{a}),\quad\phi_{2}^{2}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}$ (3)
+$\hskip 28.45274ptL_{3}:\quad$ $\phi_{3}^{1}=\Diamond(s_{b}\wedge\Diamond t_{b}),\quad\,\phi_{3}^{2}=\Diamond(s_{c}\wedge\Diamond t_{c}).\hskip 22.76228pt\square$
+Definition IV.2 (Specification hierarchy tree)
+The specification hierarchy tree, denoted as ${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$, is a tree where each node represents a specification within the H-LTLf, and an edge $(u,v)$ indicates that specification $u$ contains specification $v$ as a composite proposition.
+Any H-LTLf specifications can be turned into a specification hierarchy tree.
+(a) Specification hierarchy tree
+(b) NFA of LTLf $\Diamond a\wedge\Diamond b$
+Figure 1: The specification hierarchy tree of H-LTLf specifications in (1) where green nodes represent leaf specifications.
+The NFA corresponds to the LTLf specification in Example 2, where non-trivial decomposition states are colored in gray.
+From a tree perspective, the level $k$ of a specification node $\phi_{k}^{i}$ corresponds to its depth, defined as the number of nodes along the longest path from the root to that node.
 Definition IV.3 (Leaf and Non-leaf Specifications)
 A specification is termed as a leaf specification if the associated node in the graph ${\mathcal{G}}_{h}$ does not have any children; otherwise, it is referred to as a non-leaf specification.
-Lemma IV.4
-Any hierarchical sc-LTL specifications can be turned into a specification hierarchy graph.
+Let $\Phi_{\text{leaf}}$ denote the set of leaf specifications.
+Based on Def.
+IV.1, leaf specifications consist exclusively of atomic propositions, while non-leaf specifications consist solely of composite propositions.
 Example 1
-continued (Specification hierarchy graph)  The specification hierarchy graph is illustrated in Fig. 1(a). $\square$
-IV-B Semantics of Hierarchical sc-LTL
-Hierarchical sc-LTL cannot be interpreted by simply substituting composite propositions with their corresponding specifications.
+continued (Specification hierarchy tree)  The specification hierarchy tree is illustrated in Fig. 1(a). $\square$
+IV-B Semantics of H-LTLf
+H-LTLf cannot be interpreted by simply substituting composite propositions with their corresponding specifications.
 For example, replacing composite propositions in Eq.
-(2) at level $L_{1}$ generates a flat specification:
+(2) at level $L_{1}$ generates a standard specification:
 $\phi_{1}^{1}=\Diamond(\underbrace{\Diamond(s_{a}\wedge\Diamond t_{a})}_{\phi_{2}^{1}}\wedge\Diamond(\underbrace{\Diamond(s_{b}\wedge\Diamond t_{b})}_{\phi_{2}^{2}})\wedge\Diamond(\underbrace{\Diamond(s_{c}\wedge\Diamond t_{c})}_{\phi_{2}^{3}})),$
@@
 continued (Interpretation)  The way of interpretation is to start from the bottom level and progress upwards, verifying if the satisfaction of lower specifications allows the upper ones to move forward.
-We provide two separate cases.
-Case (i): If $s_{a}$ now holds true, then $\phi_{2}^{1}$ in (2) transitions to $\Diamond t_{a}$ based on the rules of progression [66] that generate sub-formulas to be satisfied, while $\phi_{1}^{1}$ remains unaffected as none of its composite propositions are fulfilled.
+We provide two separate cases for specifications in (2).
+Case (i): If $s_{a}$ now holds true, then $\phi_{2}^{1}$ transitions to $\Diamond t_{a}$ based on the rules of progression [65] that generate sub-formulas to be satisfied, while $\phi_{1}^{1}$ remains unaffected as none of its composite propositions are fulfilled.
 If $t_{a}$ subsequently holds true, $\phi_{2}^{1}$ is satisfied.
@@
 $\square$
-Input: Hierarchical sc-LTL specifications $\Phi$, state-specification plan $\tau$
-Output: True or False
-1 $\Phi^{acc}=\varnothing$ ;
-2 for $i\in\{0,\ldots,h\}$  do
-3       for $k\in\{K,\ldots,1\}$  do
-4             $\Phi^{acc,k}=\varnothing$ ;
-5             for $\phi\in\Phi^{k}\setminus\Phi^{acc}$  do
-6                   if $k=K$  then
-7                         $S(i)=\left\{s_{r}^{i}\,|\,(s_{r}^{i},\phi)\in\tau^{i}\right\}$ ;
-8                         if $S(i)=\varnothing$ then
-9                               continue;
-10
-11                         $\pi^{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\,|\,s_{r}^{i}\in S(i)\right\}$ ;
-12
+Given that H-LTLf includes multiple specifications, we consider multiple systems, particularly transition systems.
+At any given state, a system may be addressing a particular specification.
+We associate the state $s_{r}$ of a system ${\mathcal{T}}(r)$ with a leaf specification $\psi_{r}\in\Phi_{\text{leaf}}$.
+This association establishes a state-specification pair $(s_{r},\psi_{r})$, indicating that at state $s_{r}$, the system ${\mathcal{T}}(r)$ is engaged in satisfying the leaf specification $\psi_{r}$.
+Definition IV.4 (State-Specification Sequence)
+A state-specification sequence with a horizon $h$, represented as $\tau$, is a timed sequence $\tau=\tau_{0}\tau_{1}\tau_{2}\ldots\tau_{h}$.
+Here, $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),(s_{2}^{i},\psi_{2}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$ is the collective state-specification pairs of $N$ systems at the $i$-th timestep, where $s_{r}^{i}\in{\mathcal{S}}_{r}$, and $\psi_{r}^{i}\in\Phi_{\text{leaf}}\cup\{\epsilon\}$, with $\epsilon$ indicating the system’s non-involvement in any leaf specification at that time.
+Given the state-specification sequence $\tau$ and a leaf specification $\phi$, we generate a word as $\sigma=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$, where $\sigma_{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\ |\ (s_{r}^{i},\psi_{r}^{i})\in\tau_{i}\;\text{and}\;\psi_{r}^{i}=\phi\right\}$, represents the collective observations generated by the systems engaging in the leaf specification $\phi$ at instant $i$.
+Furthermore, given a leaf (non-leaf) specification $\phi$ and an input word $\sigma=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ with $\sigma_{i}\subseteq\texttt{Prop}(\phi)$, consisting of its atomic (composite) propositions, we construct an output word of $\phi$, denoted by $\sigma^{\prime}=\sigma^{\prime}_{0},\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{n}$, where $\sigma^{\prime}_{i}$ is either $\varnothing$ or $\{\phi\}$, tracking the satisfaction of $\phi$.
+Let $\sigma_{i,j}=\sigma_{i},\ldots,\sigma_{j}$ denote the segment of the input word between indices $i$ and $j$.
+Definition IV.5 (Output word)
+Given a specification $\phi$ and an input word $\sigma=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ with $\sigma_{i}\subseteq\texttt{Prop}(\phi)$, the output word $\sigma^{\prime}=\sigma^{\prime}_{0},\sigma_{1}^{\prime},\ldots,\sigma^{\prime}_{n}$ of $\phi$ is defined as:
+  • By default, $\sigma^{\prime}_{-1}=\{\phi\}$.
+  • $\sigma^{\prime}_{j}=\{\phi\}$ if $\sigma_{i+1,j},i+1\models\phi$, where $i$ (with $i<j$) is the most recent instant when $\phi$ was satisfied, i.e., $i=max\{k\,|\,\sigma^{\prime}_{k}=\{\phi\},k=-1,0,\ldots,j-1\}$.
+The satisfaction relation $\models$ follows the same definition as introduced in Section III, treating composite propositions as atomic propositions for non-leaf specifications.
+The progress of $\phi$ resets after it is satisfied at instant $i$, and it is considered satisfied again at instant $j$ if the input word segment $\sigma_{i+1,j}$ satisfies $\phi$.
+In the context of robots, this implies that the duration of task $\phi$ extends from $i+1$ to $j$, with the completion time at instant $j$.
+The task’s completion status is not persistent, meaning that once completed, it must start from scratch if repeated.
+The output words of specifications at level $L_{k+1}$ serve as the input words and produce output words of specifications at level $L_{k}$.
+Thus, starting from the words of leaf specifications generated by the state-specification sequence $\tau$ and proceeding in a bottom-up manner, we ultimately obtain an output word of the root specification $\phi_{1}^{1}$.
+The whole H-LTLf is satisfied by a state-specification sequence $\tau$ if the output word of $\phi_{1}^{1}$ contains $\phi_{1}^{1}$.
+Algorithmically, a post-order traversal can be used to traverse the specification hierarchy tree to compute the output word of $\phi_{1}^{1}$, as outlined in Alg.
+1.
+Definition IV.6 (Semantics)
+[…]
+Input: state-specification sequence $\tau$, LTLf  specification $\phi$, H-LTLf specifications $\Phi$
+Output: output word $\sigma^{\prime}$ of specification $\phi$
+  $\triangleright$ Get input word $\sigma$ [lines 1-1]
+  $\triangleright$ Leaf specification
+1 if $\phi\in\Phi_{\text{leaf}}$ then
+2       $\sigma(\phi)=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ where $\sigma_{i}=\left\{{\mathcal{L}}_{r}(s_{r}^{i})\ |\ (s_{r}^{i},\psi_{r}^{i})\in\tau_{i}\;\text{and}\;\psi_{r}^{i}=\phi\right\}$;
+3
+  $\triangleright$ Non-leaf specification
 4 else
-14                         $\pi^{i}=\Phi^{acc,k+1}$ ;
-15
-16                   ${\mathcal{Q}}_{\mathcal{B}}^{i+1}(\phi)=\left\{q^{\prime}\,|\,(q,\pi^{i},q^{\prime})\in\to_{\mathcal{B}}(\phi),\,\forall q\in{\mathcal{Q}}_{\mathcal{B}}^{i}(\phi),\right.$ $\left.\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{B}}(\phi)\right\}$ ;
-17                   if ${\mathcal{Q}}^{i+1}_{\mathcal{B}}(\phi)\cap{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi)\neq\varnothing$  then
-18                         $\Phi^{acc,k}=\Phi^{acc,k}\cup\{\phi\}$ ;
-19
-20            $\Phi^{acc}=\Phi^{acc}\cup\Phi^{acc,k}\,\cup\,\texttt{Children}(\Phi^{acc,k})$ ;
-21
-22      if $\phi_{1}^{1}\in\Phi^{acc}$ then
-23             return True ;
-24
-25return False ;
-Algorithm 1 Satisfaction check
-Given the presence of multiple specifications in hierarchical sc-LTL, a robot could be addressing any specification at a specific state.
-To clarify the robot’s intention and facilitate the checking of satisfaction, we associate the state $s_{r}$ of a robot $r$ with a leaf specification $\phi_{r}\in\Phi^{K}$.
-This association forms a state-specification pair $(s_{r},\phi_{r})$, signifying that the robot $r$ is engaged in fulfilling the specification $\phi_{r}$ at state $s_{r}$.
-Definition IV.5 (State-Specification Plan)
-A state-specification plan with a horizon $h$, represented as $\tau$, is a timed sequence $\tau=\tau^{0}\tau^{1}\tau^{2}\ldots\tau^{h}$.
-Here, $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ is the collective state-specification pairs of $N$ robots at the $i$-th timestep, where $s_{r}^{i}\in{\mathcal{S}}_{r}$, and $\phi_{r}^{i}\in\Phi^{K}\cup\{\epsilon\}$, with $\epsilon$ indicating the robot’s non-involvement in any leaf specification at that time.
-When considering the composite propositions in the context of long-horizon robot tasks, they correspond to sub-tasks.
-For example, one such task could be first clearing the table and then disposing of the trash, with each sub-task required to be completed exactly once.
-Another example involves retrieving a beverage from the refrigerator, such as a bottle of apple juice or orange juice, where completing just one sub-task suffices.
-Given this observation, we propose the following assumption:
-Assumption IV.6 (Composite proposition)
-For the successful completion of the robotic task, it is adequate that each composite proposition is satisfied no more than once.
-Given a robot state-specification plan, the method for determining the satisfaction of hierarchical sc-LTL specifications is shown in Alg.
-1.
-Because the truth of a non-leaf specification depends solely on the specifications at the immediately lower level, the algorithm iteratively assesses specifications in a bottom-up manner.
-Alg.
-1 will stop early and return true if the highest-level specification is met [lines 1-1].
-This is because that all specifications belong to co-safe LTL, the subsequent plan becomes irrelevant once the topmost specification is satisfied, eliminating the need for further satisfaction checks.
-Let $\Phi^{acc}$ represent the set of specifications that have been satisfied up to the current point [line 1].
-Additionally, let $\Phi^{acc,k}$ denote the set of specifications at level $L_{k}$ that have just become satisfied at the current time step $i$ [line 1].
-According to Assumption IV.6, each composite proposition is adequate to be satisfied no more than once, therefore, we iterate over each specification $\phi\in\Phi^{k}\setminus\Phi^{acc}$ at level $L_{k}$ that remain unsatisfied.
-Alg.
-1 computes the reachable set of automaton states for the next time step, designated as ${\mathcal{Q}}^{i+1}_{\mathcal{B}}(\phi)$, to evaluate the satisfaction [lines 1-1].
-This assessment varies depending on whether the specification $\phi$ is a leaf one or not.
-If the specification is a leaf specification, the algorithm aggregates the set of true propositions $\pi^{i}$, which are atomic propositions derived from the states of robots engaged in the specification $\phi$ [lines 1-1].
-Otherwise, if the specification is not a leaf, the set of true propositions is identified as the set $\Phi^{acc,k+1}$ – these are the specifications at the immediate lower level that have just been satisfied [line 1].
-The set $\Phi^{acc,k}$ gets updated if the reachable automaton states include at least one accepting state of $\phi$ [lines 1].
-Following the completion of iterations over level $L_{k}$, the set $\Phi^{acc}$ is then updated [line 1], where $\texttt{Children}(\Phi^{acc,k})$ denotes the set of specifications that are children nodes in the specification hierarchy graph of any specification in $\Phi^{acc,k}$, that is, the set of specifications for which a path exists from any specification in $\Phi^{acc,k}$.
-There is no need to verify the satisfaction of a specification if its parent specification has already been satisfied.
-For instance, there is no need to verify the satisfaction of $\phi_{2}^{1}$ for the specification $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2})$ if $\phi_{2}^{2}$ becomes true.
+5       for $\phi^{\prime}\in\texttt{Prop}(\phi)$ do
+6             $\sigma(\phi^{\prime})=\texttt{GenerateOutputWord}(\tau,\Phi,\phi^{\prime})$ ;
+7
+        $\triangleright$ Merge output words at the lower level as the input word of $\phi$
+8       $\sigma(\phi)=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$ where $\sigma_{i}=\left\{\sigma_{i}(\phi^{\prime})\;|\;\forall\phi^{\prime}\in\texttt{Prop}(\phi)\right\}$ ;
+9
+  $\triangleright$ Get output word $\sigma^{\prime}$ [lines 1-1]
+$i=-1$ ;
+  $\triangleright$ Latest time when $\phi$ is true
+10 for $j\in\{0,1,\ldots,n\}$ do
+11       if $\sigma_{i+1,j},i+1\models\phi$ then
+12             $\sigma^{\prime}_{j}=\{\phi\}$;
+13             $i=j$;
+14
+15      else
+16             $\sigma^{\prime}_{j}=\varnothing$ ;
+17
+18return $\sigma^{\prime}$;
+Algorithm 1 GenerateOutputWord($\tau,\Phi,\phi$)
+$t:$00111122223333444455556666777788889999obsv:$\{s_{a}\}$$\{t_{a}\}$$\{s_{b}\}$$\{t_{b}\}$$\{s_{c}\}$$\{t_{c}\}$ϕ21::superscriptsubscriptitalic-$\phi_{2}^{1}:$$\{\phi_{2}^{1}\}$ϕ22::superscriptsubscriptitalic-$\phi_{2}^{2}:$$\{\phi_{2}^{2}\}$ϕ23::superscriptsubscriptitalic-$\phi_{2}^{3}:$$\{\phi_{2}^{3}\}$ϕ11::superscriptsubscriptitalic-$\phi_{1}^{1}:$$\{\phi_{1}^{1}\}$
+Figure 2: Each row illustrates either the observation or the output word of a specification.
+The filled nodes indicate scenarios where an atomic or composite proposition is satisfied.
+The observations (obsv) generate the input word for leaf specifications $\phi_{2}^{1}\sim\phi_{2}^{3}$, while the combination of output words for specifications $\phi_{2}^{1}\sim\phi_{2}^{3}$ serve as the input word for the non-leaf specification $\phi_{1}^{1}$.
+The dashed arrows illustrate the correspondence between specific inputs that lead to the satisfaction of a given specification.
+Example 1
+continued (Semantics)  The output words for the specifications in (2) are illustrated in Fig. 2. $\square$
 Remark IV.7
-Every sc-LTL formula can be transformed into a Deterministic Finite Automaton (DFA).
-We opt for NBA, a form of Nondeterministic Finite Automaton (NFA), because sc-LTLs are transformable into either DFAs or NFAs, making both equally expressive for sc-LTL in this sense.
-The work from [8] and the temporal order heuristic by [21] that our approach in Sec.
-VI is based on, uses NFAs.
-Nonetheless, our approach can easily be adapted to work with DFAs as well, as graphically speaking, NFAs are more complex than DFAs.
-Remark IV.8
-The syntax and semantics for hierarchical sc-LTL here are designed for robotics and may not be applicable to other domains.
-Our experience indicates that a top-down reasoning approach is effective in constructing hierarchical sc-LTL specifications from task descriptions.
+The syntax and semantics for H-LTLf here are inspired by the domain of robots and may not be applicable to other domains.
+Our experience indicates that a top-down reasoning approach is effective in constructing H-LTLf specifications from task descriptions.
 This entails recognizing the inclusion relationships among tasks and organizing them into various levels of abstraction, ranging from general to specific.
 At each level, tasks that are closely related should be combined into single formulas, while those with weaker connections should be grouped into separate formulas.
+Remark IV.8
+The hierarchical format can be applied to syntactically co-safe LTL (sc-LTL), a syntactic subset of LTL, to formulate hierarchical sc-LTL.
+Sc-LTL formulas can be satisfied by finite sequences followed by any infinite repetitions.
 IV-C Expressiveness
@@
 A Kripke structure, denoted by ${\mathcal{K}}$, is deemed finite when it has a finite number of states.
-${\mathcal{K}}$ is considered a model for a temporal logic formula $\phi$ if every initial infinite path within ${\mathcal{K}}$ satisfies $\phi$.
+${\mathcal{K}}$ is considered a model for a temporal logic formula $\phi$ if every initial finite path within ${\mathcal{K}}$ satisfies $\phi$.
 Definition IV.9 (Expressive power [66])
@@
 Theorem IV.10 (Expressiveness)
-Hierarchical sc-LTL is more expressive than the flat sc-LTL.
+H-LTLf is more expressive than the standard LTLf.
 Proof:
-We first show that, for any flat sc-LTL, there exists an equivalent hierarchical counterpart.
-Therefore, the flat sc-LTL is subsumed by the hierarchical sc-LTL.
-Given a flat sc-LTL $\phi$, an equivalent two-level hierarchical sc-LTL can be constructed in a straightforward manner, where the top level is $\Diamond\phi$ and the bottom level is $\phi$ itself, that is,
+We first show that, for any standard LTLf, there exists an equivalent hierarchical counterpart.
+Therefore, the standard LTLf is subsumed by the H-LTLf.
+Given a standard LTLf $\phi$, an equivalent two-level H-LTLf can be constructed in a straightforward manner, where the top level is $\Diamond\phi$ and the bottom level is $\phi$ itself, that is,
 $L_{1}:\phi_{1}^{1}=\Diamond\phi_{2}^{1},\quad L_{2}:\phi_{2}^{1}=\phi.$
-Next, we introduce a counter-example to demonstrate the absence of the flat equivalent for certain hierarchical sc-LTL specifications.
-Considering a Kripke structure with two atomic propositions $\{a,b\}$, the following hierarchical sc-LTL specifications include two mutually exclusive leaf specifications:
+Next, we introduce a counter-example to demonstrate the absence of the standard equivalent for certain H-LTLf specifications.
+Considering a Kripke structure with two atomic propositions $\{a,b\}$, the following H-LTLf specifications include two mutually exclusive leaf specifications:
 $L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\,\Diamond\,\phi_{2}^{2}$
-$L_{2}:\quad$ $\phi_{2}^{1}=\neg b\,\mathcal{U}\,(a\wedge\neg b)$
-$\phi_{2}^{2}=\neg a\,\mathcal{U}\,(b\wedge\neg a).$
-While leaf specifications are mutually exclusive, the introduction of the state-specification plan enables the association of each state with one specification, facilitating the separate treatment of leaf specifications.
-For example, when these two specifications are assigned to separate robots, they can be realistically fulfilled.
-However, when solely using atomic propositions $a$ and $b$, combining two leaf specifications into a single sc-LTL specification is unfeasible, as it inevitably results in a contradiction, reducing the specification to False.
+$L_{2}:\quad$ $\phi_{2}^{1}=\neg b\,\mathcal{U}\,(a\wedge\neg b),\quad\phi_{2}^{2}=\neg a\,\mathcal{U}\,(b\wedge\neg a).$
+While leaf specifications are mutually exclusive, the introduction of the state-specification sequence enables the association of each state with one specification, facilitating the separate treatment of leaf specifications.
+For example, when these two specifications are associated with different Kripke structures, they can be realistically fulfilled.
+However, when solely using atomic propositions $a$ and $b$, combining two leaf specifications into an LTLf specification is unfeasible, as it inevitably results in a contradiction, reducing the specification to False.
 ∎
-Remark IV.11
-Hierarchical sc-LTL retains the expressive capability of sc-LTL for defining robot tasks, while also enhancing the ability to specify tasks for multiple robots.
-Moreover, as detailed in Section VIII-A, we conducted a user study that showed hierarchical sc-LTL is more user-friendly than flat sc-LTL.
 V Problem Formulation
-We first introduce a variant of the product team model defined in work [8], which aggregates the set of PBA for a given specification.
-Following this, we develop the hierarchical team model that integrates a set of product team models to address hierarchical sc-LTL specifications.
-V-A Product Team Model
+We first present an assumption on H-LTLf, and then introduce a variant of the product team model defined in work [9], which aggregates the set of PA for a given specification.
+Following this, we develop the hierarchical team model that integrates a set of product team models.
+V-A Assumptions on H-LTLf
+When considering the composite propositions in the context of robot tasks, they correspond to sub-tasks.
+For example, one such task could be retrieving a beverage from the refrigerator, such as a bottle of apple juice or orange juice, where completing just one sub-task suffices.
+Based on this observation, we assume that, for a task to be successfully completed, it is sufficient that each of its sub-tasks occurs at most once.
+This assumption applies only to non-leaf specifications.
+Assumption V.1 (Bounded One-Time Satisfaction)
+For any non-leaf specification $\phi_{k}^{i}$ at level $L_{k}$, there exists an accepting input word $\sigma=\sigma_{0},\sigma_{1},\dots,\sigma_{n}$ where $\sigma_{j}\subseteq\texttt{Prop}(\phi_{k}^{i})\subseteq\Phi_{k+1}$ for $j\in[n]$, such that $\sum_{j=0}^{n}\mathbb{I}(\phi_{k+1}^{l}\in\sigma_{j})\leq 1,$ for any $l\in[|\Phi_{k+1}|]_{+}$, where $\mathbb{I}$ is the indicator function.
+V-B Product Team Model
 Inspired by work [9], we make the following assumption.
 Assumption V.2 (Decomposition)
-Every leaf specification $\phi\in\Phi^{K}$ can be decomposed into a set of tasks such that
-  • Independence: every task can be executed independently;
+Every leaf specification $\phi\in\Phi_{\text{leaf}}$ can be decomposed into a set of tasks such that
+  • Independence: tasks can be executed independently;
   • Completeness: completion of all tasks, irrespective of the order, satisfies the leaf specification $\phi$.
 Definition V.3 (Decomposition Set [9])
-The decomposition set ${\mathcal{D}}_{\mathcal{B}}\subseteq{\mathcal{Q}}_{\mathcal{B}}$ of the NBA ${\mathcal{B}}$ contains all states $q$ such that each state leads to a decomposition of tasks.
+The decomposition set ${\mathcal{D}}_{\mathcal{A}}\subseteq{\mathcal{Q}}_{\mathcal{A}}$ of the NFA ${\mathcal{A}}$ contains all states $q$ such that each state leads to a decomposition of tasks.
 Example 2 (Decomposition Set)
-Consider an sc-LTL specification $\phi=\Diamond a\wedge\Diamond b$, with its corresponding NBA depicted in Fig.
+Consider an LTLf specification $\phi=\Diamond a\wedge\Diamond b$, with its corresponding NFA depicted in Fig.
 1(b).
@@
 Similarly, state 1 also belongs to the decomposition set.
-By default, all initial and accepting states in an NBA are considered decomposition states.
+By default, all initial and accepting states in an NFA are considered decomposition states.
 $\square$
+Definition V.4 (Product Team Model [9])
+Given a leaf specification $\phi\in\Phi_{\text{leaf}}$, the product team model ${\mathcal{P}}(\phi)$ consists of $N$ product models ${\mathcal{P}}(r,\phi)$ and is given by the tuple ${\mathcal{P}}(\phi):=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
+  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,s,q):r\in[N]_{+},(s,q)\in{\mathcal{Q}}_{\mathcal{P}}(r,\phi)\right\}$ is the set of state;
+  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:r=1,(s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(1,\phi)\right\}$ is the set of initial states;
+  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation.
+One type of transition, referred to as in-spec transitions, occurs inside a PA.
+That is, $(r,s,q)\to_{\mathcal{P}}(r,s^{\prime},q^{\prime})$ if $(s,q)\to_{{\mathcal{P}}(r,\phi)}(s^{\prime},q^{\prime})$; the other type of transition relation, denoted by $\zeta_{\text{in}}$, occurs between two PAs and is defined below in Def.
+V.5;
+  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)\}$ is the set of accepting/final states ;
+Definition V.5 (In-Spec Switch Transition (modified from [9]))
+Given a specification $\phi$, the set of in-spec switch transitions in ${\mathcal{P}}(\phi)$ is given by $\zeta_{\text{in}}\subset{\mathcal{Q}}_{\mathcal{P}}\times{\mathcal{Q}}_{\mathcal{P}}$.
+A transition $\left((r,s,q),(r^{\prime},s^{\prime},q^{\prime})\right)$ belongs to $\zeta_{\text{in}}$ if and only if it:
+  • points to the next robot, $r^{\prime}=r+1$;
+  • preserves the NFA progress, $q=q^{\prime}$;
+  • represents a decomposition choice, $q\in{\mathcal{D}}_{\mathcal{A}}(\phi)$.
 Figure 3: The product team model featuring two robots.
-PBAs ${\mathcal{P}}(1,\phi)$ and ${\mathcal{P}}(2,\phi)$ may be different reflecting the robots’ heterogeneity.
+PAs ${\mathcal{P}}(1,\phi)$ and ${\mathcal{P}}(2,\phi)$ may be different reflecting the robots’ heterogeneity.
 Initial states are indicated with incoming arrows, and accepting states are depicted with double circles.
@@
 A plan exclusive to robot 1 is traced by the green path, whereas the red path delineates a plan that involves the collaboration of both robots.
-Definition V.3 (Product Team Model (modified from [8]))
-Given a leaf specification $\phi\in\Phi^{K}$, the product team model ${\mathcal{P}}(\phi)$ consists of $N$ product models ${\mathcal{P}}(r,\phi)$ and is given by the tuple ${\mathcal{P}}(\phi):=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
-  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,s,q):r\in[N],(s,q)\in{\mathcal{Q}}_{\mathcal{P}}(r,\phi)\right\}$ is the set of state;
-  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:r=1,(s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(1,\phi)\right\}$ is the set of initial states ;
-  • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation.
-One type of transition, referred to as in-spec transitions, occurs inside a PBA.
-That is, $(r,s,q)\to_{\mathcal{P}}(r,s^{\prime},q^{\prime})$ if $(s,q)\to_{{\mathcal{P}}(r,\phi)}(s^{\prime},q^{\prime})$; the other type of transition relation, denoted by $\zeta_{\text{in}}$, occurs between two PBAs and is defined below in Def.
-V.4;
-  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)\}$ is the set of accepting/final states ;
-Definition V.4 (In-Spec Switch Transition (modified from [8]))
-Given a specification $\phi$, the set of in-spec switch transitions in ${\mathcal{P}}(\phi)$ is given by $\zeta_{\text{in}}\subset{\mathcal{Q}}_{\mathcal{P}}\times{\mathcal{Q}}_{\mathcal{P}}$.
-A transition $\left((r,s,q),(r^{\prime},s^{\prime},q^{\prime})\right)$ belongs to $\zeta_{\text{in}}$ if and only if it:
-  • points to the next robot, $r^{\prime}=r+1$;
-  • preserves the NBA progress, $q=q^{\prime}$;
-  • represents a decomposition choice, $q\in{\mathcal{D}}_{\mathcal{B}}(\phi)$.
 The distinction between in-spec transitions and in-spec switch transitions lies in their impact on task progress: in-spec transitions advance the task through the same robot, while in-spec switch transitions involve changing to a different robot without altering the task’s progress, as indicated by $q=q^{\prime}$.
 A visual representation of ${\mathcal{P}}(r,\phi)$ can be found in Fig.
-2.
+3.
 The connections between product models are unidirectional, starting from the first robot and ending at the last robot.
@@
 These action sequences can be executed in parallel as they collectively decompose the task, as illustrated in Fig.
-2.
+3.
 Remark V.6
@@
 The product team models for each leaf specification are depicted within areas shaded in light orange and green.
-To enhance clarity, some transitions within a single PBA are omitted.
+To enhance clarity, some transitions within a single PA are omitted.
 Inter-spec type I and II switch transitions are illustrated using blue and red dashed arrows, respectively.
@@
 Subsequently, robots 1 and 2 are tasked with satisfying specification $\phi_{2}$, connected through a type II inter-spec switch transition ② and involving another in-spec switch transition ③.
-V-B Hierarchical Team Models
-To extend to hierarchical sc-LTL, we construct a product team model per leaf specification.
+V-C Hierarchical Team Models
+To extend to H-LTLf, we construct a product team model per leaf specification.
 The goal here is to connect them to form a connectable space.
@@
 We note that in-spec dependence does not exist, as precluded by Asm.
-V.1.
+V.2.
 To capture inter-spec dependency and independence, we introduce two additional types of switch transitions between different product team models.
 Definition V.7 (Hierarchical Team Models)
-Given the hierarchical sc-LTL specifications $\Phi$, the hierarchical team models ${\mathcal{P}}$ consists of a set of product team models ${\mathcal{P}}(\phi)$ with $\phi\in\Phi^{K}$ and is denoted by the tuple ${\mathcal{P}}:=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
-  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,\phi,s,q):\phi\in\Phi^{K},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}(\phi)\right\}$ is the set of state;
-  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:\phi\in\Phi^{K},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(\phi)\right\}$ is the set of initial states ;
+Given the H-LTLf specifications $\Phi$, the hierarchical team models ${\mathcal{P}}$ consists of a set of product team models ${\mathcal{P}}(\phi)$ with $\phi\in\Phi_{\text{leaf}}$ and is denoted by the tuple ${\mathcal{P}}:=({\mathcal{Q}}_{{\mathcal{P}}},{\mathcal{Q}}^{0}_{\mathcal{P}},\to_{{\mathcal{P}}},{\mathcal{Q}}^{F}_{{\mathcal{P}}})$:
+  • ${\mathcal{Q}}_{{\mathcal{P}}}=\left\{(r,\phi,s,q):\phi\in\Phi_{\text{leaf}},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}(\phi)\right\}$ is the set of state;
+  • ${\mathcal{Q}}^{0}_{\mathcal{P}}=\left\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:\phi\in\Phi_{\text{leaf}},(r,s,q)\in{\mathcal{Q}}_{\mathcal{P}}^{0}(\phi)\right\}$ is the set of initial states;
   • $\to_{\mathcal{P}}\,\in{\mathcal{Q}}_{{\mathcal{P}}}\times{\mathcal{Q}}_{{\mathcal{P}}}$ is the transition relation.
 One type of transition occurs inside a product team model, that is, $(r,\phi,s,q)\to_{\mathcal{P}}(r^{\prime},\phi,s^{\prime},q^{\prime})$ if $(r,s,q)\to_{{\mathcal{P}}(\phi)}(r^{\prime},s^{\prime},q^{\prime})$; another two types of transition relation, denoted by $\zeta_{\text{inter}}^{1}$ and $\zeta_{\text{inter}}^{2}$, occur between two product teams models and are defined below in Def.
-V.7 and V.8;
-  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)\}$ is the set of accepting/final states.
+V.8 and V.9;
+  • ${\mathcal{Q}}^{F}_{{\mathcal{P}}}=\{(r,\phi,s,q)\in{\mathcal{Q}}_{\mathcal{P}}:q\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)\}$ is the set of accepting/final states.
 Definition V.8 (Inter-Spec Type I Switch Transition)
@@
   • points to the same robot state, $s=s^{\prime}$;
-  • connects decomposition states, $q\in{\mathcal{D}}_{\mathcal{B}}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})$.
+  • connects decomposition states, $q\in{\mathcal{D}}_{\mathcal{A}}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})$.
 Definition V.9 (Inter-Spec Type II Switch Transition)
@@
   • points to two different leaf specifications, $\phi\neq\phi^{\prime}$;
-  • connects an accepting state to a decomposition state, $q\in{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})$.
-Inter-spec type I switch transitions enable switching to another PBA with a different leaf specification, while retaining the same robot and pausing the current task.
-Conversely, inter-spec type II switch transitions create a connection from the accepting state of one PBA to the decomposition state of another PBA associated with the first robot, indicating the completion of one specification and the resumption of another.
+  • connects an accepting state to a decomposition state, $q\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi),q^{\prime}\in{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})$.
+Inter-spec type I switch transitions enable switching to another PA with a different leaf specification, while retaining the same robot and pausing the current task.
+Conversely, inter-spec type II switch transitions create a connection from the accepting state of one PA to the decomposition state of another PA associated with the first robot, indicating the completion of one specification and the resumption of another.
+It points toward the first robot, the starting robot of a product team model, because the in-spec switch transition is unidirectional, always pointing to the next robot.
 These hierarchical team models can be graphically represented in a 3D space, with each product team model occupying a 2D plane; see Fig.
-3.
-This structure has three axes: one representing the task progress for a specification (the $(s,q)$-axis), another indicating the task assignment among robots (the $r$-axis), and the last one tracking the overall progress of tasks across different specifications (the $\phi$-axis).
+4.
+This structure has three axes: one representing the task progress for a specification (the $(s,q)$-axis), another indicating the task assignment among robots (the $r$-axis), and the last one tracking the progress of different specifications (the $\phi$-axis).
 Remark V.10
-The construction is not a product of all NBA and robot states.
+The construction is not a product of all NFA and robot states.
 Instead, if each product team model is considered a distinct sub-space, the inter-spec switch transitions facilitate the connections between these sub-spaces.
@@
 Such a design significantly reduces the overall search space, enhancing the efficiency of the search.
-Given a state-specification plan $\tau=\tau^{0}\tau^{1}\ldots\tau^{h}$ where $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$, the cost for robot $r$, such as energy consumption or completion time, is represented as $c_{r}=\sum_{i=0}^{h-1}c_{r}(s^{i}_{r},s^{i+1}_{r})$, where $c_{r}(s^{i}_{r},s^{i+1}_{r})$ denotes the cost incurred transitioning between states for robot $r$.
+Given a state-specification sequence $\tau=\tau_{0}\tau_{1}\ldots\tau_{h}$ where $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),(s_{2}^{i},\psi_{2}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$, the cost for robot $r$, such as energy consumption or completion time, is represented as $c_{r}=\sum_{i=0}^{h-1}c_{r}(s^{i}_{r},s^{i+1}_{r})$, where $c_{r}(s^{i}_{r},s^{i+1}_{r})$ denotes the cost incurred transitioning between states for robot $r$.
 The goal is to minimize the additive cost, expressed as
@@
 Problem 1
-Given the hierarchical team models ${\mathcal{P}}$, derived from transition systems of robots and the NBAs of hierarchical sc-LTL specifications $\Phi$, find an optimal state-specification plan $\tau^{*}$ that minimizes $J(\tau^{*})$ and satisfies $\Phi$.
-VI Planning under Hierarchical sc-LTL
+Given the hierarchical team models ${\mathcal{P}}$, derived from transition systems and the H-LTLf specifications $\Phi$ that satisfy Asm.
+V.1, find an optimal state-specification sequence $\tau^{*}$ that satisfies $\Phi$ and minimizes $J(\tau^{*})$.
+VI Planning under H-LTLf
 This section begins by introducing a search-based approach, which constructs hierarchical team models on the fly, ensuring both completeness and optimality; see Sec.
@@
 2.
-The search state is defined as $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, where ${\mathbf{s}}\in{\mathcal{S}}_{1}\otimes\ldots\otimes{\mathcal{S}}_{N}$ and ${\mathbf{q}}={\mathcal{Q}}_{\mathcal{B}}(\phi_{K}^{1})\otimes\ldots\otimes{\mathcal{Q}}_{\mathcal{B}}(\phi_{1}^{1})$.
-This state is comprised of three parts: (i) $(r,\phi)$ indicates the specific PBA currently being searched; (ii) ${\mathbf{s}}$ represents the states of all robots; and (iii) ${\mathbf{q}}$ denotes the automaton states of all specifications (including leaf and non-leaf ones).
+The search state is defined as $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, where ${\mathbf{s}}\in{\mathcal{S}}_{1}\otimes\ldots\otimes{\mathcal{S}}_{N}$ and ${\mathbf{q}}={\mathcal{Q}}_{\mathcal{A}}(\phi_{K}^{1})\otimes\ldots\otimes{\mathcal{Q}}_{\mathcal{A}}(\phi_{1}^{1})$.
+This state is comprised of three parts: (i) $(r,\phi)$ indicates the specific PA currently being searched; (ii) ${\mathbf{s}}$ represents the states of all robots; and (iii) ${\mathbf{q}}$ denotes the automaton states of all specifications (including leaf and non-leaf ones).
 Two states are considered equivalent if they share the same $r$, $\phi$, ${\mathbf{s}}$, and ${\mathbf{q}}$ values.
@@
 A state $v=(r,\phi,{\mathbf{s}},{\mathbf{q}})$ is an initial state if
-  • $(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N],\forall\phi\in\Phi^{K}$;
-  • ${\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi^{K}$.
+  • $(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N]_{+},\forall\phi\in\Phi_{\text{leaf}}$;
+  • ${\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{A}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi_{\text{leaf}}$.
 That is, all robots begin at their initial states, and every specification starts from its initial automaton state [line 2].
@@
 The iteration process concludes when an accepting state of the topmost-level specification is reached.
-At this point, the function ExtractPlan is employed to parallelize the sequentially searched path, generating a state-specification plan [line 2].
+At this point, the function ExtractPlan is employed to parallelize the sequentially searched path, generating a state-specification sequence [line 2].
 If the accepting state is not yet reached, the algorithm proceeds to determine the successor states of $v$, along with their associated transition costs [line 2], through the function GetSucc, outlined in Alg.
@@
 Concurrently, the corresponding path leading to this state is recorded [lines 2-2].
-Input: Hierarchical sc-LTL specifications $\Phi$, a set of robots $[N]$
-Output: state-specification plan
+Input: H-LTLf specifications $\Phi$, a set of robots $[N]_{+}$
+Output: state-specification sequence
 1 ${\mathcal{V}}_{\text{seen}}=\varnothing,{\mathcal{V}}_{\text{explored}}=\varnothing,{\mathcal{H}}_{\text{front}}=\varnothing,\Lambda_{\text{explored}}=\varnothing$ ;
-2 ${\mathcal{V}}_{\text{init}}=\{v\,|\,(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N],\forall\phi\in\Phi^{K}\wedge{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi^{K}\}$ ;
+2 ${\mathcal{V}}_{\text{init}}=\{v\,|\,(r,\phi,{\mathbf{s}}(r),{\mathbf{q}}(\phi))\in{\mathcal{Q}}_{\mathcal{P}}^{0},\forall r\in[N]_{+},\forall\phi\in\Phi_{\text{leaf}}\wedge{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{A}}^{0}(\phi),\forall\phi\in\Phi\setminus\Phi_{\text{leaf}}\}$ ;
 3 ${\mathcal{H}}_{\text{front}}.\texttt{push}(({\mathcal{V}}_{\text{init}},0))$;
 4 $\tau_{\text{explored}}(v)=v,\forall v\in{\mathcal{V}}_{\text{init}}$;
-5 Build hierarchy specification graph ${\mathcal{G}}_{h}$;
+5 Build specification hierarchy tree ${\mathcal{G}}_{h}$;
 6 while ${\mathcal{H}}_{\text{front}}\neq\varnothing$ do
@@
 11      ${\mathcal{V}}_{\text{explored}}.\texttt{add}(v)$;
-12       if $v.{\mathbf{q}}(1,1)\in{\mathcal{Q}}_{\mathcal{B}}^{F}(1,1)$ then
+12       if $v.{\mathbf{q}}(\phi_{1}^{1})\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi_{1}^{1})$  then
 13             return $\texttt{ExtractPlan}(\Lambda_{\text{explored}}(v))$ ;
@@
 Algorithm 2 Search-based STAP
-Input: state $v$, hierarchical team models ${\mathcal{P}}$, hierarchy specification graph ${\mathcal{G}}_{h}$
+Input: state $v$, hierarchical team models ${\mathcal{P}}$, specification hierarchy tree ${\mathcal{G}}_{h}$
 Output: Set of successor states ${\mathcal{V}}_{\text{succ}}$
@@
 […]
-5 if $q\not\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)$ and $v.{\mathbf{q}}(\phi_{\text{prt}})\not\in{\mathcal{Q}}_{\mathcal{B}}^{F}(\phi_{\text{prt}}),\forall\phi_{\text{prt}}\in\Phi_{\text{prt}}$ then
+5 if $q\not\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)$ and $v.{\mathbf{q}}(\phi_{\text{prt}})\not\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi_{\text{prt}}),\forall\phi_{\text{prt}}\in\Phi_{\text{prt}}$ then
         $\triangleright$ In-spec transitions in ${\mathcal{P}}(r,\phi)$
@@
 8             $u.{\mathbf{s}}(r)=s^{\prime},\;u.{\mathbf{q}}(\phi)=q^{\prime}$ ;
-9             $\pi=\{\phi\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)$ else $\varnothing$ ;
+9             $\pi_{c}=\{\phi\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)$ else $\varnothing$ ;
 […]
@@
   $\triangleright$ Type I switch transitions in ${\mathcal{P}}$
-18 for $\phi^{\prime}\in\Phi^{K}\setminus\{\phi\}$  do
+18 for $\phi^{\prime}\in\Phi_{\text{leaf}}\setminus\{\phi\}$  do
 19       if $((r,\phi,s,q),(r,\phi^{\prime},s,v.{\mathbf{q}}(\phi^{\prime}))\in\zeta_{\text{inter}}^{1}$ then
@@
   $\triangleright$ Type II switch transitions in ${\mathcal{P}}$
-24 for $\phi^{\prime}\in\Phi^{K}\setminus\{\phi\}$  do
+24 for $\phi^{\prime}\in\Phi_{\text{leaf}}\setminus\{\phi\}$  do
 25       if $((r,\phi,s,q),(1,\phi^{\prime},v.{\mathbf{s}}(1),v.{\mathbf{q}}(\phi^{\prime}))\in\zeta_{\text{inter}}^{2}$ then
@@
 To identify in-spec transitions in ${\mathcal{P}}(r,\phi)$, the function initially checks if the specification $\phi$ or any of its non-leaf parent specifications are satisfied.
-This is achieved using Parents(${\mathcal{G}}_{h}$, $\phi$), which returns the path from the leaf specification $\phi$ to the topmost specification in the hierarchical specification graph ${\mathcal{G}}_{h}$, excluding $\phi$ itself [line 3].
+This is achieved using Parents(${\mathcal{G}}_{h}$, $\phi$), which returns the path from the leaf specification $\phi$ to the root specification in the specification hierarchy tree ${\mathcal{G}}_{h}$, excluding $\phi$ itself [line 3].
 If $\phi$ or any of its parent specifications are satisfied, the search in ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(\phi)$ ceases, as further exploration is unnecessary.
-This is based on condition (2) in Definition IV.1 and Assumption IV.6, which state that each leaf specification appears only in one non-leaf specification and satisfying each specification at most once is adequate.
-If a parent specification is satisfied, the leaf specification no longer contributes to task progress.
-Input: True composite propositions $\pi$, state $u$, parent specifications $\Phi_{\text{prt}}$
+This is based on condition (3) in Definition IV.1 and Assumption V.1, which state that each leaf specification appears only in one non-leaf specification and satisfying each specification at most once is adequate.
+Input: True composite propositions $\pi_{c}$, state $u$, parent specifications $\Phi_{\text{prt}}$
 Output: Updated set of states ${\mathcal{V}}_{u}$
 1 $\phi_{\text{prt}}=\Phi_{\text{prt}}.\texttt{pop}()$ ;
-2 ${\mathcal{Q}}_{\mathcal{B}}^{\text{succ}}=\left\{q^{\prime}\,|\,(u.{\mathbf{q}}(\phi_{\text{prt}}),\pi,q^{\prime})\in\to_{\mathcal{B}}(\phi_{\text{prt}}),\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{B}}(\phi_{\text{prt}})\right\}$;
-3 for $q^{\prime}\in{\mathcal{Q}}^{\text{succ}}_{\mathcal{B}}$ do
+2 ${\mathcal{Q}}_{\mathcal{A}}^{\text{succ}}=\left\{q^{\prime}\,|\,(u.{\mathbf{q}}(\phi_{\text{prt}}),\pi_{c},q^{\prime})\in\to_{\mathcal{A}}(\phi_{\text{prt}}),\forall q^{\prime}\in{\mathcal{Q}}_{\mathcal{A}}(\phi_{\text{prt}})\right\}$;
+3 for $q^{\prime}\in{\mathcal{Q}}^{\text{succ}}_{\mathcal{A}}$ do
 4       $u.{\mathbf{q}}(\phi_{\text{prt}})=q^{\prime}$ ;
@@
 8      else
-9             $\pi=\{\phi_{\text{prt}}\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi_{\text{prt}})$ else $\varnothing$;
+9             $\pi_{c}=\{\phi_{\text{prt}}\}$ if $q^{\prime}\in{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi_{\text{prt}})$ else $\varnothing$;
 […]
@@
 For the three types of switch transitions, only the robot $r$ (in the case of in-spec switch transitions [line 3)], the leaf specification $\phi$ (for inter-spec type I switch transitions [line 3]), or both robot $r$ and specification $\phi$ (for inter-spec type II switch transitions [line 3]) are updated.
-This implies a shift in the search space to a different PBA, without modifying $v.{\mathbf{s}}$ and $v.{\mathbf{q}}$.
+This implies a shift in the search space to a different PA, without modifying $v.{\mathbf{s}}$ and $v.{\mathbf{q}}$.
 Since these transitions do not change robot states and automaton states, the function UpdateNonLeafSpecs is not invoked.
@@
 The automaton state of $\phi_{\text{prt}}$ in $u.{\mathbf{q}}$ is then updated using these reachable states [line 4].
-This process continues until the automaton state of the topmost specification is updated [line 4].
+This process continues until the automaton state of the root specification is updated [line 4].
 Input: Searched path $\Lambda$
-Output: state-specification plan $\tau$
+Output: state-specification sequence $\tau$
 1 Remove switch states from $\Lambda$ and divide $\Lambda$ into $l$ path segments with each sharing the same specification;
-2 for $k\in[l]$ do
+2 for $k\in[l]_{+}$ do
 3       for $v\in\Lambda_{k}$ do
-4             $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),\ldots,(s_{r}^{i},\phi_{r}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ where $(s_{r}^{i},\phi_{r}^{i})=(v.{\mathbf{s}}(r),v.\phi)$ and $(s_{r^{\prime}}^{i},\phi_{r^{\prime}}^{i})=(v.{\mathbf{s}}(r^{\prime}),\epsilon),\forall r^{\prime}\not=v.r$ ;
+4             $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),\ldots,(s_{r}^{i},\psi_{r}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$ where $(s_{r}^{i},\psi_{r}^{i})=(v.{\mathbf{s}}(r),v.\phi)$ and $(s_{r^{\prime}}^{i},\psi_{r^{\prime}}^{i})=(v.{\mathbf{s}}(r^{\prime}),\epsilon),\forall r^{\prime}\not=v.r$ ;
 5
@@
 The ExtractPlan function in Alg.
-5 processes the path $\Lambda=v_{0}v_{1}\ldots v_{n}$, which leads to the state satisfying the topmost specification, and outputs a state-specification plan $\tau$.
+5 processes the path $\Lambda=v_{0}v_{1}\ldots v_{n}$, which leads to the state satisfying the root specification, and outputs a state-specification sequence $\tau$.
 This process is visually represented in Fig.
-4.
+5.
 A state $v$ within $\Lambda$ is identified as a switch state if it is the end state of an inter-spec switch transition $\zeta_{\text{inter}}^{1}\cup\zeta_{\text{inter}}^{2}$, which has the same robot state and automaton state as the corresponding start state.
@@
 Remark VI.2
+The search state includes the joint robot states to keep track of the states of idle robots, allowing the search to resume from their previous states when necessary.
+In Fig.
+4, after completing the search for robot 2 with respect to specification $\phi_{1}$, the search transitions to robot 1 while robot 2 remains idle.
+It then returns to robot 2 for specification $\phi_{2}$ via an in-spec switch transition ③.
+At this stage, knowing the state of robot 2 is essential.
+Remark VI.3
 Our method aligns with the decoupled approaches in the domain MAPF [67, 68], where initially, individual agents independently plan their paths without considering potential collisions.
 Subsequently, the system assesses for collisions and modifies the paths to prevent conflicts.
-Although this approach is more scalable and demands less computational effort than fully coupled methods, it may not ensure an optimal outcome, could lead to deadlocks, and might face challenges in dense areas.
 Resolving conflicts in temporal logic planning proves more difficult because, unlike MAPF, the actions of robots may be temporally correlated.
 When addressing collisions, it is crucial not only to consider robots that are physically close but also those that are temporally correlated.
-Therefore, we leave the systemic exploration of a complete collision resolution strategy to future studies.
+Moreover, the modified path must be carefully adjusted to ensure that the temporal logic task remains satisfied.
+Therefore, we leave the exploration of a collision resolution strategy to future studies.
 VI-B Search Acceleration with Heuristics
@@
 VI-B2 Essential Switch Transitions
-We introduce essential switch transitions to refine the connections between PBAs.
-Definition VI.3 (Essential State)
+We introduce essential switch transitions to refine the connections between PAs.
+Definition VI.4 (Essential State)
 An essential state for robot $r$, denoted as $(r,s)$, is defined either (a) $s=s_{r}^{0}$ or (b) there exists a leaf specification $\phi$ fulfilling the conditions:
@@
   • $q\neq q^{\prime}$;
-  • $q^{\prime}\in{\mathcal{D}}_{\mathcal{B}}(\phi)$.
+  • $q^{\prime}\in{\mathcal{D}}_{\mathcal{A}}(\phi)$.
 This definition implies that an essential state is either (a) the initial state of robot $r$ or (b) a state where robot $r$ causes an in-spec transition to a decomposition state of a leaf specification for the first time.
 In the latter case, the intuition is that essential states are pivotal for achieving substantial progress in the task.
-Definition VI.4 (Essential Switch Transition)
+Definition VI.5 (Essential Switch Transition)
 A switch transition $\zeta=\left((\phi,r,s,q),(\phi^{\prime},r^{\prime},s^{\prime},q^{\prime})\right)\in\zeta_{\text{in}}\cup\zeta_{\text{inter}}^{1}\cup\zeta_{\text{inter}}^{2}$ is an essential switch transition if both $(r,s)$ and $(r^{\prime},s^{\prime})$ are essential states.
@@
 For a state $u=(r,\phi,{\mathbf{s}},{\mathbf{q}})$, we define its task-level cost as summation over all leaf specifications:
-$c_{q}({\mathbf{q}})&=\sum_{\phi^{\prime}\in\Phi^{K}}c_{q}({\mathbf{q}}(\phi^{\prime}))\\ =\sum_{\phi^{\prime}\in\Phi^{K}}&\max\left\{\texttt{len}(q^{0}_{\mathcal{B}}\to{\mathbf{q}}(\phi^{\prime}))\,|\,\forall\,q^{0}_{\mathcal{B}}\in{\mathcal{Q}}_{\mathcal{B}}^{0}(\phi^{\prime})\right\}.\end{aligned}$ (4)
-The cost $c_{q}({\mathbf{q}}(\phi^{\prime}))$ is determined by the length of the simplest path from any initial state $q^{0}_{\mathcal{B}}$ to the automaton state ${\mathbf{q}}(\phi^{\prime})$ within the NBA ${\mathcal{B}}(\phi)$.
+$c_{q}({\mathbf{q}})&=\sum_{\phi^{\prime}\in\Phi_{\text{leaf}}}c_{q}({\mathbf{q}}(\phi^{\prime}))\\ =\sum_{\phi^{\prime}\in\Phi_{\text{leaf}}}&\max\left\{\texttt{len}(q^{0}_{\mathcal{A}}\to{\mathbf{q}}(\phi^{\prime}))\,|\,\forall\,q^{0}_{\mathcal{A}}\in{\mathcal{Q}}_{\mathcal{A}}^{0}(\phi^{\prime})\right\}.\end{aligned}$ (5)
+The cost $c_{q}({\mathbf{q}}(\phi^{\prime}))$ is determined by the length of the simplest path from any initial state $q^{0}_{\mathcal{A}}$ to the automaton state ${\mathbf{q}}(\phi^{\prime})$ within the NFA ${\mathcal{A}}(\phi)$.
 A simple path is defined as one that does not include any repeated states, and the function len measures the path’s length in terms of the number of transitions.
@@
 where $c_{r}$ is the cost between robot states, $c_{q}(u.{\mathbf{q}})-c_{q}(v.{\mathbf{q}})$ is the cost between automaton states and $w$ is the weight.
-Remark VI.5
+Remark VI.6
 The first two heuristics essentially alter the structure of the hierarchical team models.
@@
 However, employing exclusively the third heuristic still maintains the completeness of the search, as it ensures the full exploration of the search space eventually.
-Additionally, this approach with heuristic of automaton state remains optimal with respect to the newly defined cost (5).
 Extensive simulations presented in Sec.
@@
 The rationale behind this empirical completeness is twofold.
-Firstly, [28] prove that the inference on temporal order is correct for a broad range of flat sc-LTL specifications.
+Firstly, [29] prove that the inference on temporal order is correct for a broad range of standard LTL specifications.
 By eliminating switch transitions that conflict with precedence orders, transitions that are in accordance with temporal orders remain unaffected.
@@
 By default, these are recognized as decomposition states.
-When a PBA ${\mathcal{P}}(r,\phi)$ is chosen for exploration, only the corresponding automaton state ${\mathbf{q}}(\phi)$ is updated, leaving the automaton states from other leaf specifications unaffected.
+When a PA ${\mathcal{P}}(r,\phi)$ is chosen for exploration, only the corresponding automaton state ${\mathbf{q}}(\phi)$ is updated, leaving the automaton states from other leaf specifications unaffected.
 The search can exit ${\mathcal{P}}(r,\phi)$ only through three types of switch transitions, each, by definition, associated with decomposition states.
-Upon switching to a different PBA ${\mathcal{P}}(r^{\prime},\phi^{\prime})$, the automaton state of $\phi$, which is a decomposition state, remains unchanged.
+Upon switching to a different PA ${\mathcal{P}}(r^{\prime},\phi^{\prime})$, the automaton state of $\phi$, which is a decomposition state, remains unchanged.
 Therefore, by inductive reasoning, it follows that at any given point, at most one automaton state from the leaf specifications does not belong to the decomposition set.
@@
 The total number of transitions in the hierarchical team models ${\mathcal{P}}$ is given as follows:
-$|\to_{\mathcal{P}}|=\sum_{\phi\in\Phi^{K}}\sum_{r\in[N]}\left|\to_{{\mathcal{P}}(r,\phi)}\right|+\underbrace{|\zeta_{\text{in}}|+|\zeta_{\text{inter}}^{1}|+|\zeta_{\text{inter}}^{2}|}_{|\zeta|},$ (6)
+$|\to_{\mathcal{P}}|=\sum_{\phi\in\Phi_{\text{leaf}}}\sum_{r\in[N]_{+}}\left|\to_{{\mathcal{P}}(r,\phi)}\right|+\underbrace{|\zeta_{\text{in}}|+|\zeta_{\text{inter}}^{1}|+|\zeta_{\text{inter}}^{2}|}_{|\zeta|},$ (7)
 which is the total number of in-spec transitions and switch transitions.
-According to [8], $|\to_{{\mathcal{P}}(r,\phi)}|$ is bounded by $|\to_{{\mathcal{P}}(r,\phi)}|\leq\left(|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right).$ Furthermore, $|\zeta_{\text{in}}|$, $|\zeta_{\text{inter}}^{1}|$ and $|\zeta_{\text{inter}}^{2}|$ are, respectively, upper bounded by
-$|\zeta_{\text{in}}|$ $\leq\sum_{\phi\in\Phi^{K}}\sum_{r\in[N-1]}\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r+1}|\right),$
-$|\zeta_{\text{inter}}^{1}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi^{K}}\sum_{r\in[N]}\left(|{\mathcal{D}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{r}|\right),$ (7)
-$|\zeta_{\text{inter}}^{2}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi^{K}}\sum_{r\in[N]}\left(|{\mathcal{Q}}^{F}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{B}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{1}|\right).$
+According to [9], $|\to_{{\mathcal{P}}(r,\phi)}|$ is bounded by $|\to_{{\mathcal{P}}(r,\phi)}|\leq\left(|{\mathcal{Q}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{Q}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right).$ Furthermore, $|\zeta_{\text{in}}|$, $|\zeta_{\text{inter}}^{1}|$ and $|\zeta_{\text{inter}}^{2}|$ are, respectively, upper bounded by
+$|\zeta_{\text{in}}|$ $\leq\sum_{\phi\in\Phi_{\text{leaf}}}\sum_{r\in[N-1]_{+}}\left(|{\mathcal{D}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r+1}|\right),$
+$|\zeta_{\text{inter}}^{1}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi_{\text{leaf}}}\sum_{r\in[N]_{+}}\left(|{\mathcal{D}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{r}|\right),$ (8)
+$|\zeta_{\text{inter}}^{2}|$ $\leq\sum_{\phi,\phi^{\prime}\in\Phi_{\text{leaf}}}\sum_{r\in[N]_{+}}\left(|{\mathcal{Q}}^{F}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|\right)\left(|{\mathcal{D}}_{\mathcal{A}}(\phi^{\prime})|\cdot|{\mathcal{S}}_{1}|\right).$
 Proof:
-For a given PBA ${\mathcal{P}}(r,\phi)$, the number of states is upper bounded by $|{\mathcal{Q}}_{\mathcal{B}}(\phi)|\cdot|{\mathcal{S}}_{r}|$.
-Its squared term gives the upper bound for the number of in-spec transitions $|\to_{{\mathcal{P}}(r,\phi)}|$, since any pair of states within the PBA can potentially be connected.
-Regarding in-spec switch transitions between two consecutive PBAs, ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(r+1,\phi)$, each decomposition state can be associated with any robot state, which sets the upper bound on the number of possible starting states to $|{\mathcal{D}}_{\mathcal{B}}(\phi)|\times|{\mathcal{S}}_{r}|$.
-Similarly, the maximum number of end states is determined by $|{\mathcal{D}}_{\mathcal{B}}(\phi)|\times|{\mathcal{S}}_{r+1}|$, subsequently defining the upper bound for $|\zeta_{\text{in}}|$.
+For a given PA ${\mathcal{P}}(r,\phi)$, the number of states is upper bounded by $|{\mathcal{Q}}_{\mathcal{A}}(\phi)|\cdot|{\mathcal{S}}_{r}|$.
+Its squared term gives the upper bound for the number of in-spec transitions $|\to_{{\mathcal{P}}(r,\phi)}|$, since any pair of states within the PA can potentially be connected.
+Regarding in-spec switch transitions between two consecutive PAs, ${\mathcal{P}}(r,\phi)$ and ${\mathcal{P}}(r+1,\phi)$, each decomposition state can be associated with any robot state, which sets the upper bound on the number of possible starting states to $|{\mathcal{D}}_{\mathcal{A}}(\phi)|\times|{\mathcal{S}}_{r}|$.
+Similarly, the maximum number of end states is determined by $|{\mathcal{D}}_{\mathcal{A}}(\phi)|\times|{\mathcal{S}}_{r+1}|$, subsequently defining the upper bound for $|\zeta_{\text{in}}|$.
 Following the same logic, the upper bound for inter-spec switch transitions can be deduced as outlined in Eq.
@@
 Theorem VII.4 (Soundness)
-The state-specification plan given by Alg. 2 satisfies hierarchical sc-LTL specifications $\Phi$.
+The state-specification sequence given by Alg. 2 satisfies H-LTLf specifications $\Phi$.
 Proof:
 The soundness of Alg.
-2 stems directly from its correct-by-construction design.
-To prove it, we use Alg.
-1 to validate the correctness of the returned state-specification plan $\tau$.
-This plan $\tau$ is divided into segments $\tau=\tau_{1}\tau_{2}\ldots\tau_{l}$, with each $\tau_{k}$ aligning with a corresponding path segment $\Lambda_{k}$ from the searched path $\Lambda$.
-Beginning with the first segment $\tau_{1}$, and as the algorithm iterates over each $\phi\in\Phi^{K}\setminus\Phi^{acc}$ in line 1, the loop continues as long as $\phi\not=\phi_{1}$, where $\phi_{1}$ is the leaf specification associated with $\tau_{1}$.
-This is because $S(i)=\varnothing$ in line 1 for any $\phi\not=\phi_{1}$.
-When $\phi=\phi_{1}$, the set of reachable states ${\mathcal{Q}}_{\mathcal{B}}^{i+1}$ in NBA ${\mathcal{B}}(\phi_{1})$ are updated accordingly.
-Consequently, progress is only made for $\phi_{1}$, while all other leaf specifications remain idle.
-Upon completing the iteration over $\tau_{1}$ segment, $\phi_{1}$ reaches a decomposition state in NBA ${\mathcal{B}}(\phi_{1})$, as per the design of the hierarchical team models.
-In line 1 of Alg.
-1, the non-leaf parent specifications of $\phi_{1}$ are updated based on the progress made solely in $\phi_{1}$.
-This update mirrors the operations performed by the UpdateNonLeafSpecs function in Alg.
+2 follows directly from its correct-by-construction design.
+The correctness of the state-specification sequence aligns with the correctness of the computed path $\Lambda=v_{0},v_{1},\ldots,v_{n}$.
+Each state $v\in\Lambda$ maintains its specification-related information through its component $v.{\mathbf{q}}$, which tracks the automaton states for all specifications.
+Consequently, for each specification $\phi\in\Phi$, we can derive an output word $\sigma(\phi)=\sigma_{0},\sigma_{1},\ldots,\sigma_{n}$, where $\sigma_{i}=\{\phi\}$ if $v_{i}.{\mathbf{q}}(\phi)\in{\mathcal{Q}}_{\mathcal{A}}^{F}(\phi)$ and $\sigma_{i}=\varnothing$ otherwise.
+Output words at lower levels propagate upward, generating those at the next higher level according to the bottom-up updates of non-leaf specifications in Alg.
 4.
-The same procedure is applied to the subsequent segments $\tau_{2},\tau_{3},\ldots$, with each segment facilitating progress on a single leaf specification while pausing others.
-When an accepting state for a particular specification is reached, all PBAs associated with its child leaf specifications cease to be searched, as specified in Alg.
-3, line 3.
-This is in alignment with the process in Alg.
-1, line 1, where child leaf specifications of a satisfied parent specification are not further examined.
-Ultimately, upon reaching the end of $\tau_{l}$, the topmost specification is fulfilled, as noted in Alg.
-2, line 2.
-This completion signifies that the hierarchical sc-LTL specifications are satisfied in accordance with Alg.
-1, line 1, thereby confirming the correctness of the entire process.
-∎
+Ultimately, the output word for the root specification $\phi_{1}^{1}$ includes $\phi_{1}^{1}$ itself, which meets the termination condition of Alg.
+2 at line 2.
+As a result, the approach conforms to the semantic definition in Def.
+IV.6.
 Assumption VII.5 (Segmented Plan)
-Assuming that there exists a feasible state-specification plan $\tau=\tau^{0}\tau^{1}\ldots\tau^{h}$ that can be sequentially divided into $\ell$ segments with $\ell\leq|\Phi^{K}|$, represented as $\tau=\langle\tau\rangle_{1}\langle\tau\rangle_{2}\ldots\langle\tau\rangle_{\ell}$ where no two segments share the same specification.
+Assuming that there exists a feasible state-specification sequence $\tau=\tau_{0}\tau_{1}\ldots\tau_{h}$ that can be sequentially divided into $\ell$ segments with $\ell\leq|\Phi_{\text{leaf}}|$, represented as $\tau=\langle\tau\rangle_{1}\langle\tau\rangle_{2}\ldots\langle\tau\rangle_{\ell}$ where no two segments share the same specification.
 Each segment $\langle\tau\rangle_{l}$ fulfills a single leaf specification.
-That is, for every aggregated state-specification pair $\tau^{i}=((s_{1}^{i},\phi_{1}^{i}),(s_{2}^{i},\phi_{2}^{i}),\ldots,(s_{N}^{i},\phi_{N}^{i}))$ in the segment $\langle\tau\rangle_{l}=\tau^{m}\tau^{m+1}\ldots\tau^{n}$, each $\phi^{i}_{j}$ is either the same leaf specification $\phi$ that $\langle\tau\rangle_{l}$ satisfies, or it is the null specification $\epsilon$.
-Note that the number of segments $\ell$ might be less than the number of leaf specifications $|\Phi^{K}|$, since it is possible that not all leaf specifications are required to be satisfied, such as leaf specifications that are connected by disjunction.
+That is, for every aggregated state-specification pair $\tau_{i}=((s_{1}^{i},\psi_{1}^{i}),(s_{2}^{i},\psi_{2}^{i}),\ldots,(s_{N}^{i},\psi_{N}^{i}))$ in the segment $\langle\tau\rangle_{l}=\tau_{m}\tau_{m+1}\ldots\tau_{n}$, each $\psi^{i}_{r}$ is either the same leaf specification $\phi$ that $\langle\tau\rangle_{l}$ satisfies, or it is the null specification $\epsilon$.
+Note that the number of segments $\ell$ might be less than the number of leaf specifications $|\Phi_{\text{leaf}}|$, since it is possible that not all leaf specifications are required to be satisfied, such as leaf specifications that are connected by disjunction.
 The following properties only apply to Alg.
@@
 Theorem VII.6 (Completeness)
-Assuming that a state-specification plan $\tau$ satisfying Asm.
-VII.5 exists, Alg.
-2 is complete in that it is guaranteed to return a feasible plan that fulfills the hierarchical sc-LTL specifications $\Phi$.
+Assuming that a state-specification sequence $\tau$ satisfying Asm. VII.5 exists, Alg. 2 is complete in that it is guaranteed to return a feasible plan that fulfills the H-LTLf specifications $\Phi$.
 Proof:
@@
 In this case, only inter-spec type II switch transitions exist between team models of any two leaf specifications $\phi$ and $\phi^{\prime}$.
-Type II transitions connect accepting states of every PBA within the product team model of $\phi$ (or $\phi^{\prime}$) to decomposition states of the first PBA of the team model of $\phi^{\prime}$ (or $\phi$).
+Type II transitions connect accepting states of every PA within the product team model of $\phi$ (or $\phi^{\prime}$) to decomposition states of the first PA of the team model of $\phi^{\prime}$ (or $\phi$).
 As a result, the search does not advance to product team models of other leaf specifications until it discovers a path segment that satisfies the current one, e.g., $\phi$.
-Furthermore, once an accepting state of a PBA for $\phi$ is reached, the search only moves to initial states, not other decomposition states, of the first PBA in another team model, e.g., $\phi^{\prime}$, as the search on this first PBA has not started yet.
+Furthermore, once an accepting state of a PA for $\phi$ is reached, the search only moves to initial states, not other decomposition states, of the first PA in another team model, e.g., $\phi^{\prime}$, as the search on this first PA has not started yet.
 Following this logic, Alg.
@@
 Theorem VII.7 (Optimality)
-Assuming the existence of a state-specification plan $\tau^{*}$ that meets Asm.
+Assuming the existence of a state-specification sequence $\tau^{*}$ that meets Asm.
 VII.5 with least cost, Alg.
@@
 VIII Simulation Experiments
+We use Python 3.10.12 on a computer with 3.5 GHz Apple M2 Pro and 16G RAM.
+Our code can be available via the https://github.com/XushengLuo92/Hierarchical-LTL-STAP.
+The video is accessible via the link https://www.youtube.com/watch?v=N3f8pUHDPF4&t=4s.
+VIII-A User Study on Understanding H-LTLf
+[…]
 Figure 6: Pick-and-place platform for user study.
-We use Python 3.10.12 on a computer with 3.5 GHz Apple M2 Pro and 16G RAM. Our code can be available via the https://github.com/XushengLuo92/Hierarchical-LTL-STAP.
-VIII-A User Study on Understanding Hierarchical sc-LTL
-[…]
 VIII-A1 Survey
@@
 The subjective section consists of 7 questions where participants are presented with natural language instructions for 7 tasks.
-For each task, we provide a flat sc-LTL formula and hierarchical formulas.
-The participants are asked to choose which formula is easier to understand and clearer in representing the task described, of which task 5 is, “First place the circle, and then place diamond and star in any order, and lastly place the heart.” The flat sc-LTL is
+For each task, we provide a standard LTLf formula and hierarchical formulas.
+The participants are asked to choose which formula is easier to understand and clearer in representing the task described, of which task 5 is, “First place the circle, and then place diamond and star in any order, and lastly place the heart.” The standard LTLf is
 $&\Diamond(s_{\text{cc}}\wedge\Diamond t_{\text{cc}})\wedge\Diamond(s_{\text{dm}}\wedge\Diamond t_{\text{dm}})\wedge\Diamond(s_{\text{st}}\wedge\Diamond t_{\text{st}})\wedge\Diamond(s_{\text{ht}}\wedge\Diamond t_{\text{ht}})\\ &\wedge\neg s_{\text{dm}}\mathcal{U}t_{\text{cc}}\wedge\neg s_{\text{dm}}\mathcal{U}t_{\text{cc}}\wedge\neg s_{\text{st}}\mathcal{U}t_{\text{cc}}\wedge\neg s_{\text{ht}}\mathcal{U}t_{\text{dm}}\wedge\neg s_{\text{ht}}\mathcal{U}t_{\text{st}}\end{aligned}$
-and the hierarchical sc-LTL is
+and the H-LTLf specifications are
 $L_{1}:\quad&\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond(\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}))\\ L_{2}:\quad&\phi_{2}^{1}=\Diamond(s_{\text{cc}}\wedge\Diamond t_{\text{cc}})\\ &\phi_{2}^{2}=\Diamond(s_{\text{dm}}\wedge\Diamond t_{\text{dm}})\wedge\Diamond(s_{\text{st}}\wedge\Diamond t_{\text{st}})\\ &\phi_{2}^{3}=\Diamond(s_{\text{ht}}\wedge\Diamond t_{\text{ht}})\end{aligned}$ (9)
-For the objective section, we prepare 4 pairs of flat and hierarchical formulas, each pair corresponding to the same task.
-This results in a total of 8 questions, with each one featuring either a flat sc-LTL formula or hierarchical sc-LTL formulas.
+Note that standard LTLf and H-LTLf formulas are not equivalent, as achieving such equivalence is nontrivial.
+However, both formalisms are capable of capturing task descriptions to a large extent.
+For the objective section, we prepare 4 pairs of standard and hierarchical formulas, each pair corresponding to the same task.
+This results in a total of 8 questions, with each one featuring either a standard LTLf formula or H-LTLf formulas.
 We randomly order the formulas so that adjacent questions do not necessarily involve the same task.
@@
 This results in 20 valid surveys.
-For the subjective questions, we compute the percentage of participants who prefer flat versus hierarchical formulas across 140 instances (20 participants, 7 questions each).
-For the objective questions, we calculate the accuracy for each type of sc-LTL across 80 instances per sc-LTL type (20 participants, 4 questions per sc-LTL type).
+For the subjective questions, we compute the percentage of participants who prefer standard versus hierarchical formulas across 140 instances (20 participants, 7 questions each).
+For the objective questions, we calculate the accuracy for each type of LTLf across 80 instances per LTLf type (20 participants, 4 questions per LTLf type).
 We track the time taken by each participant to respond to each question.
-A detailed analysis of the survey results is provided below.
-For the subjective questions, 27.1% of all responses show a preference for flat sc-LTL, while 72.9% favor hierarchical sc-LTL.
-A detailed count of preferences for each task is illustrated in Fig.
-6, where respondents showed a clear preference over hierarchical sc-LTL except for task 4.
+For the subjective questions, 27.1% of all responses show a preference for standard LTLf, while 72.9% favor H-LTLf.
+A detailed count of preferences is illustrated in Fig.
+7, where respondents showed a clear preference over H-LTLf except for task 4.
 The specification is that “Place the circle, diamond, star and heart.
-Diamond should be after circle and heart should be after star.” The temporal constraints are simpler compared to those in task 5 in (8), making flat sc-LTL adequate for this scenario.
-Regarding the objective questions, the accuracy rates are 65.0% for flat sc-LTL and 86.3% for hierarchical sc-LTL.
-Additionally, the average response time for questions involving flat sc-LTL formulas is 78.3 seconds, compared to 47.1 seconds for hierarchical sc-LTL formulas.
-These findings further validate the user-friendliness of hierarchical sc-LTL, as evidenced by both higher accuracy and reduced response times.
+Diamond should be after circle and heart should be after star.” The temporal constraints are simpler compared to those in task 5 in (9), making standard LTLf adequate for this scenario.
+Regarding the objective questions, the accuracy rates are 65.0% for standard LTLf and 86.3% for H-LTLf.
+Additionally, the average response time for questions involving standard LTLf formulas is 78.3 seconds, compared to 47.1 seconds for H-LTLf formulas.
+These findings further validate the user-friendliness of H-LTLf, as evidenced by both higher accuracy and reduced response times.
 Details are presented in Fig.
-7, where respondents achieved higher accuracy rates in all tasks specified under hierarchical sc-LTL.
+8, where respondents achieved higher accuracy rates in all tasks specified under H-LTLf.
 Figure 8: The accuracy and response time per task.
@@
 We use the office environment described in [9] for service tasks, as depicted in Fig.
-8.
+9.
 This 30×$\times 7$ grid map features 14 desk areas, 6 meeting rooms, and several other functional rooms.
-Additionally, we distribute $m$ robots at various locations within this simulated environment.
+Additionally, we distribute $m$ robots at various locations.
 The first three scenarios are derived from the work [9].
-We outline the hierarchical specifications for these scenarios below, while the corresponding flat versions of these specifications are referred to [8].
-Note that the hierarchical specifications were derived directly from task descriptions, rather than from flat specifications.
-In Def.
-IV.1, we limit the sc-LTL specifications to co-safe LTL.
-However, all hierarchical sc-LTL specifications below contain a sub-formula expressed as $\square(a\Rightarrow b)$.
-Even with this sub-formula, it is still sufficient to assert satisfaction upon reaching the accepting states.
+We outline the hierarchical specifications for these scenarios below, while the corresponding standard versions of these specifications are referred to [9].
+Note that the hierarchical specifications were derived directly from task descriptions, rather than from standard LTLf specifications.
 VIII-B1 Scenario 1
-Empty a paper bin located at desk $d_{5}$. The robot must avoid the public area while transporting the bin.
+Transport the paper bin from desk $d_{5}$ to area $g$ for emptying, avoiding the public area while it is full.
+Return an empty bin from $g$ to desk $d_{5}$.
+The atomic propositions are: $\mathsf{default}$: the robot is not carrying any object, $\mathsf{carrybin}$: the robot is carrying a full paper bin, $\mathsf{dispose}$: the robot is disposing of garbage, $\mathsf{emptybin}$: the robot is carrying an empty paper bin, and $\mathsf{public}$: the robot is located in a public area.
 $L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}$
@@
 $\quad\quad\quad\quad\wedge\square(\mathsf{carrybin\Rightarrow\neg\mathsf{public}})$
-$\phi_{2}^{2}=\Diamond(d_{5}\wedge\mathsf{emptybin}\wedge\bigcirc(d_{5}\wedge\mathsf{default}))$
+$\phi_{2}^{2}=\Diamond(g\wedge\bigcirc(g\wedge\mathsf{emptybin})\wedge\Diamond(d_{5}\wedge\bigcirc(d_{5}\wedge\mathsf{default})))$
 Figure 9: The office building in a grid-based layout, where areas $d_{1}$ to $d_{14}$ represent desks, $m_{1}$ to $m_{6}$ are meeting rooms, $e$ stands for the elevator, $g$ for the garbage room, $p$ for the printer room, and $k$ for the coffee kitchen.
@@
 VIII-B2 Scenario 2
-Distribute documents to desks $d_{10}$, $d_{7}$, and $d_{5}$, and avoid public areas while carrying the document.
+Distribute documents to desks $d_{10}$, $d_{7}$, and $d_{5}$, and avoid public areas while carrying the document. Let $\mathsf{carry}$ denote the action of the robot carrying the document.
 $L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$
@@
 Guide a person waiting at desk $d_{11}$ to meeting room $m_{6}$.
+Let $\mathsf{guide}$, $\mathsf{photo}$, and $\mathsf{camera}$ denote the actions of the robot guiding a person, capturing a photo, and activating the camera, respectively.
 $L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$
@@
 VIII-B4 Combinations of scenarios 1, 2 and 3
-We examine combinations of any two of these tasks as well as the combination of all three. Due to space constraints, we only detail the scenario involving the combination of all three tasks below.
+We examine combinations of any two of these tasks as well as the combination of all three. Due to space constraints, we only detail the scenario involving the final occurrence of all three tasks.
 $L_{1}:\quad$ $\phi_{1}^{1}=\Diamond\phi_{2}^{1}\wedge\Diamond\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}$
-$L_{2}:\quad$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{1}=\Diamond\phi_{3}^{1}\wedge\Diamond\phi_{3}^{2}}$
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{2}=\Diamond\phi_{3}^{3}\wedge\Diamond\phi_{3}^{4}\wedge\Diamond\phi_{3}^{5}}$
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{2}^{3}=\Diamond\phi_{3}^{6}\wedge\Diamond\phi_{3}^{7}\wedge\Diamond\phi_{3}^{8}}$
+[…]
+[…]
+[…]
 $L_{3}:\quad$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{1}=\Diamond(d_{5}\wedge\mathsf{default}}\wedge\bigcirc((\mathsf{carrybin}\;\mathcal{U}\;\mathsf{dispose})\wedge\Diamond\mathsf{default}))$
 $\quad\quad\quad\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\wedge\square(\mathsf{carrybin\Rightarrow\neg\mathsf{public}})}$
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{2}=\Diamond(d_{5}\wedge\mathsf{emptybin}\wedge\bigcirc(d_{5}\wedge\mathsf{default})})$
+$\phi_{3}^{2}=\Diamond(g\wedge\bigcirc(g\wedge\mathsf{emptybin})\wedge\Diamond(d_{5}\wedge\bigcirc(d_{5}\wedge\mathsf{default})))$
 ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{3}^{3}=\Diamond(p\wedge\mathsf{carry}\;\mathcal{U}\;(d_{10}\wedge\bigcirc\neg\mathsf{carry}))}\wedge\mathsf{notpublic}$ (10)
@@
 ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{4}^{3}=\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})}$
-scenario $l_{\text{flat}}$ $l_{\text{hier}}$ ${\mathcal{A}}_{\text{flat}}$ ${\mathcal{A}}_{\text{hier}}$ $t_{\text{flat}}$ $t_{\text{hier}}$ $c_{\text{flat}}$ $c_{\text{hier}}$
+If the objective is to accomplish either scenario, this can be indicated by replacing the formula at level $L_{1}$ with $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2}\vee\phi_{2}^{3})$.
+Furthermore, the expression $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\vee\phi_{2}^{2})\wedge\Diamond\phi_{2}^{3}$ specifies that either task 1 or task 2 must be completed, in addition to task 3.
+scenario $l_{\text{std}}$ $l_{\text{hier}}$ ${\mathcal{A}}_{\text{std}}$ ${\mathcal{A}}_{\text{hier}}$ $t_{\text{std}}$ $t_{\text{hier}}$ $c_{\text{std}}$ $c_{\text{hier}}$
 1 18 19 (17, 39) (12, 13) 14.1$\pm$3.7 4.9$\pm$2.3 71.0$\pm$6.3 69.0$\pm$5.7
@@
 2 $\wedge$ 3 60 90 (6056, 325745) (54, 85) timeout 42.5$\pm$16.9 175.4$\pm$6.9
-1 $\wedge$ 2 $\wedge$ 3 79 111 timeout (70, 112) timeout 89.6$\pm$28.1 246.6$\pm$8.3
-TABLE III: The comparative analysis focuses on two different types of sc-LTL specifications.
-We denote the lengths of the flat and hierarchical sc-LTL specifications as $l_{\text{flat}}$ and $l_{\text{hier}}$, respectively.
-The sizes of the corresponding NBAs are represented by ${\mathcal{A}}_{\text{flat}}$ and ${\mathcal{A}}_{\text{hier}}$, which detail the number of nodes and edges, with the node count listed first.
-In terms of solutions, the runtimes for the flat and hierarchical sc-LTL specifications are indicated by $t_{\text{flat}}$ and $t_{\text{hier}}$, respectively.
-Additionally, the plan horizons, or the lengths of the solutions, for the flat and hierarchical specifications are denoted by $c_{\text{flat}}$ and $c_{\text{hier}}$, respectively.
+1 $\wedge$ 2 $\wedge$ 3 79 111 (70, 112) timeout 89.6$\pm$28.1 246.6$\pm$8.3
+(1 $\vee$ 2) $\wedge$ 3 79 110 (66, 98) timeout 71.5$\pm$17.6 213.1$\pm$23.4
+1 $\vee$ 2 $\vee$ 3 79 109 (64, 94) timeout 63.4$\pm$20.9 163.9$\pm$40.5
+TABLE III: The comparative analysis focuses on two different types of LTLf specifications.
+We denote the lengths of the standard and H-LTLf specifications as $l_{\text{std}}$ and $l_{\text{hier}}$, respectively.
+The sizes of the corresponding NFAs are represented by ${\mathcal{A}}_{\text{std}}$ and ${\mathcal{A}}_{\text{hier}}$, which detail the number of nodes and edges, with the node count listed first.
+In terms of solutions, the runtimes for the standard and H-LTLf specifications are indicated by $t_{\text{std}}$ and $t_{\text{hier}}$, respectively.
+Additionally, the plan horizons, or the lengths of the solutions, for the standard and hierarchical specifications are denoted by $c_{\text{std}}$ and $c_{\text{hier}}$, respectively.
 VIII-C Ablation Study on Effects of Heuristics
@@
 VIII-D Comparison with Existing Works
-We use $m=6$ robots and compare our method with the approach in [8], both using all heuristics.
-Note that [8]’s method is tailored to flat sc-LTL specifications.
+We use $m=6$ robots and compare with the approach in [9], both using all heuristics.
+Note that [9]’s method is tailored to standard LTLf specifications.
 Due to the absence of open-source code, we implemented their method.
@@
 The performance, in terms of average runtimes and costs over 20 runs, is detailed in Tab.
-III and includes the length of formulas and sizes of automatons.
+III and includes the length of formulas and sizes of automata.
 The length of a formula is the total number of logical and temporal operators.
-We observed that, contrary to simulation findings in [42], where hierarchical sc-LTL shows syntactical conciseness compared to flat sc-LTL, our hierarchical sc-LTL specifications are longer in length.
-We hypothesize that the syntactical succinctness of hierarchical sc-LTL becomes more evident in scenarios with complex interdependencies between tasks at non-bottom levels, which are more challenging to express through flat sc-LTL specifications.
-This is not the case in our scenarios, where tasks at non-bottom levels are temporally independent.
-Upon reviewing the results, [8]’s method failed to generate solutions for the last four tasks within the one-hour limit.
+Upon reviewing the results, [9]’s method failed to generate solutions for the last seven tasks within the one-hour limit.
 For tasks 1 and 2, our method produced solutions more quickly and with comparable costs.
-The failure of  [8]’s approach is attributed to the excessively large automatons it generates, sometimes with hundreds of thousands of edges, e.g., 325745 edges for scenario $2\wedge 3$, making the computation of the decomposition set time-consuming as it requires iterating over all possible runs.
-For the most complex scenario $1\wedge 2\wedge 3$, generating an automaton within one hour is impossible.
+The failure of  [9]’s approach is attributed to the excessively large automata it generates, sometimes with hundreds of thousands of edges, e.g., 325745 edges for scenario $2\wedge 3$, making the computation of the decomposition set time-consuming as it requires iterating over all possible runs.
+For the most complex scenario, such as $1\wedge 2\wedge 3$, generating an automaton within one hour is impossible.
 In contrast, our method was able to find a solution in around 90 seconds.
+Moreover, considering the last three tasks, which require completing all, two, or only one task, both the runtime and cost decrease as the number of required tasks is reduced.
 flattened specifications ${\mathcal{A}}_{\text{hier}}$ $t_{\text{hier}}$ $c_{\text{hier}}$
@@
 VIII-E Effects of Granularity of Hierarchy
-For the same task, the performance of the planning algorithm is influenced by the granularity of hierarchical specifications, given the variability in humans’ levels of hierarchical reasoning granularity.
+For the same task, the performance of the planning algorithm is influenced by the granularity of hierarchical specifications.
 To investigate this, we evaluate the most complex task, which combines scenarios 1, 2, and 3 as defined in (VIII-B4).
@@
 We then progressively increase the coarseness by flattening more non-leaf specifications through the merging of their respective leaf specifications, thus gradually transforming the hierarchical structure towards a more flattened form.
-Eventually, this process results in a single flat specification.
+Eventually, this process results in a single standard specification.
 For example, flattening specification $\phi_{3}^{6}$ results in $\phi_{3}^{6}=\Diamond(m_{1}\wedge\mathsf{photo})\wedge\Diamond(m_{4}\wedge\mathsf{photo})\wedge\Diamond(m_{6}\wedge\mathsf{photo})\wedge\square(\neg\mathsf{meeting}\Rightarrow\neg\mathsf{camera})$.
@@
 IV.
-Generally, as more non-leaf specifications are flattened, the automatons increase in size, resulting in longer runtimes, while the costs remain relatively stable.
+Generally, as more non-leaf specifications are flattened, the automata increase in size, resulting in longer runtimes, while the costs remain relatively stable.
 Consequently, finer granularity in the hierarchy is associated with enhanced planning performance.
@@
 The statistical findings are presented in Fig.
-9.
+10.
 Across different scenarios, a consistent pattern emerged: as the number of robots increased, the runtimes tended to increase, while the overall costs decreased.
@@
 Additionally, 20 robots are capable of handling tasks of this complexity effectively, with the marginal benefit of increasing to 30 robots being minor.
-Figure 10: The presented keyframes illustrate the execution of the derivative task described in Eq.
+Figure 11: The keyframes illustrate the execution of the derivative task described in Eq.
 (VIII-G) within the AI2-THOR framework, indicating the crucial phases in the completion of task.
 Each of the four robots is inside yellow boxes with distinct numbers.
-The upper-left corner of the keyframes marks the time steps, and the objects being manipulated are encircled in blue.
-Note that the robots 2 and 4 are simultaneously manipulating the box and the laptop.
+The upper-left corner of the keyframes indicates the time instants, while the bottom-left corner displays the working robot and its corresponding action.
+The working robots are enclosed in blue boxes, with dashed lines in a magenta gradient color connecting the position where they performed most recent manipulation action to their current location.
+White arrows highlight the objects with which the robots are interacting.
+Robots 2 and 4 are simultaneously manipulating the box and the laptop.
 VIII-G Evaluation on simulator AI2-THOR
@@
 After that, move the box to a sofa and the laptop to an armchair.
-Finally, dispose of the credit card in the garbage can.” The hierarchical sc-LTL formulas are:
+Finally, dispose of the credit card in the garbage can.” The H-LTLf formulas are:
 $L_{1}:\quad$ $\phi_{1}^{1}=\Diamond(\phi_{2}^{1}\wedge\Diamond(\phi_{2}^{2}\wedge\Diamond\phi_{2}^{3}))$
@@
 In this part, we modified the goal (4) as
-$J(\tau)=\omega\max_{r\in[N]}c_{r}+(1-\omega)\sum_{r=1}^{N}c_{r}$ (11)
+$J(\tau)=\omega\max_{r\in[N]_{+}}c_{r}+(1-\omega)\sum_{r=1}^{N}c_{r}$ (12)
 where $\omega\in[0,1]$.
 Here, setting $\omega=0$ simplifies (12) to objective (4), while $\omega=1$ seeks to reduce the maximum cost among robots, with the secondary term regularizing the costs of the remaining robots.
-We set $\omega=0.9$ to prioritize minimizing the maximum path length among the robots.
+In the evaluation, we set $\omega=0.9$ to prioritize minimizing the completion time (makespan).
 The results, detailed in Table V, indicate an increase in runtime for solutions as the number of base tasks or robots increases.
-Additionally, with more robots engaged, the maximum cost per robot is reduced due to the ability to distribute tasks more evenly among multiple robots.
+Additionally, with more robots engaged, the completion time is reduced due to the ability to distribute tasks more evenly among multiple robots.
 The execution of derivative task (VIII-G) is illustrated in Fig.
-10.
-# base tasks # robots $t$ $c$
-1 1 2.6$\pm$0.9 11.7$\pm$8.7
-2 3.0$\pm$1.2 11.7$\pm$9.8
-4 4.1$\pm$1.3 9.0$\pm$7.6
-2 1 4.3$\pm$2.9 23.4$\pm$21.0
-2 4.8$\pm$3.2 17.1$\pm$15.4
-4 6.5$\pm$4.7 14.8$\pm$12.7
-3 1 10.8$\pm$7.7 38.2$\pm$25.4
-2 12.1$\pm$9.2 27.9$\pm$21.6
-4 14.6$\pm$9.1 22.0$\pm$14.2
-4 1 13.0$\pm$13.9 50.1$\pm$29.6
-2 13.3$\pm$14.9 30.7$\pm$16.9
-4 14.9$\pm$11.4 23.9$\pm$14.0
-TABLE V: Results comparing performance across different numbers of base tasks and robots.
-Here, $t$ represents the runtimes required to return solutions, and $c$ denotes the maximal cost incurred by any robot.
+11.
+# robots # tasks $t$ $c$ tasks $t$ $c$
+1 1 2.6$\pm$0.9 11.7$\pm$8.7 3 10.8$\pm$7.7 38.2$\pm$25.4
+2 3.0$\pm$1.2 11.7$\pm$9.8 12.1$\pm$9.2 27.9$\pm$21.6
+4 4.1$\pm$1.3 9.0$\pm$7.6 14.6$\pm$9.1 22.0$\pm$14.2
+1 2 4.3$\pm$2.9 23.4$\pm$21.0 4 13.0$\pm$13.9 50.1$\pm$29.6
+2 4.8$\pm$3.2 17.1$\pm$15.4 13.3$\pm$14.9 30.7$\pm$16.9
+4 6.5$\pm$4.7 14.8$\pm$12.7 14.9$\pm$11.4 23.9$\pm$14.0
+TABLE V: Results across different numbers of base tasks and robots. Here, $t$ represents the runtimes required to return solutions, and $c$ denotes the completion time.
 IX Conclusion and Discussion
-In this work, we addressed the problem of STAP for multiple robots subject to hierarchical sc-LTL specifications.
-First, we presented the syntax and semantics for hierarchical sc-LTL.
-Next, we developed a search-based planning algorithm.
-To further expedite the search, we introduced several heuristics leveraging the task structure.
-We also provided theoretical analysis on the completeness and optimality.
+We addressed the problem of STAP for multiple robots subject to H-LTLf specifications.
+First, we presented the syntax and semantics for H-LTLf.
+Next, we developed a search-based planning algorithm, and introduced several heuristics leveraging the task structure to expedite it.
+We provided theoretical analysis on the completeness and optimality.
 Our simulations, focused on service tasks involving navigation and manipulation, included an ablation study on heuristics.
@@
 III.
-We contend that hierarchical specifications do not necessarily increase cognitive load, given that humans naturally excel at hierarchical thinking [6, 7].
-The results from the user study also highlighted the user-friendliness of hierarchical sc-LTL.
-The second factor is the design of the planning algorithm, particularly its heuristics.
+The results from the user study also highlighted the user-friendliness of H-LTLf.
+The second factor is the planning algorithm, particularly its heuristics.
 As shown in Tab.
@@
 Furthermore, employing a combination of all three heuristics results in shorter runtimes compared to using any single heuristic.
-While the hierarchical sc-LTL introduced has proven effective, there are still several areas that require additional investigation.
-Firstly, the planning algorithm assumes that robots operate independently, without considering scenarios where coordination, such as jointly carrying heavy items, is necessary.
-Therefore, it is crucial to include such situations in the algorithm.
-Another area of interest is addressing uncertainty arising from multiple sources, including partially observable environments, or the costs of actions, which might not be predetermined.
-Since hierarchical sc-LTL can encompass multiple conflicting specifications, a future research avenue could be to define and maximize degree of satisfaction when all specifications cannot be concurrently met.
-Lastly, integrating natural languages, as in work [72], to embed task hierarchies from human instructions into structured planning languages suitable for existing planners, presents an intriguing prospect.
+While the H-LTLf introduced has proven effective, there are several areas that require investigation.
+An open challenge is developing an algorithmic approach to convert standard LTLf formulas into H-LTLf formulas and exploring how to introduce the hierarchical form to full LTL.
+The second aspect relates to the planning algorithm.
+Currently, the planning algorithm assumes that robots operate independently, without considering scenarios where coordination, such as jointly carrying heavy items, is necessary.
+Therefore, it is crucial to include such situations, as in work [71].
+Since H-LTLf can encompass multiple conflicting specifications, a future direction could be to define and maximize degree of satisfaction when not all specifications can be met.
+Finally, integrating natural languages, as in work [72], to embed task hierarchies from instructions into structured formal languages, presents an intriguing prospect.